### v1

# Long-range nonstabilizerness from topology and correlation

Abstract
Understanding nonstabilizerness (aka quantum magic) in many-body quantum systems, particularly its interplay with entanglement, represents an important quest in quantum computation and many-body physics. Drawing natural motivations from the study of quantum phases of matter and entanglement, we systematically investigate the notion of long-range magic (LRM), defined as nonstabilizerness that cannot be erased by constant-depth local circuits. By establishing connections to the Bravyi–König theorem concerning the limitation of fault-tolerant logical gates, we show that certain families of topological stabilizer code states exhibit LRM. Then, we show that all ground states of topological orders that cannot be realized by topological stabilizer codes, such as Fibonacci topological order, exhibit LRM, which can be viewed as a “no lowest-energy trivial magic” result. Building on our considerations of LRM, we discuss the classicality of short-range magic from e.g. preparation and learning perspectives, and put forward a “no low-energy trivial magic” (NLTM) conjecture that has key motivation in the quantum PCP context. We also connect two-point correlations with LRM, demonstrating certain LRM state families by correlation properties. Most of our proof techniques do not depend on geometric locality and can potentially be extended to systems with general connectivity. Our study leverages and sheds new light on the interactions between quantum resources, coding and fault tolerance theory, complexity theory, and many-body physics.
I Introduction
A pivotal mission of quantum information science is to understand and characterize the origin of potential advantages of quantum computers over classical ones. As indicated by the Gottesman–Knill theorem Gottesman (1998); Aaronson and Gottesman (2004); Nielsen and Chuang (2011), stabilizer quantum computation can be efficiently simulated by classical methods. Consequently, nonstabilizerness (magic) Bravyi and Kitaev (2005); Veitch et al. (2014) can be understood as a “resource” for quantum computational speedups, with its characterizations and meanings having been actively studied from e.g. resource theory Veitch et al. (2014); Howard and Campbell (2017) and classical simulation Bravyi et al. (2016); Bravyi and Gosset (2016); Bravyi et al. (2019); Seddon et al. (2021); Bu and Koh (2019) perspectives for a long time. Recently, nonstabilizerness in many-body quantum systems has attracted substantial interest from both quantum information and physics communities, with extensive efforts devoted to quantitatively characterizing many-body nonstabilizerness and its interplay with entanglement (see e.g. Refs. Liu and Winter (2022); Leone et al. (2022); Haug and Piroli (2023); Lami and Collura (2023); Tarabunga et al. (2024); Frau et al. (2024); Gu et al. ) and understanding many-body physical phenomena through the lens of nonstabilizerness (see e.g. Refs. Sarkar et al. (2020); White et al. (2021); Liu and Winter (2022); Tarabunga et al. (2023); Gu et al. ).
In the development of condensed matter physics, the concept of long-range entanglement Chen et al. (2010); Wen (2013); Hastings and Wen (2005) has played a crucial role, underpinning intrinsically quantum notions of order and phases of matter that extend beyond the traditional Landau paradigm. Here “long-range” captures the global, topological nature of the feature, requiring it to be preserved under constant (equivalently, $\mathcal{O}(1)$ or finite) depth local quantum circuits representing “smooth” deformations that do not alter fundamentally global properties.
Drawing inspiration from this, we adopt a computational perspective and consider long-range magic (LRM) White et al. (2021); Ellison et al. (2021) defined as nonstabilizerness that cannot be erased by constant-depth local circuits, which capture the “topological” type of nonstabilizerness embedded in global structures and robust against local perturbations. Several insights are worth noting. First, it is evident that LRM is a strictly stronger notion than long-range entanglement in the sense that the former is a sufficient but not necessary condition for the latter. As a result, LRM can induce a more refined classification of phases in the long-range entangled regime with computational underpinnings. Second, from a computational complexity perspective, LRM can be understood as high-complexity nonstabilizerness beyond efficient classical computation, thus holding natural relevance in quantum complexity theory, particularly to the prominent quantum probabilistically checkable proofs (PCP) conjecture Aharonov et al. . Third, as will be seen, the study of LRM bridges the quantum error correction, fault tolerance, computational resources, and many-body physics in new interesting ways, drawing on insights from these fields to inform one another. Given the wide importance of these areas and connections, systematic theories and constructions of LRM would be highly valuable but have so far remained largely elusive.
In this work, we address this gap by presenting a rigorous theory of LRM and studying its various aspects including construction and significance from diverse perspectives. Note that establishing LRM essentially entails the challenging problem of proving nontrivial circuit lower bounds for erasing nonstabilizerness. First, we demonstrate examples of LRM rooted in topological codes and topologically ordered systems, highlighting the insight that long-range properties can originate from topology. Leveraging Bravyi and König’s insight in the fault tolerance context Bravyi and König (2013) that certain logical operations cannot be realized by constant-depth local circuits in topological stabilizer codes, we can show that they generate LRM state families, a basic example being logical nonstabilizer states in the 2D toric code. Furthermore, utilizing the local indistinguishability property of the ground states of general topologically ordered systems Bravyi et al. (2006, 2010), we establish general conditions for “LRM phases” in which the entire ground spaces of entire phases of topological orders have LRM. Examples include non-Abelian topological orders based on arguments in the literature that they cannot be realized by stabilizer models Potter and Vasseur (2016); Ellison et al. (2022). We can also establish a ground space dimension condition which, for instance, indicate a robust notion of LRM for Fibonacci topological order on certain manifolds. Then, we point out that LRM can stem from correlation properties: by characterizing the restrictions on the correlations between two distant subsystems in SRM states—specifically, the correlations resemble those of EPR pairs—we show by the violation of these restrictions that simple families of states including $\mathrm{C}^{n-1}Z$ states and GHZ-type states that exhibit LRM. After presenting these LRM constructions, we discuss in more depth the classical characteristics of systems without LRM from various perspectives including preparation, description, and learning complexities. In particular, building upon our LRM phase results, we formulate a “no low-energy trivial magic” (NLTM) conjecture, which refines the recently validated no low-energy trivial state (NLTS) theorem Freedman and Hastings ; Anshu et al. (2023) by unifying the trivial entanglement and nonstabilizerness types of efficient classical proofs, potentially moving closer to quantum PCP.
II Preliminaries
A family of constant-depth local circuits consists of unitary gates acting on an increasing number of qubits. These circuits are formed by a constant number of layers of local gates, specifically two-qubit gates applied to neighboring qubits. The connectivity of the qubits depends on the specific model under consideration.
Definition 1.
A family of states $\{|\psi_{n}\rangle\}$, where $|\psi_{n}\rangle$ is an $n$-qubit state, is said to have long-range magic (LRM), if for any family of constant-depth local circuits $\{U_{n}\}$, $U_{n}|\psi_{n}\rangle$ remains a magic state for sufficiently large $n$. Conversely, say $\{|\psi_{n}\rangle\}$ has short-range magic (SRM) if there exists a family of constant-depth local circuits $\{U_{n}\}$ such that $\{U_{n}|\psi_{n}\rangle\}$ are stabilizer states. Furthermore, $\{|\psi_{n}\rangle\}$ is said to have strong LRM, if for any family of constant-depth local circuits $\{U_{n}\}$ and any family of stabilizer states $\{|S_{n}\rangle\}$, it holds that $\big{\|}U_{n}|\psi_{n}\rangle\langle\psi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=\Omega(1/n)$.
Note that if $\{|\psi_{n}\rangle\}$ doesn’t have LRM, there exists a subsequence $\{n_{q}\}_{q=1}^{\infty}\subset\mathbb{Z}^{+}$, such that $\{|\psi_{n_{q}}\rangle\}_{q=1}^{\infty}$ has SRM. If $\{|\psi_{n}\rangle\}$ has LRM, then it must have long-range entanglement, meaning they cannot be transformed into product states by constant-depth local circuits. We sometimes say a state $|\psi\rangle$ has LRM, implicitly referring to a family of such states. The existence of LRM does not depend on the choice of local bases.
The Clifford hierarchy for $n$-qubits is defined as follows. The first level consists of the Pauli group, $\mathcal{C}_{n}^{(1)}:=\mathcal{P}_{n}$. For $l\geq 2$, the $l$-th level is recursively defined as $\mathcal{C}_{n}^{(l)}:=\{U\in\operatorname{U}(2^{n})\mid UPU^{\dagger}\in\mathcal{C}_{n}^{(l-1)},\forall P\in\mathcal{P}_{n}\}$, where $\operatorname{U}(2^{n})$ is the unitary group.
A $D$-dimensional topological stabilizer code (TSC) family is a family of codes that encode a fixed number of logical qubits. The physical qubits are arranged on a lattice embedded in a $D$-dimensional manifold, where the Pauli stabilizers have constant-diameter support. The code distance can be made arbitrarily large by increasing the lattice size.
Figure 1: (a) A set of $n$ stabilizer generators of the logical state $|\overline{T0}\rangle$ in toric code, one of them is non-Pauli. We define $|T\rangle:=\operatorname{cos}(\pi/8)|0\rangle+\operatorname{sin}(\pi/8)|1\rangle$, the $+1$ eigenstate of $\frac{1}{\sqrt{2}}(X+Z)$. The Pauli generators (blue) consist of $n-2$ local ones and a global $\overline{Z}_{2}$, which is a tensor product of Pauli $Z$’s around the blue handle. The orange one is a global non-Pauli generator $\frac{1}{\sqrt{2}}(\overline{X}_{1}+\overline{Z}_{1})$, which introduces global $\frac{1}{\sqrt{2}}$’s into the Pauli spectrum of $|\overline{T0}\rangle$, producing “topological nonstabilizerness”. Tracing out the entire region outside an arbitrary gray “cross” will not decrease the amount of nonstabilizerness contained in $|\overline{T0}\rangle$. (b) The bipartite reduced density matrix of a state with SRM on distant subsystems $A$ and $B$ is only correlated by a stabilizer state $|S\rangle$, as the light cones of $A$ and $B$ do not intersect.
III Long-range magic from topological codes
Our first observation is that LRM state families can arise naturally from TSCs, which provide topological protection for logical information and can thereby protect logical nonstabilizerness against local perturbations. The restriction of “topology” and “constant-depth local” provides a nice connection between the LRM property and the limitations on fault-tolerant logical gates established by Bravyi–König Bravyi and König (2013). Leveraging this observation, we show that:
Theorem 1.
For a $D$-dimensional ($D\geq 2$) TSC family, if $|\phi\rangle\notin\big{\{}V^{\dagger}|0^{k}\rangle\mid V\in\mathcal{C}_{k}^{(D)}\big{\}}$, where $k$ is the number of logical qubits, then the corresponding logical state family $|\overline{\phi}\rangle$ has LRM.
We demonstrate the proof of this theorem through a concrete example of toric code, which captures the essential ideas. Full details can be found in Appendix A.
Example 1.
The logical nonstabilizer state $|\overline{T0}\rangle$ in toric code has LRM.
All code states in the toric code Kitaev (2003a) on a $L\times L$ lattice ($n=2L^{2}$ physical qubits, $2$ logical qubits) has long-range entanglement, requiring local circuits of depth $\Theta(L)$ to prepare Dennis et al. (2002); Bravyi et al. (2006) from a product state. Notice that while $|\overline{T0}\rangle$ has circuit depth complexity $\Theta(L)$, we cannot immediately rule out the possibility that it is mapped to a stabilizer state, thus losing all magic, by a constant-depth local circuit. The “topological nonstabilizerness” of $|\overline{T0}\rangle$ can be visualized by its Pauli spectrum. Recall that for any stabilizer state $|\phi\rangle$ and Pauli string $P$ with $+1$ phase, we have $\langle\phi|P|\phi\rangle=\pm 1$ when ±$\pm P$ belongs to the stabilizer group of $|\phi\rangle$, and $\langle\phi|P|\phi\rangle=0$ otherwise. Due to the global non-Pauli stabilizer generator $\frac{1}{\sqrt{2}}(\overline{X}_{1}+\overline{Z}_{1})$ for $|\overline{T0}\rangle$ (see Fig. 1(a)), we find $\langle\overline{T0}|P|\overline{T0}\rangle=\pm\frac{1}{\sqrt{2}}$ for exponentially many $P$ with support wrapping around the torus. This suggests that a constant-depth local circuit on the torus cannot organize its Pauli spectrum to be $-1,0,1$ to match that of a stabilizer state.
We prove $|\overline{T0}\rangle$ in toric code has LRM by contradiction. Suppose on the contrary that $|\overline{T0}\rangle$ does not have LRM. Then, for any lattice size $L_{q}$ in a subsequence of sizes $\{L_{q}\}_{q=1}^{\infty}\subset\mathbb{Z}^{+}$, there exists a constant-depth local circuit $U$ such that $U|\overline{T0}\rangle=|S\rangle$, where $|S\rangle$ is a stabilizer state. Denote $C$ as a constant that is independent of $L_{q}$, which upper bounds the diameter of the $n-2$ deformed local generators $\{UA_{v}U^{\dagger}\}$, $\{UB_{p}U^{\dagger}\}$. There are $3$ “brother” states of $|S\rangle=U|\overline{T0}\rangle$, namely $U\overline{Y}_{1}|\overline{T0}\rangle$, $U\overline{Y}_{2}|\overline{T0}\rangle$, and $U\overline{Y}_{1}\overline{Y}_{2}|\overline{T0}\rangle$. These brother states and $|S\rangle$ are mutually orthogonal, and their reduced density matrices are identical when restricted to systems of diameter $\leq C$. At most $3$ such brother states of $|S\rangle$ exist since these states are all invariant under the action of $n-2$ deformed generators $\{UA_{v}U^{\dagger}\}$, $\{UB_{p}U^{\dagger}\}$.
Now we show that the deformed code space $U\Pi U^{\dagger}$ is also a TSC, where $\Pi$ denotes the projector onto the original toric code. Consider the stabilizer group $G$ of $|S\rangle$. We claim that there must exist $n-2$ independent Pauli strings in $G$ with diameters $\leq C$, where “independent” means they cannot multiply to $\mathbb{I}_{2^{n}}$. Suppose for example that, there exists and exists at most $n-3$ such local Pauli strings, then we can construct $7$ brother states for $|S\rangle$ by finding $3$ Pauli strings $Q_{1},Q_{2},Q_{3}$ in $G$, satisfying that $Q_{1},Q_{2},Q_{3}$ together with the $n-3$ local Pauli strings generate $G$. The support of $Q_{1},Q_{2},Q_{3}$ must have diameter $>C$ by assumption. We can then examine the states stabilized by these $n-3$ local Pauli strings and ±$\pm Q_{1},\pm Q_{2},\pm Q_{3}$, get $7$ brother states for $|S\rangle$ and leading to a contradiction. Therefore, we proved the claim that there are $n-2$ independent $C$-local Pauli strings that stabilize $|S\rangle$. Since the deformed code states in $U\Pi U^{\dagger}$ are locally indistinguishable from $|S\rangle$, these $n-2$ local Pauli strings also fixed other code states. Additionally, the constant-depth local circuit $U$ will degrade the distance of toric code, $L_{q}$, by a constant factor, so the deformed code family $U\Pi U^{\dagger}$ also has macroscopic distance.
Now we know that $U$ is a constant-depth local circuit connecting toric code and another TSC, such that the logical magic state $|\overline{T0}\rangle$ is mapped to a stabilizer state $|S\rangle$, thus a logical stabilizer state in the deformed TSC. However, this is impossible by the restriction of constant-depth local circuits mapping between $2$D TSCs. The celebrated result of Bravyi and König goes as follows:
Lemma 2 (Bravyi–König Bravyi and König (2013)).
Suppose two $D$-dimensional ($D\geq 2$) TSCs are defined on the same lattice, both have $k$ logical qubits, with code projectors $\Pi_{1}=J_{1}J_{1}^{\dagger}$ and $\Pi_{2}=J_{2}J_{2}^{\dagger}$ respectively. If a constant-depth local unitary $U$ satisfies $U\Pi_{1}U^{\dagger}=\Pi_{2}$, then the induced logical unitary $J_{2}^{\dagger}UJ_{1}$ belongs to $\mathcal{C}_{k}^{(D)}$, when the lattice size is sufficiently large.
Specifically, for $2$D TSCs, a constant-depth local circuit can only induce a logical Clifford between them, which makes it impossible to transform $|\overline{T0}\rangle$ into $|S\rangle$. Therefore, we conclude that $|\overline{T0}\rangle$ in toric code has LRM.
We comment that LRM can be generated by only one non-Clifford gate: $|\overline{T0}\rangle$ can be obtained by first preparing 2-qubit state $|T0\rangle$ with a non-Clifford gate $T=\operatorname{diag}(1,e^{i\pi/4})$, and then perform the Clifford encoding circuit for toric code. We also notice an intriguing property of topological nonstabilizerness of TCSs: tracing out qubits from the union of several contractable regions on the torus (see Fig. 1(a)) does not reduce the amount of nonstabilizerness in $|\overline{T0}\rangle$, because the erasure of such regions is correctable by Clifford recovery operations.
IV Long-range magic from topological orders
The study of LRM provides new insights into topological orders, which are quantum phases of matter characterized by long-range entanglement that depend on the system’s topology Ellison et al. (2022). We focus on topological orders realized on qubits in this section, while our result also holds for qudits in prime dimensions, on which stabilizer codes can be well-defined. Toric code serves as a representative example of topological order, and all translation invariant 2D TSCs belong to the same phase as the decoupled copies of toric codes Bombin et al. (2012); Bombín (2014); Haah (2021). The LRM/SRM property of states in a topological order further reveals the many-body entanglement structure and can be instrumental for the classification of quantum phases.
The quantum double model Kitaev (2003b) and the string-net model Levin and Wen (2005) provide local Hamiltonian realizations of topological orders, describing the topological order by ground spaces of local Hamiltonians. Utilizing the TQO condition Bravyi et al. (2006, 2010) which asserts that the ground space describing a topological order forms a quantum error-correcting code with macroscopic distance, we can demonstrate that if a family of states in ground space exhibits SRM–that is, they are mapped to stabilizer states by constant-depth local circuits–then the entire ground space is mapped to a TSC by same circuits. This implies that the topological order can be described by this TSC. We arrive at the following theorem, with the rigorous proof provided in Appendix B.
Theorem 3.
A topological order cannot be realized by any TSC, if and only if for any local Hamiltonian realization of this topological order, all ground states exhibit LRM.
Previous arguments that stabilizer model cannot realize non-Abelian topological orders Potter and Vasseur (2016); Ellison et al. (2022) imply that ground states in the entire phases of all non-Abelian topological orders have LRM. As a direct corollary of Theorem 3, a topological order can be realized by a TCS, if and only if it admits a stabilizer state (i.e., there exists a local Hamiltonian realization of this order whose ground space has a stabilizer state). Based on the difference in ground space degeneracy between a topological order and TSCs, we can further rigorously establish a stronger result:
Theorem 4.
Suppose the ground space dimension of a topological order is not a power of 2, then for any local Hamiltonian realization of this topological order, all ground states exhibit strong LRM.
The proof of this theorem is given in Appendix C. Intuitively, the “strong” property is ensured by the energy gap of conventional topological order. Specifically, if a family of ground states $\{|\phi_{n}\rangle\}$ of a topological order described by Hamiltonians $\{H^{(n)}\}$ satisfies $\big{\|}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=o(1/n)$ for constant-depth local circuits $\{U_{n}\}$ and stabilizer states $\{|S_{n}\rangle\}$, then $U_{n}$ maps the ground space of $H^{(n)}$ close to a TSC given by the code state $|S_{n}\rangle$. The $o(1/n)$ distance of $U_{n}|\phi_{n}\rangle$ and $|S_{n}\rangle$ implies all code states of the TSC have energy at most $o(1)$ above the ground energy of the deformed Hamiltonian $U_{n}H^{(n)}U_{n}^{\dagger}$. Since $\{H^{(n)}\}$ is gapped, the logical dimension of the TSC is at most the ground space degeneracy of $H^{(n)}$. The converse also holds similarly, implying that the dimension of the TSC equals the ground space degeneracy of $H^{(n)}$.
Example 2.
(Informal) For the entire phase of Fibonacci topological order on certain manifolds, all ground states exhibit strong LRM.
Consider the Fibonacci code defined on a hollow cylinder $\Sigma_{2}=S^{2}-(A^{1}\cup A^{2})$ with discs $A^{1}$ and $A^{2}$ both labeled by $1$ Koenig et al. (2010). The ribbon graphs on $\Sigma_{2}$ consistent with labeling $(1,1)$ span a subspace $\mathcal{H}_{\Sigma_{2}}^{(1,1)}$ with $\dim(\mathcal{H}_{\Sigma_{2}}^{(1,1)})=3$. The Fibonacci code realized by the ground space of the Levin-Wen Hamiltonian $H=\sum_{\text{plaquettes }p}B_{p}-\sum_{\text{vertices }v}Q_{v}$ is isomorphic to $\mathcal{H}_{\Sigma_{2}}^{(1,1)}$ Koenig et al. (2010). Therefore, as the lattice becomes denser, all ground states of $H$ have strong LRM. Also, the degeneracy of Fibonacci code on closed higher-genus surfaces is not generally a power of $2$ Schotte et al. (2022), implying that their ground states also exhibit strong LRM.
V Long-range magic from correlation
Two-point correlations play a crucial role in studying long-range entanglement Zeng et al. (2015), as they serve to reveal and quantify the quantum correlations between distant points in a many-body system. In this section, we show that two-point correlations are also instrumental in identifying the LRM property, specifically by showcasing the super-constant all-to-all complexities of magic in state families with “nonstabilizer” correlations. We begin by examining the constraints on correlations generated by SRM states and then utilize the violation of these constraints to demonstrate the LRM property in families such as ${|\mathrm{C}^{n-1}Z\rangle}$ and GHZ-type states.
For a quantum circuit $U$ acting on $n$ qubits and a subset $A\subset[n]:=\{1,2,\cdots,n\}$, we define the backward light-cone $\mathcal{L}(A)\subset[n]$ as the set of qubits that can influence the output of $U$ at $A$.
Theorem 5.
For a family of SRM states $\{|\psi_{n}\rangle=U_{n}|S_{n}\rangle\}$, where $\{U_{n}\}$ are constant-depth local circuits and $\{|S_{n}\rangle\}$ are stabilizer states, the following holds: For all subsystems $A,B\subset[n]$ of size $\mathcal{O}(1)$, satisfying $\mathcal{L}(A)\cap\mathcal{L}(B)=\emptyset$, the bipartite reduced state $\operatorname{Tr}_{\overline{A\cup B}}|\psi_{n}\rangle\langle\psi_{n}|$ can be obtained by local operations on $\mathcal{O}(1)$ EPR pairs $|\Phi^{+}\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$ shared by the two parties, where $\overline{A\cup B}:=[n]-(A\cup B)$.
This theorem can be applied to constant-depth circuits with arbitrary connectivity. To prove this theorem, note that (see Fig. 1(b))
$\operatorname{Tr}_{\overline{A\cup B}}|\psi_{n}\rangle\langle\psi_{n}|=\mathcal{E}_{A}\otimes\mathcal{E}_{B}\big{(}\operatorname{Tr}_{\overline{\mathcal{L}(A)\cup\mathcal{L}(B)}}|S\rangle\langle S|\big{)},$ (1)
where $\mathcal{E}_{A}(\sigma):=\operatorname{Tr}_{\mathcal{L}(A)-A}(U_{A}\sigma U_{A}^{\dagger})$ is a completely-positive tracing preserving (CPTP) map form $\mathcal{L}(A)$ to $A$, with $U_{A}$ being the circuit within the light-cone of $A$ (colored orange in Fig. 1(b)), and $\mathcal{E}_{B}$ is defined similarly. $\operatorname{Tr}_{\overline{\mathcal{L}(A)\cup\mathcal{L}(B)}}|S\rangle\langle S|$ is a reduced state of a stabilizer state, and thus is a bipartite stabilizer mixed state Aaronson and Gottesman (2004) on $M=\left|\mathcal{L}(A)\right|+\left|\mathcal{L}(B)\right|=\mathcal{O}(1)$ qubits. It can only create rigid correlations between $\mathcal{L}(A)$ and $\mathcal{L}(B)$, similar to those of EPR pairs, thereby restricting the possible states that can be created by performing local operations on it. By Lemma 13 in Appendix D, $\operatorname{Tr}_{\overline{\mathcal{L}(A)\cup\mathcal{L}(B)}}|S\rangle\langle S|$ can be created by local operations on $\mathcal{O}(1)$ EPR pairs, and thus so can $\operatorname{Tr}_{\overline{A\cup B}}|\psi_{n}\rangle\langle\psi_{n}|$.
Consider the family of hypergraph states $|\mathrm{C}^{n-1}Z\rangle:=\mathrm{C}^{n-1}Z|+^{n}\rangle$, where $\mathrm{C}^{n-1}Z=\operatorname{diag}(1,\cdots,1,-1)$ is the $n$-qubit multi-controlled-$Z$ gate. Such gates are native gates on several physical platforms Kim et al. (2022); Nguyen et al. (2024); Bluvstein et al. (2024); Wang et al. (2024). By measuring $X$ on arbitrary two qubits $\{k,l\}\subset[n]$, we get a correlation $\langle X_{k}X_{l}\rangle:=\langle\mathrm{C}^{n-1}Z|X_{k}X_{l}|\mathrm{C}^{n-1}Z\rangle=1-2^{2-n}$. On single qubits we also have $\langle X_{k}\rangle=\langle X_{l}\rangle=1-2^{2-n}$. Detailed derivations are left to Appendix D.
We demonstrate in Proposition 14 in Appendix D that, for $P$ a single-qubit Pauli operator, $K\geq 1$ EPR pairs can generate a two-qubit state $\rho$ by local CPTP maps, with $\langle P\otimes P\rangle_{\rho}=b$ and $\langle P\otimes\mathbb{I}\rangle_{\rho}=\langle\mathbb{I}\otimes P\rangle_{\rho}=c$, if and only if $(b,c)$ lies within the feasible regions surrounded by the blue lines shown in Fig. 2. For $P=X$ and $\rho$ the two-point reduced density matrix of $|\mathrm{C}^{n-1}Z\rangle$, we have $b=c=1-2^{2-n}$. For arbitrary constant $K$, $(1-2^{2-n},1-2^{2-n})$ can always stay outside the feasible region when $n$ is sufficiently large.
Figure 2: The “feasible region” is the area bounded by the blue lines. If a bipartite two-qubit state $\rho$ has $(b,c)$ coordinate located outside the feasible region, then $\rho$ cannot be obtained by local CPTP maps on $K$ EPR pairs. The blue boundary lines consist of two linear functions and $2^{K}$ quadratic functions. We plot the cases for $K=2,3$ as examples, with the general cases being similar. The feasible region converges to a triangle when $K\rightarrow\infty$. The coordinates $(b,c)=(1-2^{2-n},1-2^{2-n})$ given by $|\mathrm{C}^{n-1}Z\rangle$ are plotted on the orange line $b=c$, which moves outside the feasible region for any fixed $K$ when $n$ is sufficiently large.
Corollary 6.
$\{|\mathrm{C}^{n-1}Z\rangle\}_{n=3}^{\infty}$ has LRM.
Suppose on the contrary that there exists $\{n_{q}\}_{q=1}^{\infty}\subset\mathbb{Z}^{+}$ such that $\{|\mathrm{C}^{n_{q}-1}Z\rangle=U_{n_{q}}|S_{n_{q}}\rangle\}$ has SRM. Denote $M_{q}=\{j\in[n_{q}]:\mathcal{L}(j)\cap\mathcal{L}(1)\neq\emptyset\}$ be the set of qubits whose light-cone with regard to $U_{n_{q}}$ intersects with that of qubit $1$. We have $\left|M_{q}\right|=\mathcal{O}(1)$. By Theorem 5, for all qubits $k\notin M_{q}$, the reduced state $\operatorname{Tr}_{\overline{\{1,k\}}}|\mathrm{C}^{n_{q}-1}Z\rangle\langle\mathrm{C}^{n_{q}-1}Z|$ can be generated by local CPTP maps on $\mathcal{O}(1)$ EPR pairs. However, we know that when $n_{q}$ is sufficiently large, $\operatorname{Tr}_{\overline{\{1,k\}}}|\mathrm{C}^{n_{q}-1}Z\rangle\langle\mathrm{C}^{n_{q}-1}Z|$ cannot be generated by local CPTP maps on $\mathcal{O}(1)$ EPR pairs, for all $k\neq 1$, leading to a contradiction. Therefore, $\{|\mathrm{C}^{n-1}Z\rangle\}$ has LRM.
For GHZ-type state family $\left\{\sqrt{1-\alpha}|0^{n}\rangle+\sqrt{\alpha}e^{i\theta}|1^{n}\rangle\right\}$, we have $\langle Z_{k}Z_{l}\rangle=1$ and $\langle Z_{k}\rangle=1-2\alpha$ for all positions $\{k,l\}\subset[n]$ and all values of $\theta$. When $1-2\alpha$ is not a finite-length binary decimal, the feasible region shown in Fig. 2 does not contain $(1,1-2\alpha)$ for arbitrary $K\geq 1$, implying that this family exhibits LRM.
VI Classicality of short-range magic and NLTM conjecture
A family of SRM states $\{|\psi_{n}\rangle=U_{n}|S_{n}\rangle\}$ admits an efficient classical description by storing the quantum circuits implementing $U_{n}$ and the stabilizer tableaux of $|S_{n}\rangle$. The expectation value of any $\mathcal{O}(1)$-weight observable $O$ can be computed exactly on classical computers in time $\mathrm{poly}(n)$. This follows from the fact that $U_{n}^{\dagger}OU_{n}$ acts nontrivially on only $\mathcal{O}(1)$ qubits, can be decomposed into the sum of $\mathcal{O}(1)$ Pauli strings, and the expectation value of $|S_{n}\rangle$ on Pauli strings can be computed efficiently Aaronson and Gottesman (2004). Consequently, SRM states can serve as classical proofs for quantum PCP Hamiltonians Aharonov et al. . Based on this, we propose the no low-energy trivial magic (NLTM) conjecture, formulated as follows:
Conjecture 1 (NLTM).
There exists a constant $\epsilon>0$ and a family of $\mathcal{O}(1)$-local commuting Hamiltonians $\{H^{(n)}=\sum_{i=1}^{m^{(n)}}H_{i}^{(n)}\}$ acting on $n$ qubits, with $m^{(n)}=\Theta(n)$ and $\|H_{i}^{(n)}\|_{\infty}\leq 1$, such that any family of states $\{|\psi_{n}\rangle\}$ satisfying $\langle\psi_{n}|H^{(n)}|\psi_{n}\rangle<\lambda_{\min}(H^{(n)})+\epsilon n$ has LRM.
As demonstrated in Appendix E, NLTM is a necessary consequence of the quantum PCP conjecture. A resolution of NLTM would provide further insights and evidence for quantum PCP. We extend the notion of SRM states to mixed SRM states, defined as the reduced density matrices of pure SRM states. Like pure SRM states, mixed SRM states can also serve as classical proofs. NLTM is strictly stronger than NLTS Freedman and Hastings ; Anshu et al. (2023), as SRM states, which serve as classical proofs, generalize short-range entangled states. In a sense, NLTM unifies different forms of efficient classical proofs enabled by trivial entanglement and classical computational complexity Nirkhe (2022); Coble et al. (2023a), thereby providing a more refined condition for the quantum PCP conjecture. Our findings on Fibonacci topological order provide a realization of “no lowest-energy trivial magic”, which may be further strengthened to fully satisfy the NLTM conjecture potentially by generalizing Fibonacci-type topological orders to systems with general connectivity.
A family of SRM states can be prepared efficiently. If the graph representing the connectivity of qubits is connected, any $n$-qubit Clifford unitary can be synthesized using $\mathcal{O}(n^{2})$ gates Aaronson and Gottesman (2004); Nash et al. (2020). Therefore, all $n$-qubit states with SRM can be prepared with $\mathcal{O}(n^{2})$ gates, and states with gate complexity $\omega(n^{2})$ exhibit LRM. With the gate complexity upper bounded by $\mathcal{O}(n^{2})$, it follows from Ref. Zhao et al. (2024) that all SRM states can be learned to $\epsilon$-precision in trace distance with high probability, using sample complexity $\widetilde{\mathcal{O}}(n^{2}\epsilon^{-2})$. Note that even with all-to-all connectivity, there exist states that require $\Omega(2^{n})$ gates Plesch and Brukner (2011); Sun et al. (2023) to prepare. Haar random states exhibit LRM with probability approaching $1$ as $n\rightarrow\infty$. All-to-all random quantum circuits (Chen et al. (2024) Corollary 1.9) with $\omega(n^{6})$ gates can create a state with gate complexity $\omega(n^{2})$, and thus LRM, with probability approaching $1$.
We say $\{|\psi_{n}\rangle\}$ has $\mathrm{SRM}_{0}$, if their magic can be removed by tensor products of single-qubit gates. For example, the product of single-qubit magic states, the logical $|T\rangle^{\otimes k}$ state in codes with transversal $T$ gates, and “no low-energy stabilizer states” constructed by Coble et al. (2023b) have $\mathrm{SRM}_{0}$. For any $n$-qubit state $\rho$, unitary $U=U_{1}\otimes\cdots\otimes U_{n}$ and any subset $A\subset[n]$, we have $\sum_{P:\operatorname{supp}(P)=A}\operatorname{Tr}(P\rho)^{2}=\sum_{P:\operatorname{supp}(P)=A}\operatorname{Tr}(PU\rho U^{\dagger})^{2}$. Therefore, for $|\psi_{n}\rangle$ with $\mathrm{SRM}_{0}$, $\sum_{P:\operatorname{supp}(P)=A}\langle\psi_{n}|P|\psi_{n}\rangle^{2}$ are integers for all $A\subset[n]$. This provides an easily verified necessary condition for a state family to have $\mathrm{SRM}_{0}$. For example, consider the logical $|\overline{T0}\rangle$ in toric code, and take $A$ be a handle which is the support of a logical $\overline{Z}_{1}$. Then we have $\sum_{P:\operatorname{supp}(P)=A}\langle\overline{T0}|P|\overline{T0}\rangle^{2}=1/2$, which rules out $|\overline{T0}\rangle$ from having $\mathrm{SRM}_{0}$ and suggests that it has LRM.
VII Concluding remarks
Motivated by the intriguing interplay between entanglement and nonstabilizerness, as well as the critical role of long-range entanglement in the study of many-body physics, we explore the concept of LRM. We rigorously demonstrate its presence in topological codes, topological orders, and long-range nonstabilizer correlations from various perspectives. Our study provides a systematic theory of the properties of LRM that builds on and generates valuable insights for wide-ranging areas in quantum information, complexity, and many-body physics.
Various intriguing open questions remain. The relationship between LRM and nonlocal magic merits further study, and we provide a preliminary discussion in Appendix F. The study of local state transformations will further contribute to proving LRM from correlations, utilizing more information from the two-point reduced density matrix beyond merely the diagonal elements. From the perspective of LRM itself, valuable future tasks include e.g. developing quantitative measures for LRM, bounding the minimal circuit depth required to erase LRM in different systems, and exploring LRM in mixed states. Also, the question of learning SRM states with polynomial computational complexity Huang et al. (2024); Landau and Liu (2024) warrants further investigation. Other directions that we consider particularly worth further exploring include: a more complete understanding of LRM, stabilizer models and topological order in general qudit systems; relation with the measurement-assisted complexity of topological orders Tantivasadakarn et al. (2023); possible connections the theory of approximate QEC codes (e.g. considering the similar $1/n$ critical error scaling that also emerge in fundamental ways Faist et al. (2020); Liu and Zhou (2023); Yi et al. (2024)); and potential implications to computational notions of phases of matter. Further enriching the theory of LRM from these perspectives is expected to yield deeper insights into the diverse fields connected by LRM and the interactions between them.
Acknowledgements.
We thank Tyler Ellison, Zimu Li, Yuguo Shao, Jinmin Yi and Zishuo Zhao for valuable discussions and feedback. Z.-W.L. is supported in part by a startup funding from YMSC, Tsinghua University, and NSFC under Grant No. 12475023. While finalizing this paper, we noticed the release of an independent work Korbany et al. which considers similar notions of LRM but adopts different perspectives and focuses on different types of systems (1D matrix product states).
Appendix A LRM from limitations of logical gates for topological stabilizer codes
In this section, we present a rigorous proof of the following theorem, which generalizes the toric code case discussed in the main text. We define two edges on a lattice as adjacent if they share a common vertex. We define the distance between two edges to be the length of the shortest path along the lattice connecting the two edges. For a set $E$ of edges on a lattice, the diameter $\operatorname{Diam}(E)$ is defined as the smallest integer $\delta$ such that every pair of edges in $E$ is within distance $\delta$ of each other. For an operator $O$ acting on qubits associated with the edges, we also denote $\operatorname{Diam}(O):=\operatorname{Diam}(\operatorname{supp}(O))$, where $\operatorname{supp}(O)$ denotes the support of $O$.
Theorem 7.
For a $D$-dimensional ($D\geq 2$) TSC family and a $k$-qubit state $|\phi\rangle$, if $|\phi\rangle\notin\big{\{}V^{\dagger}|0^{k}\rangle\mid V\in\mathcal{C}_{k}^{(D)}\big{\}}$, where $k$ is the number of logical qubits, then the corresponding logical state $|\overline{\phi}\rangle$ has LRM.
Proof.
Suppose, for the sake of contradiction, that $|\overline{\phi}\rangle$ does not have LRM. Then, for a sub-family of this TSC with lattice size $\{L_{q}\}_{q=1}^{\infty}\subset\mathbb{Z}^{+}$ and code space projectors $\{\Pi^{(L_{q})}\}$, there exists a family of constant-depth local circuits $\{U^{(L_{q})}\}$ acting on the physical qubits such that $U^{(L_{q})}|\overline{\phi}^{(L_{q})}\rangle=|S^{(L_{q})}\rangle$, where $\{|S^{(L_{q})}\rangle\}$ are stabilizer states. Suppose this family of codes have code parameters $[\![n^{(L_{q})},k,d^{(L_{q})}]\!]$. Suppose the light-cone size of $\{U^{(L_{q})}\}$ are upper bounded by a constant $\eta$. For brevity, we omit the superscript $(L_{q})$ and implicitly consider these families in the following proof.
We claim that $U\Pi U^{\dagger}$ is also a TSC. By Lemma 8 below, we know that there exist independent Pauli strings $\{P_{1},\cdots,P_{n-k}\}$, each with diameter $\mathcal{O}(1)$, that fix $|S\rangle$. Suppose $|a\rangle$ is an arbitrary code state in $U\Pi U^{\dagger}$. Since the reduced density matrices of $|a\rangle$ and $|S\rangle$ on $\operatorname{supp}(P_{i})$ are the same, we have
$\langle a|P_{i}|a\rangle=\operatorname{Tr}(|a\rangle\langle a|P_{i})=\operatorname{Tr}\left(\operatorname{Tr}_{\overline{\operatorname{supp}(P_{i})}}(|a\rangle\langle a|)P_{i}|_{\operatorname{supp}(P_{i})}\right)$ (2)
$=$ $\operatorname{Tr}\left(\operatorname{Tr}_{\overline{\operatorname{supp}(P_{i})}}(|S\rangle\langle S|)P_{i}|_{\operatorname{supp}(P_{i})}\right)=\operatorname{Tr}(|S\rangle\langle S|P_{i})=\langle S|P_{i}|S\rangle=1,$ (3)
where $P_{i}|_{\operatorname{supp}(P_{i})}$ denotes the restriction of $P_{i}$ to its support. Therefore, $P_{i}|a\rangle=|a\rangle$, meaning that these $n-k$ local Pauli strings also fix other code states in $U\Pi U^{\dagger}$. The deformed code $U\Pi U^{\dagger}$ is a stabilizer code with $n-k$ local Pauli stabilizers. Additionally, $U\Pi U^{\dagger}$ has distance $\gtrsim d/\eta$, which can be made arbitrarily large by increasing lattice size $L_{q}$. Therefore, we have proved the claim that $U\Pi U^{\dagger}$ is a TSC.
Denote $\Pi=J_{1}J_{1}^{\dagger}$ and $U\Pi U^{\dagger}=J_{2}J_{2}^{\dagger}$ as projectors onto the original and deformed code space, where $J_{1}$ and $J_{2}$ are Clifford encoding isometries. We have $|\overline{\phi}\rangle=J_{1}|\phi\rangle$. The set of all pure stabilizer states in $U\Pi U^{\dagger}$ is given by $\{J_{2}W|0^{k}\rangle\mid W\in\mathcal{C}_{k}^{(2)}\}$. Since both $\Pi$ and $U\Pi U^{\dagger}$ are $D$-dimensional TSCs, by Bravyi–König (Lemma 2), the induced logical operation $J_{2}^{\dagger}UJ_{1}\in\mathcal{C}_{k}^{(D)}$. Moreover, from Zeng et al. (2008), we know that $\{V^{\dagger}\mid V\in\mathcal{C}_{k}^{(D)}\}=\{V^{\dagger}W\mid V\in\mathcal{C}_{k}^{(D)},W\in\mathcal{C}_{k}^{(2)}\}$, thus
$|\phi\rangle\notin\big{\{}V^{\dagger}|0^{k}\rangle\mid V\in\mathcal{C}_{k}^{(D)}\big{\}}=\big{\{}V^{\dagger}W|0^{k}\rangle\mid V\in\mathcal{C}_{k}^{(D)},W\in\mathcal{C}_{k}^{(2)}\big{\}}.$ (4)
Therefore, $(J_{2}^{\dagger}UJ_{1})|\phi\rangle\notin\{W|0^{k}\rangle\mid W\in\mathcal{C}_{k}^{(2)}\}$, which implies that $U|\overline{\phi}\rangle=UJ_{1}|\phi\rangle$ remains a nonstabilizer state. This leads to a contradiction with $U|\overline{\phi}\rangle=|S\rangle$. ∎
Lemma 8.
Suppose $|S\rangle$ is a stabilizer state, and $U$ is a constant-depth local circuit such that that $U^{\dagger}|S\rangle$ lies in the code space of a TSC with code projector $\Pi$ and code parameters $[\![n,k,d]\!]$, and local Pauli generators $\{g_{1},\cdots,g_{n-k}\}$. Denote by $G$ the stabilizer group of $|S\rangle$, and let $C$ be a constant that upper bounds the diameter of deformed local generators $\{Ug_{1}U^{\dagger},\cdots,Ug_{n-k}U^{\dagger}\}$. Then there exist $n-k$ independent Pauli strings $\{P_{1},\cdots,P_{n-k}\}\subset G$ such that $\operatorname{Diam}(P_{i})\leq C$ for $i=1,\cdots,n-k$.
Proof.
Denote the set of states that are $C$-locally indistinguishable from $|S\rangle$ as
:=$\mathcal{B}:=\Big{\{}|\varphi\rangle\Big{|}\operatorname{Tr}_{\overline{J}}|\varphi\rangle\langle\varphi|=\operatorname{Tr}_{\overline{J}}|S\rangle\langle S|,\forall J\subset[n]\text{ with }\operatorname{Diam}(J)\leq C\Big{\}}.$ (5)
For an arbitrary $|\psi\rangle$ satisfying $Ug_{i}U^{\dagger}|\psi\rangle=|\psi\rangle,i=1,\cdots,n-k$, $U^{\dagger}|\psi\rangle$ is a logical state of $\Pi$, and thus locally indistinguishable from $U^{\dagger}|S\rangle$. Since $U$ is constant-depth local, $|\psi\rangle$ is also locally indistinguishable from $|S\rangle$, meaning that $|\psi\rangle\in\mathcal{B}$. Also, all $|\psi\rangle$ in $\mathcal{B}$ are fixed by all $Ug_{i}U^{\dagger},i=1,\cdots,n-k$. Therefore, we have =$\mathcal{B}=\{|\psi\rangle\mid Ug_{i}U^{\dagger}|\psi\rangle=|\psi\rangle,i=1,\cdots,n-k\}$ and $\dim(\operatorname{span}(\mathcal{B}))=2^{k}$.
Let $r$ be the maximum number of independent Pauli strings in $\{h\in G\mid\operatorname{Diam}(h)\leq C\}$. Pick such an independent set $\{P_{1},\cdots,P_{r}\}\subset\{h\in G\mid\operatorname{Diam}(h)\leq C\}$. We must have
$\{h\in G\mid\operatorname{Diam}(h)\leq C\}\subset\langle P_{1},\cdots,P_{r}\rangle,$ (6)
because otherwise would be able to find $r+1$ independent Pauli strings in $\{h\in G\mid\operatorname{Diam}(h)\leq C\}$.
We take $\{P_{r+1},\cdots,P_{n}\}\subset G$ such that $\{P_{1},\cdots,P_{n}\}$ are independent, which implies that $\langle P_{1},\cdots,P_{n}\rangle=G$. For $j=r+1,\cdots,n$, since $P_{j}\notin\langle P_{1},\cdots,P_{r}\rangle$, we have $\operatorname{Diam}(P_{j})>C$. For $\mathbf{b}=(b_{r+1},\cdots,b_{n})\in\{0,1\}^{n-r}$, consider the states
$|\psi_{\mathbf{b}}\rangle\langle\psi_{\mathbf{b}}|$ $:=\prod_{i=1}^{r}\frac{\mathbb{I}+P_{i}}{2}\prod_{j=r+1}^{n}\frac{\mathbb{I}+(-1)^{b_{j}}P_{j}}{2}$ (7)
$=\frac{1}{2^{n}}\sum_{\mathbf{x}\in\{0,1\}^{n}}\prod_{i=1}^{r}P_{i}^{x_{i}}\prod_{j=r+1}^{r}((-1)^{b_{j}}P_{j})^{x_{j}},$ (8)
where $\mathbf{x}=(x_{1},\cdots,x_{n})$. These states are orthogonal to each other. In particular, we have $|S\rangle=|\psi_{\mathbf{0}}\rangle$.
Consider an arbitrary subset $J\subset[n]$ with $\operatorname{Diam}(J)\leq C$. For $\mathbf{x}$ such that $(x_{r+1},\cdots,x_{n})\neq 0^{n-r}$, we must have
$\operatorname{Diam}\left(\prod_{i=1}^{r}P_{i}^{x_{i}}\prod_{j=r+1}^{r}((-1)^{b_{j}}P_{j})^{x_{j}}\right)=\operatorname{Diam}\left(\prod_{i=1}^{r}P_{i}^{x_{i}}\prod_{j=r+1}^{r}P_{j}^{x_{j}}\right)>C,$ (9)
otherwise $\prod_{i=1}^{r}P_{i}^{x_{i}}\prod_{j=r+1}^{r}P_{j}^{x_{j}}\in\langle P_{1},\cdots,P_{r}\rangle$, which would contradict the independence $\{P_{1},\cdots,P_{n}\}$. Therefore, we have
$\operatorname{Tr}_{\overline{J}}\left(\prod_{i=1}^{r}P_{i}^{x_{i}}\prod_{j=r+1}^{r}((-1)^{b_{j}}P_{j})^{x_{j}}\right)=0.$ (10)
Denote $\mathbf{y}=(y_{1},\cdots,y_{r})$, and we get
$\operatorname{Tr}_{\overline{J}}|\psi_{\mathbf{b}}\rangle\langle\psi_{\mathbf{b}}|=\frac{1}{2^{n}}\operatorname{Tr}_{\overline{J}}\left(\sum_{\mathbf{y}\in\{0,1\}^{r}}\prod_{i=1}^{r}P_{i}^{y_{i}}\right)=\operatorname{Tr}_{\overline{J}}|S\rangle\langle S|,$ (11)
which holds for all $\mathbf{b}\in\{0,1\}^{n-r}$. Therefore, we have $\{|\psi_{\mathbf{b}}\rangle\mid\mathbf{b}\in\{0,1\}^{n-r}\}\subset\mathcal{B}$, which means $2^{k}=\dim(\operatorname{span}(\mathcal{B}))\geq 2^{n-r}$. That is, $r\geq n-k$. Actually, we have $r=n-k$ because $\mathcal{B}\subset\{|\psi\rangle\mid P_{i}|\psi\rangle=|\psi\rangle,i=1,\cdots,r\}$. ∎
Appendix B LRM of topological orders
Theorem 9.
A topological order cannot be realized by any TSC, if and only if for any local Hamiltonian realization of this topological order, all ground states exhibit LRM.
Proof.
($\Leftarrow$) Since for any local Hamiltonian realization of this topological order, there are no stabilizer states in its ground states, we know that this topological order cannot be realized by any TSC.
($\Rightarrow$) Suppose there exists a local Hamiltonian realization of this topological order, and a family of states in the ground space exhibiting SRM. Then, by Proposition 10, we know that a TSC family realizes this topological order, which leads to a contradiction. ∎
Proposition 10.
Suppose a topological order is realized by a family of local Hamiltonians $\{H^{(n)}=\sum_{i\in I^{(n)}}H_{i}^{(n)}\}$ acting on $n$ qubits located on a $D$-dimensional manifold $(D\geq 2)$. If there exists a family of ground states $\{|\phi^{(n)}\rangle\}$ of $\{H^{(n)}\}$ and a family of constant-depth local circuits $\{U^{(n)}\}$, such that $U^{(n)}|\phi^{(n)}\rangle=|S^{(n)}\rangle$ for some stabilizer states $|S^{(n)}\rangle$, then the ground space of $\{U^{(n)}H^{(n)}{U^{(n)}}^{\dagger}\}$ is a family of TSC.
Proof.
For clarity, we omit the superscript “$(n)$” in the proof. By the TQO condition, the states in the ground space of $H$ are locally indistinguishable.
Denote $C$ as a constant independent of $n$ that upper bounds the diameters of deformed local terms $\{UH_{i}U^{\dagger}\}$. Let $\mathcal{B}$ be the set of states that are $C$-locally indistinguishable from $|S\rangle$, defined as
:=$\mathcal{B}:=\Big{\{}|\psi\rangle\Big{|}\operatorname{Tr}_{\overline{J}}|\psi\rangle\langle\psi|=\operatorname{Tr}_{\overline{J}}|S\rangle\langle S|,\forall J\subset[n]\text{ with }\operatorname{Diam}(J)\leq C\Big{\}}.$ (12)
We show that $\mathcal{B}$ is exactly the set of pure ground states of $UHU^{\dagger}$. On one hand, every pure ground state of $UHU^{\dagger}$ can be written as $U|\phi^{\prime}\rangle$, where $|\phi^{\prime}\rangle$ is a ground state of $H$. Since $U$ is constant-depth local, the reduced density matrices of $U|\phi^{\prime}\rangle$ and $|S\rangle=U|\phi\rangle$ are the same locally. Therefore, $U|\phi^{\prime}\rangle\in\mathcal{B}$. On the other hand, the ground state energies of $UHU^{\dagger}$ and $H$ are the same. For any $|b\rangle\in\mathcal{B}$, we have
$\langle b|UH_{i}U^{\dagger}|b\rangle=\langle S|UH_{i}U^{\dagger}|S\rangle=\langle\phi|H_{i}|\phi\rangle$ (13)
for all local terms $H_{i}$, thus $|b\rangle$ is a ground state of $UHU^{\dagger}=\sum_{i}UH_{i}U^{\dagger}$.
Let $r$ be the maximum number of independent Pauli strings in $\{h\in G\mid\operatorname{Diam}(h)\leq C\}$, where $G$ is the stabilizer group of $|S\rangle$. Pick such an independent set $\{P_{1},\cdots,P_{r}\}\subset\{h\in G\mid\operatorname{Diam}(h)\leq C\}$, we must have
$\{h\in G\mid\operatorname{Diam}(h)\leq C\}\subset\langle P_{1},\cdots,P_{r}\rangle,$ (14)
since otherwise we can find $r+1$ independent Pauli strings in $\{h\in G\mid\operatorname{Diam}(h)\leq C\}$. Let
$\mathcal{A}:=\{|\psi\rangle\mid P_{i}|\psi\rangle=|\psi\rangle,i=1,\cdots,r\}.$ (15)
Now we show that $\mathcal{B}=\mathcal{A}$. For $|b\rangle\in\mathcal{B}$, we have $\langle b|P_{i}|b\rangle=\langle S|P_{i}|S\rangle=1$, that is, $P_{i}|b\rangle=|b\rangle$, thus $\mathcal{B}\subset\mathcal{A}$. We take $\{P_{r+1},\cdots,P_{n}\}\subset G$ such that $\{P_{1},\cdots,P_{n}\}$ are independent, and $\langle P_{1},\cdots,P_{n}\rangle=G$. For $j=r+1,\cdots,n$, since $P_{j}\notin\langle P_{1},\cdots,P_{r}\rangle$, we have $\operatorname{Diam}(P_{j})>C$. For $\mathbf{b}=(b_{r+1},\cdots,b_{n})\in\{0,1\}^{n-r}$, consider the states
$|\psi_{\mathbf{b}}\rangle\langle\psi_{\mathbf{b}}|$ $:=\prod_{i=1}^{r}\frac{\mathbb{I}+P_{i}}{2}\prod_{j=r+1}^{n}\frac{\mathbb{I}+(-1)^{b_{j}}P_{j}}{2}.$ (16)
They are orthogonal to each other. In particular, we have $|S\rangle=|\psi_{\mathbf{0}}\rangle$. Following the same procedure as in the proof of Lemma 8, we obtain $\{|\psi_{\mathbf{b}}\rangle\mid\mathbf{b}\in\{0,1\}^{n-r}\}\subset\mathcal{B}$, which means $\dim(\operatorname{span}(\mathcal{B}))\geq 2^{n-r}$. Note that $\dim(\operatorname{span}(\mathcal{A}))=2^{n-r}$, and $\operatorname{span}(\mathcal{B})\subset\operatorname{span}(\mathcal{A})$. We deduce that $\operatorname{span}(\mathcal{B})=\operatorname{span}(\mathcal{A})$. Since $\mathcal{B}$ and $\mathcal{A}$ are the sets of unit vectors in $\operatorname{span}(\mathcal{B})$ and $\operatorname{span}(\mathcal{A})$ respectively, we conclude that $\mathcal{B}=\mathcal{A}$.
To conclude, we identify a set of independent Pauli strings $\{P_{1},\cdots,P_{r}\}$ (with $\langle P_{1},\cdots,P_{r}\rangle$ not containing $-\mathbb{I}$) such that each $P_{i}$ has diameter $\leq C$, and the set of pure ground states of $UHU^{\dagger}$ are exactly $\{|\psi\rangle\mid P_{i}|\psi\rangle=|\psi\rangle,i=1,\cdots,r\}$, which forms a TSC. Therefore, we know that the ground space of $UHU^{\dagger}$ (and also $H$) must have degeneracy $2^{k}$ for some integer $k\geq 0$, and we have $r=n-k$. ∎
Appendix C Strong LRM of topological orders
Lemma 11.
Suppose $H$ is a Hamiltonian acting on a finite-dimensional Hilbert space $\mathcal{H}$, with eigenvalues $E_{0}<E_{1}<E_{2}<\cdots$. Let $D_{0}$ be the degeneracy of the ground-state space with energy $E_{0}$. For any subspace $\mathcal{S}\subset\mathcal{H}$, if every state $|\phi\rangle\in\mathcal{S}$ satisfies $\langle\phi|H|\phi\rangle<E_{1}$, then the dimension of $\mathcal{S}$ satisfies $\dim(\mathcal{S})\leq D_{0}$.
Proof.
Assume for contradiction that $\dim(\mathcal{S})>D_{0}$. Let $\{|\phi_{1}\rangle,\cdots,|\phi_{\dim(\mathcal{S})}\rangle\}$ be an orthonormal basis of $\mathcal{S}$. Let $\Pi_{0}$ denote the projector onto the ground space of $H$. Since $\{\Pi_{0}|\phi_{j}\rangle\}_{j=1}^{\dim(\mathcal{S})}$ are $\dim(\mathcal{S})$ vectors in a space of dimension $D_{0}$, they must be linearly dependent. Therefore, there exist scalars $a_{1},\cdots,a_{\dim(\mathcal{S})}\in\mathbb{C}$, satisfying $\left|a_{1}\right|^{2}+\cdots+\left|a_{\dim(\mathcal{S})}\right|^{2}=1$, such that
$a_{1}\Pi_{0}|\phi_{1}\rangle+\cdots+a_{\dim(\mathcal{S})}\Pi_{0}|\phi_{\dim(\mathcal{S})}\rangle=0.$ (17)
Define $|\psi\rangle:=a_{1}|\phi_{1}\rangle+\cdots+a_{\dim(\mathcal{S})}|\phi_{\dim(\mathcal{S})}\rangle\in\mathcal{S}$, we have $\Pi_{0}|\psi\rangle=0$. Denote $\Pi_{1},\Pi_{2},\cdots$ be the projectors onto the eigenspaces of $E_{1},E_{2},\cdots$, we have
$\langle\psi|H|\psi\rangle=\sum_{i\geq 1}E_{i}\langle\psi|\Pi_{i}|\psi\rangle\geq E_{1}\sum_{i\geq 1}\langle\psi|\Pi_{i}|\psi\rangle=E_{1},$ (18)
a contradiction. We conclude that $\dim(\mathcal{S})\leq D_{0}$. ∎
Theorem 12.
Suppose a topological order is realized by a family of local Hamiltonians $\{H^{(n)}=\sum_{i\in I^{(n)}}H_{i}^{(n)}\}$ acting on $n$ qubits located on a $D$-dimensional manifold $(D\geq 2)$. Suppose $|I^{(n)}|=\Theta(n)$, $\|H_{i}^{(n)}\|_{\infty}\leq 1$. If the dimension of the ground space of $\{H^{(n)}\}$ is not a power of 2, then any family of ground states $\big{\{}|\phi_{n}\rangle\big{\}}$ of $\{H^{(n)}\}$ has strong LRM.
Proof.
Assume for contradiction that there exists a subsequence $T\subset\mathbb{Z}^{+}$, a family of ground states $\{|\phi_{n}\rangle\}_{n\in T}$, a family of constant-depth local circuits $\{U_{n}\}_{n\in T}$ and a family of stabilizer states $\{|S_{n}\rangle\}_{n\in T}$, such that for $n\in T$,
$\big{\|}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=o(1/n).$ (19)
Denote $C$ as a constant independent of $n$ that upper bounds the diameters of all local terms $H_{i}^{(n)}$ and all deformed local terms $UH_{i}^{(n)}U^{\dagger}$. Define $\mathcal{B}_{n}$ as the set of states that are $C$-locally indistinguishable from $U_{n}|\phi_{n}\rangle$,
$\mathcal{B}_{n}:=$$\Big{\{}|\psi\rangle\Big{|}\operatorname{Tr}_{\overline{J}}|\psi\rangle\langle\psi|=\operatorname{Tr}_{\overline{J}}(U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}),\forall J\subset[n]\text{ with }\operatorname{Diam}(J)\leq C\Big{\}}$ (20)
[…]
where the second equality follows from the argument in the proof of Proposition 10.
Let $r_{n}$ be the maximum number of independent Pauli strings in $\{h\in G_{n}\mid\operatorname{Diam}(h)\leq C\}$, where $G_{n}$ is the stabilizer group of $|S\rangle$. Pick such an independent set $\{P_{1}^{(n)},\cdots,P_{r_{n}}^{(n)}\}\subset\{h\in G_{n}\mid\operatorname{Diam}(h)\leq C\}$. Let
$\mathcal{A}_{n}:=$$\Big{\{}\text{pure ground states of }\widetilde{H}^{(n)}:=-\sum_{j=1}^{r_{n}}P_{j}^{(n)}\Big{\}}$ (22)
$=$ $\big{\{}|\psi\rangle\mid P_{j}^{(n)}|\psi\rangle=|\psi\rangle,j=1,\cdots,r_{n}\big{\}}$ (23)
$=$ $\Big{\{}|\psi\rangle\Big{|}\operatorname{Tr}_{\overline{J}}|\psi\rangle\langle\psi|=\operatorname{Tr}_{\overline{J}}|S_{n}\rangle\langle S_{n}|,\forall J\subset[n]\text{ with }\operatorname{Diam}(J)\leq C\Big{\}},$ (24)
where the third equality is given by the proof of Proposition 10.
We claim that $\dim(\operatorname{span}(\mathcal{A}_{n}))=\dim(\operatorname{span}(\mathcal{B}_{n}))$ when $n\in T$ is sufficiently large. For arbitrary $|a\rangle\in\mathcal{A}_{n}$,
$\left|\langle a|U_{n}H^{(n)}U_{n}^{\dagger}|a\rangle-\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle\right|=$ $\left|\langle S_{n}|U_{n}H^{(n)}U_{n}^{\dagger}|S_{n}\rangle-\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle\right|$ (25)
$=$ $\left|\operatorname{Tr}\big{[}H^{(n)}\big{(}U_{n}^{\dagger}|S_{n}\rangle\langle S_{n}|U_{n}-|\phi_{n}\rangle\langle\phi_{n}|\big{)}\big{]}\right|$ (26)
$\leq$ $\sum_{i\in I^{(n)}}\left|\operatorname{Tr}\big{[}H_{i}^{(n)}\big{(}U_{n}^{\dagger}|S_{n}\rangle\langle S_{n}|U_{n}-|\phi_{n}\rangle\langle\phi_{n}|\big{)}\big{]}\right|$ (27)
$\leq$ $\sum_{i\in I^{(n)}}\big{\|}H_{i}^{(n)}\big{\|}_{\infty}\big{\|}|S_{n}\rangle\langle S_{n}|-U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}\big{\|}_{1}$ (28)
$\leq$ $|I^{(n)}|o(1/n)=o(1).$ (29)
Therefore, for arbitrary $|a\rangle\in\mathcal{A}_{n}$ we have
$\langle a|U_{n}H^{(n)}U_{n}^{\dagger}|a\rangle\leq\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle+o(1)=\lambda_{\min}(U_{n}H^{(n)}U_{n}^{\dagger})+o(1).$ (30)
Since $\{U_{n}H^{(n)}U_{n}^{\dagger}\}$ is topological ordered and thus gapped, by Lemma 11 we know $\dim(\operatorname{span}(\mathcal{A}_{n}))\leq\dim(\operatorname{span}(\mathcal{B}_{n}))$ when $n\in T$ is sufficiently large. For arbitrary $|b\rangle\in\mathcal{B}_{n}$,
$\left|\langle b|\widetilde{H}^{(n)}|b\rangle-\langle S_{n}|\widetilde{H}^{(n)}|S_{n}\rangle\right|=$ $\left|\langle\phi_{n}|U_{n}^{\dagger}\widetilde{H}^{(n)}U_{n}|\phi_{n}\rangle-\langle S_{n}|\widetilde{H}^{(n)}|S_{n}\rangle\right|$ (31)
$=$ $\left|\operatorname{Tr}\big{[}\widetilde{H}^{(n)}\big{(}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{)}\big{]}\right|$ (32)
$\leq$ $\sum_{j=1}^{r_{n}}\left|\operatorname{Tr}\big{[}P_{j}^{(n)}\big{(}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{)}\big{]}\right|$ (33)
$\leq$ $\sum_{j=1}^{r_{n}}\big{\|}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}$ (34)
$\leq$ $r_{n}o(1/n)=o(1).$ (35)
Therefore, for arbitrary $|b\rangle\in\mathcal{B}_{n}$ we have
$\langle b|\widetilde{H}^{(n)}|b\rangle\leq\langle S_{n}|\widetilde{H}^{(n)}|S_{n}\rangle+o(1)=\lambda_{\min}(\widetilde{H}^{(n)})+o(1).$ (36)
Since $\{\widetilde{H}^{(n)}\}$ is gapped, by Lemma 11 we know $\dim(\operatorname{span}(\mathcal{B}_{n}))\leq\dim(\operatorname{span}(\mathcal{A}_{n}))$ when $n\in T$ is sufficiently large.
Therefore, when $n\in T$ is sufficiently large, we have
$2^{n-r_{n}}=\dim(\operatorname{span}(\mathcal{A}_{n}))=\dim(\operatorname{span}(\mathcal{B}_{n}))=\text{ground space dimension of $\{H^{(n)}\}$}.$ (37)
Since we assume the ground space dimension of $\{H^{(n)}\}$ is not a power of 2, we get a contradiction. This allows us to conclude that for arbitrary ground state family $\big{\{}|\phi_{n}\rangle\big{\}}$ of $\{H^{(n)}\}$, any family of constant-depth local circuits $\{U_{n}\}$ and any family of stabilizer states $\{|S_{n}\rangle\}$, we have
$\big{\|}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=\Omega(1/n),$ (38)
that is, $\big{\{}|\phi_{n}\rangle\big{\}}$ has strong LRM. ∎
Appendix D LRM from correlation
D.1 Bipartite stabilizer mixed states
Given the generating set $\{g_{1},\cdots,g_{r}\}$ for an $n$-qubit stabilizer code, where $r\leq n$ and $g_{1},\cdots,g_{r}$ are independent, mutually commuting Pauli strings such that $\langle g_{1},\cdots,g_{r}\rangle$ does not contain $-\mathbb{I}$, the corresponding stabilizer mixed state Aaronson and Gottesman (2004) is defined as $\frac{1}{2^{r}}\prod_{i=1}^{r}(\mathbb{I}+g_{i})$, which is the uniform distribution over all states within the code space. Stabilizer mixed states are in one-to-one correspondence with stabilizer codes.
Lemma 13.
A bipartite stabilizer mixed state $\rho$ on $N$ qubits can be obtained by local operations on at most $N$ EPR pairs.
Proof.
By Theorem 1 in Audenaert and Plenio (2005), every bipartite stabilizer mixed state $\rho_{AB}$ on $N=N_{A}+N_{B}$ qubits with $r$ independent generators is locally unitary equivalent to $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes p}\otimes\rho^{\prime}$, where $|\Phi^{+}\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$ is EPR pair shared by $A$ and $B$, and
$\rho^{\prime}=\frac{1}{2^{N-2p}}\prod_{l=2p+1}^{r}(\mathbb{I}+h_{l})$ (39)
is a bipartite separable stabilizer mixed state acting on $(N_{A}-p)+(N_{B}-p)$ qubits, with $h_{l}$ a tensor product of single-qubit $\mathbb{I}$ and $X$, acting on $N-2p$ qubits. The number of EPR pairs $p$ is upper bounded by $p\leq\min(\lfloor r/2\rfloor,N_{A},N_{B})$.
[…]
D.2 Generate correlations by EPR pairs
We take $P=Z$ without loss of generality.
Proposition 14.
For integer $K\geq 1$, $K$ EPR pairs can generate a two-qubit state $\rho$ by local CPTP maps, with $\langle Z\otimes Z\rangle_{\rho}=b$ and $\langle Z\otimes\mathbb{I}\rangle_{\rho}=\langle\mathbb{I}\otimes Z\rangle_{\rho}=c$, if and only if $(b,c)$ lies in the region given by
$1+2c+b\geq 0,$ (40)
$1-2c+b\geq 0,$ (41)
$\frac{1+2c+b}{4}2^{K}\leq l+(\frac{1+c}{2}2^{K}-l)^{2},\text{ for}l=0,1,\cdots,2^{K}-1.$ (42)
See Fig. 2 for the region given by constraints Eqs. (40)–(42).
Proof.
($\Rightarrow$) Suppose $\rho=\mathcal{E}_{A}\otimes\mathcal{E}_{B}\big{(}|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\big{)}$ for some local CPTP maps $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$. Measure $\rho$ in the computational basis, we get $|00\rangle$ with probability $(1+2c+b)/4$, get $|11\rangle$ with probability $(1-2c+b)/4$, and get $|01\rangle$ or $|10\rangle$ with probability $(1-b)/4$. Therefore, we get a classical correlation Jain et al. (2013)
$\frac{1}{4}\begin{bmatrix}1+2c+b&1-b\\ 1-b&1-2c+b\\ \end{bmatrix}.$ (43)
To ensure (43) a valid correlation, each element should be non-negative. By measuring local POVMs $\{\mathcal{E}_{A}^{\dagger}(|0\rangle\langle 0|),\mathcal{E}_{A}^{\dagger}(|1\rangle\langle 1|)\}$ and $\{\mathcal{E}_{B}^{\dagger}(|0\rangle\langle 0|),\mathcal{E}_{B}^{\dagger}(|1\rangle\langle 1|)\}$ on $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}$, we can get the correlation (43). All $2\times 2$ classical correlations $M$ that can be obtained by $K$ EPR pairs, can be obtained by measuring local POVMs $\{E,\mathbb{I}_{2^{K}}-E\}$ and $\{F,\mathbb{I}_{2^{K}}-F\}$ acting on $K$ qubits on $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}$, and can be written as
$M=$ $\begin{bmatrix}\operatorname{Tr}\left(E\otimes F|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\right)&\operatorname{Tr}\left(E\otimes(\mathbb{I}_{2^{K}}-F)|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\right)\\ \operatorname{Tr}\left((\mathbb{I}_{2^{K}}-E)\otimes F|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\right)&\operatorname{Tr}\left((\mathbb{I}_{2^{K}}-E)\otimes(\mathbb{I}_{2^{K}}-F)|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\right)\end{bmatrix}$ (44)
$=$ $2^{-K}\begin{bmatrix}\operatorname{Tr}\left(E^{\mathrm{T}}F\right)&\operatorname{Tr}\left(E\right)-\operatorname{Tr}\left(E^{\mathrm{T}}F\right)\\ \operatorname{Tr}\left(F\right)-\operatorname{Tr}\left(E^{\mathrm{T}}F\right)&2^{K}-\operatorname{Tr}\left(E\right)-\operatorname{Tr}\left(F\right)+\operatorname{Tr}\left(E^{\mathrm{T}}F\right)\end{bmatrix}.$ (45)
Letting (45) $=$ (43), we get
$\operatorname{Tr}(E)=\frac{1+c}{2}2^{K},$ (46)
$\operatorname{Tr}(F)=\frac{1+c}{2}2^{K},$ (47)
$\operatorname{Tr}(E^{\mathrm{T}}F)=\frac{1+2c+b}{4}2^{K},$ (48)
$0\leq E,F\leq\mathbb{I}_{2^{K}}.$ (49)
Note that $\operatorname{Tr}(E^{\mathrm{T}}F)=\operatorname{Tr}((E^{*})^{\dagger}F)\leq\left\|E^{*}\right\|_{2}\left\|F\right\|_{2}=\left\|E\right\|_{2}\left\|F\right\|_{2}$, where $E^{*}$ is the complex conjugate of $E$. For $l=0,1,\cdots,2^{K}-1$, when $l\leq\frac{1+c}{2}2^{K}\leq l+1$, we have $\left\|E\right\|_{2},\left\|F\right\|_{2}\leq\sqrt{l+(\frac{1+c}{2}2^{K}-l)^{2}}$. Therefore, we obtain
$\operatorname{Tr}(E^{\mathrm{T}}F)\leq l+(\frac{1+c}{2}2^{K}-l)^{2},$ (50)
giving the constraint $\frac{1+2c+b}{4}2^{K}\leq l+(\frac{1+c}{2}2^{K}-l)^{2}$. These $2^{K}$ constraints also imply $1-b\geq 0$.
($\Leftarrow$) Suppose $(b,c)$ lies in the region given by Eqs. (40)–(42). We claim that we can generate the correlation (43), by measuring some POVMs $\{E,\mathbb{I}_{2^{K}}-E\}$ and $\{F,\mathbb{I}_{2^{K}}-F\}$ on $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}$, that is, we can find $E$ and $F$ satisfying the constraints in Eqs. (46)–(49).
For $l=0,1,\cdots,2^{K-1}-1$, when $c$ satisfies $l\leq\frac{1+c}{2}2^{K}\leq l+1$, we have
$0\leq\frac{1+2c+b}{4}2^{K}\leq l+(\frac{1+c}{2}2^{K}-l)^{2}.$ (51)
Since $\operatorname{Tr}(E)=\operatorname{Tr}(F)=\frac{1+c}{2}2^{K}\leq 2^{K-1}$, we can take $E^{*}$ and $F$ such their support are disjoint and get $\operatorname{Tr}(E^{\mathrm{T}}F)=0$. By taking $E=F=\operatorname{diag}(1,\cdots,1,\frac{1+c}{2}2^{K}-l,0,\cdots,0)$ with $l$ 1’s, we get $\operatorname{Tr}(E^{\mathrm{T}}F)=l+(\frac{1+c}{2}2^{K}-l)^{2}$. Moreover, $\operatorname{Tr}(E^{\mathrm{T}}F)$ takes arbitrary values in the region $[0,l+(\frac{1+c}{2}2^{K}-l)^{2}]$ by appropriately choosing $0\leq E,F\leq\mathbb{I}_{2^{K}}$. Therefore, we get $E$ and $F$ satisfying the constraints in Eqs. (46)–(49).
For $l=2^{K-1},1,\cdots,2^{K}-1$, when $c$ satisfies $l\leq\frac{1+c}{2}2^{K}\leq l+1$, we have
$c2^{K}\leq\frac{1+2c+b}{4}2^{K}\leq l+(\frac{1+c}{2}2^{K}-l)^{2}.$ (52)
By taking $E=\operatorname{diag}(1,\cdots,1,c,\cdots,c)$ and $F=\operatorname{diag}(c,\cdots,c,1,\cdots,1)$ with $2^{K-1}$ 1’s and $2^{K-1}$ $c$’s, we get $\operatorname{Tr}(E)=\operatorname{Tr}(F)=\frac{1+c}{2}2^{K}$ and $\operatorname{Tr}(E^{\mathrm{T}}F)=c2^{K}$. By taking $E=F=\operatorname{diag}(1,\cdots,1,\frac{1+c}{2}2^{K}-l,0,\cdots,0)$ with $l$ 1’s, we get $\operatorname{Tr}(E^{\mathrm{T}}F)=l+(\frac{1+c}{2}2^{K}-l)^{2}$. Moreover, $\operatorname{Tr}(E^{\mathrm{T}}F)$ takes arbitrary values in the region $[c2^{K},l+(\frac{1+c}{2}2^{K}-l)^{2}]$ by appropriately choosing $0\leq E,F\leq\mathbb{I}_{2^{K}}$. Therefore, we get $E$ and $F$ satisfying the constraints in Eqs. (46)–(49).
Given $(b,c)$, measuring the corresponding POVMs $\{E,\mathbb{I}_{2^{K}}-E\}$ and $\{F,\mathbb{I}_{2^{K}}-F\}$ on $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}$, get the correlation (43), and then prepare the diagonal state
$\rho:=\frac{1+2c+b}{4}|00\rangle\langle 00|+\frac{1-b}{4}|01\rangle\langle 01|+\frac{1-b}{4}|10\rangle\langle 10|+\frac{1-2c+b}{4}|00\rangle\langle 00|$ (53)
conditioned on the measurement results. We have $\langle Z\otimes Z\rangle_{\rho}=b$ and $\langle Z\otimes\mathbb{I}\rangle_{\rho}=\langle\mathbb{I}\otimes Z\rangle_{\rho}=c$, and $\rho$ is prepared by local CPTP maps on $K$ EPR pairs. ∎
D.3 $|\mathrm{C}^{n-1}Z\rangle$
The reduced density matrix of $|\mathrm{C}^{n-1}Z\rangle$ Wei and Liu (2024) on two qubits $\{k,l\}\subset[n]$ is
$\operatorname{Tr}_{\overline{\{k,l\}}}|\mathrm{C}^{n-1}Z\rangle\langle\mathrm{C}^{n-1}Z|=\big{(}1-\frac{1}{2^{n-2}}\big{)}|++\rangle\langle++|+\frac{1}{2^{n-2}}|\mathrm{C}Z\rangle\langle\mathrm{C}Z|.$ (54)
We have $\langle\mathrm{C}Z|X\otimes X|\mathrm{C}Z\rangle=0$, thus
$\langle X_{k}X_{l}\rangle=\operatorname{Tr}\left(X\otimes X\operatorname{Tr}_{\overline{\{k,l\}}}|\mathrm{C}^{n-1}Z\rangle\langle\mathrm{C}^{n-1}Z|\right)=1-\frac{1}{2^{n-2}}.$ (55)
On one qubit we have
$\operatorname{Tr}_{\overline{\{k\}}}|\mathrm{C}^{n-1}Z\rangle\langle\mathrm{C}^{n-1}Z|=\big{(}1-\frac{1}{2^{n-1}}\big{)}|+\rangle\langle+|+\frac{1}{2^{n-1}}|-\rangle\langle-|,$ (56)
thus
$\langle X_{k}\rangle=1-\frac{1}{2^{n-1}}-\frac{1}{2^{n-1}}=1-\frac{1}{2^{n-2}}.$ (57)
Appendix E NLTM as a strengthened necessary condition for quantum PCP
Here we present a straightforward argument that the validity of NLTM conjecture is necessary for quantum PCP to hold. That is, as remarked in the main text, NLTM is a stronger version of NLTS that gets closer to quantum PCP since it takes into account more general classical witnesses.
Proposition 15.
Suppose $\mathrm{NP}\neq\mathrm{QMA}$ and the quantum PCP conjecture holds, then the NLTM conjecture holds.
The proof is adapted from Ref. Nirkhe (2022). Various necessary preliminary definitions can be found therein.
Proof.
Assume the contrary, namely, assume that NLTM is false while $\mathrm{NP}\neq\mathrm{QMA}$ and the quantum PCP conjecture holds.
Take an arbitrary QMA decision problem $L$. If the quantum PCP conjecture is true, there is a polynomial-time quantum reduction that maps an input instance $x$ of $L$ to a local Hamiltonian $H^{(n)}=\sum_{i=1}^{m^{(n)}}H_{i}^{(n)}$ acting on $n$ qubits ($H^{(n)}$ depends on $x$, $n$ relies on the size of $x$), with $m^{(n)}=\Theta(n)$ and $\|H_{i}^{(n)}\|_{\infty}\leq 1$. There are two energy thresholds $a^{(n)}$ and $b^{(n)}$ with $b^{(n)}-a^{(n)}\geq\gamma n$ for some constant $\gamma>0$. The promise for $H^{(n)}$ is:
  1. 1.
    YES instance: If $x\in L$, then $\lambda_{\min}(H^{(n)})\leq a^{(n)}$;
  2. 2.
    NO instance: If $x\notin L$, then $\lambda_{\min}(H^{(n)})\geq b^{(n)}$.
By false-NLTM assumption, we now show that $L\in\mathrm{NP}$.
Consider a family of YES instance $\{x\}$ and the corresponding $\{H^{(n)}\}$. For constant $\gamma>0$, by false-NLTM, for this family of Hamiltonian $\{H^{(n)}\}$, there exists a family of states $\{|\phi_{n}\rangle\}$ with
$\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle<\lambda_{\min}(H^{(n)})+\gamma n\leq a^{(n)}+\gamma n=b^{(n)}$ (58)
that has SRM. Suppose they can be written as $|\phi_{n}\rangle=U_{n}|S_{n}\rangle$ for families of constant-depth circuits $\{U_{n}\}$ and stabilizer states $\{|S_{n}\rangle\}$. The classical description of $U_{n}$ and $|S_{n}\rangle$ can serve as the NP witness for YES instance $H^{(n)}$. To see this, first notice that $\langle\phi_{n}|H^{(n)}|\phi^{(n)}\rangle<b^{(n)}$ is a convincing proof that $\lambda_{\min}(H^{(n)})\leq a^{(n)}$, due to the promise. Also, $\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle$ can be computed efficiently classically by the classical description of $U_{n}$ and $|S_{n}\rangle$, because
$\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle=\sum_{i=1}^{m^{(n)}}\langle S_{n}|U_{n}^{\dagger}H_{i}^{(n)}U_{n}|S_{n}\rangle,$ (59)
where $\big{|}\operatorname{supp}\big{(}U_{n}^{\dagger}H_{i}^{(n)}U_{n}\big{)}\big{|}=\mathcal{O}(1)$ implies $\langle S_{n}|U_{n}^{\dagger}H_{i}^{(n)}U_{n}|S_{n}\rangle$ can be computed efficiently by e.g. decompose $U_{n}^{\dagger}H_{i}^{(n)}U_{n}$ into the sum of $\mathcal{O}(1)$ Pauli strings Aaronson and Gottesman (2004).
For a family of NO instance $\{x\}$ and the corresponding $\{H^{(n)}\}$, since $\lambda_{\min}(H^{(n)})\geq b^{(n)}$, arbitrary NP witnesses will not be accepted.
Therefore we conclude that $L\in\mathrm{NP}$. Since $L\in\mathrm{QMA}$ is taken arbitrarily, we have $\mathrm{QMA}\subset\mathrm{NP}$, contrary to the assumption that $\mathrm{NP}\neq\mathrm{QMA}$. ∎
Appendix F LRM and nonlocal magic
Given a partition of the whole system into several subsystems, we say a nonstabilizer state has completely nonlocal magic if its local reduced density matrices on subsystems have no magic.
We show that LRM is stronger than completely nonlocal magic.
Suppose $\{|\psi_{n}\rangle\}$ has LRM, then for arbitrary disjoint local partition $[n]=R_{1}\cup\cdots\cup R_{t_{n}}$, with $\left|R_{i}\right|=\mathcal{O}(1)$, we can use a constant-depth local circuit $U^{(n)}=U_{1}\otimes\cdots\otimes U_{t_{n}}$ to rotate the reduced density matrix on each $R_{i}$ into a matrix that is diagonal in the computational basis, thereby exhibiting no magic on each $R_{i}$. Meanwhile, $U^{(n)}|\psi_{n}\rangle$ remains magical, because $U^{(n)}$ cannot remove all the magic of $|\psi_{n}\rangle$ when $n$ is sufficiently large.
On the other hand, some families of states with SRM can have completely nonlocal magic. The logical $|T\rangle^{\otimes k}$ state in 3D color codes can be obtained by performing a single layer of $T$ gates on logical $|+\rangle^{\otimes k}$, a stabilizer state, thus having SRM. Note that its magic is completely nonlocal because it exhibits no magic when restricted to regions smaller than the code distance.