### v2

# Long-range nonstabilizerness from quantum codes, orders, and correlations

Abstract
Understanding nonstabilizerness (aka quantum magic) in many-body quantum systems, particularly its interplay with entanglement, represents an important quest in quantum computation and many-body physics. Drawing natural motivations from the study of quantum phases of matter and entanglement, we systematically investigate the notion of long-range magic (LRM), defined as nonstabilizerness that cannot be (approximately) erased by shallow local unitary circuits. In doing so, we prove a robust generalization of the Bravyi–König theorem. By establishing connections to the theory of fault-tolerant logical gates on quantum error-correcting codes, we show that certain families of topological stabilizer code states exhibit LRM. Then, we show that all ground states of topological orders that cannot be realized by topological stabilizer codes, such as Fibonacci topological order, exhibit LRM, which yields a “no lowest-energy trivial magic” result. Building on our considerations of LRM, we discuss the classicality of short-range magic from e.g. preparation and learning perspectives, and put forward a “no low-energy trivial magic” (NLTM) conjecture that has key motivation in the quantum PCP context. We also connect correlation functions with LRM, demonstrating certain LRM state families by correlation properties. Most of our concepts and techniques do not rely on geometric locality and can be extended to systems with general connectivity. Our study leverages and sheds new light on the interplay between quantum resources, error correction and fault tolerance, complexity theory, and many-body physics.
I Introduction
A pivotal mission of quantum information science is to understand and characterize the origin of potential advantages of quantum computers over classical ones. As indicated by the Gottesman–Knill theorem Gottesman (1998); Aaronson and Gottesman (2004); Nielsen and Chuang (2011), stabilizer quantum computation can be efficiently simulated by classical methods. Consequently, nonstabilizerness (magic) Bravyi and Kitaev (2005); Veitch et al. (2014) can be understood as a “resource” for quantum computational speedups, with its characterizations and meanings having been actively studied from e.g. resource theory Veitch et al. (2014); Howard and Campbell (2017) and classical simulation Bravyi et al. (2016); Bravyi and Gosset (2016); Bravyi et al. (2019); Seddon et al. (2021); Bu and Koh (2019) perspectives for a long time. Recently, nonstabilizerness in many-body quantum systems has attracted substantial interest from both quantum information and physics communities, with extensive efforts devoted to quantitatively characterizing many-body nonstabilizerness and its interplay with entanglement (see e.g. Refs. Liu and Winter (2022); Leone et al. (2022); Haug and Piroli (2023); Lami and Collura (2023); Tarabunga et al. (2024); Frau et al. (2024); Gu et al. ; Liu and Clark (2025)) and understanding many-body physical phenomena through the lens of nonstabilizerness (see e.g. Refs. Sarkar et al. (2020); White et al. (2021); Liu and Winter (2022); Tarabunga et al. (2023); Gu et al. ; Bejan et al. (2024)).
In the development of condensed matter physics, the concept of long-range entanglement Chen et al. (2010); Wen (2013); Hastings and Wen (2005) has played a crucial role, underpinning intrinsically quantum notions of order and phases of matter that extend beyond the traditional Landau paradigm. Here “long-range” captures the global, topological nature of the feature, requiring it to be preserved under constant (equivalently, $O(1)$ or finite) depth local quantum circuits representing “smooth” deformations that do not alter fundamentally global properties.
Drawing inspiration from this, we adopt a computational perspective and consider long-range magic (LRM) White et al. (2021); Ellison et al. (2021), defined as nonstabilizerness that cannot be erased by constant-depth local circuits, which captures the “topological” type of nonstabilizerness embedded in global structures and robust against local perturbations. Several insights are worth noting. First, it is evident that LRM is a strictly stronger notion than long-range entanglement, suggesting that LRM induces a refined characterization of many-body quantumness with computational underpinnings. Second, from a computational complexity perspective, LRM can be understood as high-complexity nonstabilizerness intrinsically far from efficient classical computation, thus holding natural relevance in quantum complexity theory, particularly to the prominent quantum probabilistically checkable proofs (PCP) conjecture Aharonov et al. . Third, as we will demonstrate, the study of LRM bridges many fields including quantum error correction and fault tolerance, computation, complexity, and many-body physics in new interesting ways, drawing on insights from these fields to inform one another. Given the wide importance of these areas and connections, systematic theories and constructions of LRM would be highly valuable but have so far remained largely elusive.
In this work, we address this gap by developing a rigorous theory of LRM and studying its multiple facets, including construction and significance, from diverse perspectives. Note that establishing LRM essentially entails the challenging problem of proving nontrivial circuit lower bounds for erasing nonstabilizerness. First, we demonstrate examples of LRM rooted in topological codes and topologically ordered systems, highlighting the insight that long-range properties can originate from topology. Building on Bravyi and König’s insight in the fault tolerance context that certain logical operations cannot be implemented by constant-depth circuits on topological stabilizer codes Bravyi and König (2013), we show that they generate robust LRM state families, a basic example being logical nonstabilizer states in the toric code. Along the way, we generalize the Bravyi–König theorem to an approximate form and present a new method for understanding transversal gates, which may find independent applications in quantum error correction. Furthermore, utilizing the local indistinguishability property of the ground states of general topologically ordered systems Bravyi et al. (2006, 2010), we establish general conditions for “LRM phases” in which the entire ground spaces of entire phases of topological orders have LRM. Examples include non-Abelian topological orders based on arguments in the literature that they cannot be realized by stabilizer models Potter and Vasseur (2016); Ellison et al. (2022). We also establish a simple ground space dimension condition which indicates robust LRM in e.g., Fibonacci topological order on certain manifolds. Such LRM phases are remarkable as they naturally evade all stabilizer states although they pervade the Hilbert space, establishing LRM as a new emergent property with physical significance. Then, we bridge LRM with correlation properties and present a general scheme for establishing LRM from correlation functions, which can be used to establish LRM properties of e.g., $\mathrm{C}^{n-1}Z$ states and GHZ-type states. We note that most of our LRM results are robust even against $O(1/n)$ approximation and do not need geometric locality unless specified. After presenting these LRM constructions, we examine in more depth the classical characteristics of systems without LRM from various perspectives including preparation, description, and learning complexities. In particular, building upon our LRM phase results, we formulate a “no low-energy trivial magic” (NLTM) conjecture, which strengthens the recently validated no low-energy trivial state (NLTS) theorem Freedman and Hastings ; Anshu et al. (2023) by unifying the trivial entanglement and nonstabilizerness types of efficient classical proofs, potentially yielding important progress towards quantum PCP.
II Preliminaries
Key to our study is the notion of shallow circuits, i.e., families of local quantum circuits consisting of a constant ($O(1)$ as the system size grows) number of layers of local unitary gates which only interact a finite number of neighboring qubits as determined by certain connectivity/adjacency graphs. Note that we shall stick to constant-depth circuits throughout for the simplicity of technical statements, but generalizations to logarithmic-depth circuits are straightforward.
Definition 1.
A family of states $\{|\psi_{n}\rangle\}$, where $|\psi_{n}\rangle$ is an $n$-qubit state, is said to have long-range magic (LRM), if for any family of constant-depth local circuits $\{U_{n}\}$, $U_{n}|\psi_{n}\rangle$ remains a magic state for sufficiently large $n$. Conversely, we say $\{|\psi_{n}\rangle\}$ has short-range magic (SRM) if there exists a family of constant-depth local circuits $\{U_{n}\}$ such that $\{U_{n}|\psi_{n}\rangle\}$ are stabilizer states. Furthermore, $\{|\psi_{n}\rangle\}$ is said to have strong LRM, if for any family of constant-depth local circuits $\{U_{n}\}$ and any family of stabilizer states $\{|S_{n}\rangle\}$, it holds that $\big{\|}U_{n}|\psi_{n}\rangle\langle\psi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=\Omega(1/n)$.
If $\{|\psi_{n}\rangle\}$ doesn’t have LRM, there exists a subsequence $\{n_{q}\}_{q=1}^{\infty}\subset\mathbb{Z}_{>0}$, such that $\{|\psi_{n_{q}}\rangle\}_{q=1}^{\infty}$ has SRM. If $\{|\psi_{n}\rangle\}$ has LRM, then it must have long-range entanglement, i.e., cannot be transformed into product states by constant-depth local circuits. Sometimes we just say a state $|\psi\rangle$ has LRM, implicitly referring to a family of such states. The existence of LRM does not depend on the choice of local bases. Also note that if $\{|\phi_{n}\rangle\}$ has strong LRM, then $|\phi_{n}\rangle$ must have distance $\Omega(1/n)$ to any SRM states. As will be seen, this $\Omega(1/n)$ tolerance is highly lenient and fundamental.
The Clifford hierarchy for $n$-qubits is defined as follows. The first level consists of the Pauli group, $\mathcal{C}_{n}^{(1)}:=\mathcal{P}_{n}$. For $l\geq 2$, the $l$-th level is recursively defined as $\mathcal{C}_{n}^{(l)}:=\{U\in\operatorname{U}(2^{n})\mid UPU^{\dagger}\in\mathcal{C}_{n}^{(l-1)},\forall P\in\mathcal{P}_{n}\}$, where $\operatorname{U}(2^{n})$ is the unitary group.
A $D$-dimensional topological stabilizer code (TSC) family is a family of codes that encode a fixed number of logical qubits. The physical qubits are arranged on a lattice embedded in a $D$-dimensional manifold, where the Pauli stabilizers have constant-diameter support. The code distance can be made arbitrarily large by increasing the lattice size.
Figure 1: (a) A set of $n$ stabilizer generators of the logical state $|\overline{T0}\rangle$ in toric code, one of them is non-Pauli. We define $|T\rangle:=\operatorname{cos}(\pi/8)|0\rangle+\operatorname{sin}(\pi/8)|1\rangle$, the $+1$ eigenstate of $\frac{1}{\sqrt{2}}(X+Z)$. The Pauli generators (blue) consist of $n-2$ local ones and a global $\overline{Z}_{2}$, which is a tensor product of Pauli $Z$’s around the blue handle. The orange one is a global non-Pauli generator $\frac{1}{\sqrt{2}}(\overline{X}_{1}+\overline{Z}_{1})$, which introduces global $\frac{1}{\sqrt{2}}$’s into the Pauli spectrum of $|\overline{T0}\rangle$, producing “topological nonstabilizerness”. Tracing out the entire region outside an arbitrary gray “cross” will not decrease the amount of nonstabilizerness contained in $|\overline{T0}\rangle$. (b) The bipartite reduced density matrix of a state with SRM on distant subsystems $A$ and $B$ is only correlated by a stabilizer state $|S\rangle$, as the light cones of $A$ and $B$ do not intersect.
III Long-range magic from topological codes
Our first observation is that LRM state families can arise naturally from TSCs due to their topological features. In particular, the topology and locality constraints underpin a nice connection between the LRM property and limitations on fault-tolerant logical gates of topological codes as indicated by the Bravyi–König theorem Bravyi and König (2013). Building on this insight, we show:
Theorem 1.
For a $D$-dimensional ($D\geq 2$) TSC family, if $|\phi\rangle\notin\big{\{}V^{\dagger}|0^{k}\rangle\mid V\in\mathcal{C}_{k}^{(D)}\big{\}}$ (where $k$ is the number of logical qubits), then the corresponding logical state family $|\overline{\phi}\rangle$ has strong LRM.
We illustrate the essential ideas of the proof of this theorem through the following concrete example of toric code. The fully general proof can be found in Appendix A.
Example 1.
The logical nonstabilizer state $|\overline{T0}\rangle$ in toric code has LRM.
As a fundamental topological feature, all code states in the toric code Kitaev (2003a) on a $L\times L$ lattice ($n=2L^{2}$ physical qubits, $2$ logical qubits) has long-range entanglement, requiring local circuits of depth $\Theta(L)$ to prepare Dennis et al. (2002); Bravyi et al. (2006) from a product state. Notice that while $|\overline{T0}\rangle$ has circuit complexity $\Theta(L)$, we cannot immediately rule out the possibility that it is mapped to a stabilizer state, thus losing all magic, by a constant-depth local circuit. The “topological nonstabilizerness” of $|\overline{T0}\rangle$ can be visualized by its Pauli spectrum. Recall that for any stabilizer state $|\phi\rangle$ and Pauli string $P$ with $+1$ phase, we have $\langle\phi|P|\phi\rangle=\pm 1$ when ±$\pm P$ belongs to the stabilizer group of $|\phi\rangle$, and $\langle\phi|P|\phi\rangle=0$ otherwise. Due to the global non-Pauli stabilizer generator $\frac{1}{\sqrt{2}}(\overline{X}_{1}+\overline{Z}_{1})$ for $|\overline{T0}\rangle$ (see Fig. 1(a)), we find $\langle\overline{T0}|P|\overline{T0}\rangle=\pm\frac{1}{\sqrt{2}}$ for exponentially many $P$ with support wrapping around the torus. This suggests that a constant-depth local circuit on the torus cannot organize its Pauli spectrum to be $-1,0,1$ to match that of a stabilizer state.
We prove $|\overline{T0}\rangle$ in toric code has LRM by contradiction. Suppose on the contrary that $|\overline{T0}\rangle$ does not have LRM. Then, for any lattice size $L_{q}$ in a subsequence of sizes $\{L_{q}\}_{q=1}^{\infty}\subset\mathbb{Z}_{>0}$, there exists a constant-depth local circuit $U$ such that $U|\overline{T0}\rangle=|S\rangle$, where $|S\rangle$ is a stabilizer state. Denote $C$ as a constant that is independent of $L_{q}$, which upper bounds the diameters of the $n-2$ deformed local generators $\{UA_{v}U^{\dagger}\}$, $\{UB_{p}U^{\dagger}\}$. There are $3$ “brother” states of $|S\rangle=U|\overline{T0}\rangle$, namely $U\overline{Y}_{1}|\overline{T0}\rangle$, $U\overline{Y}_{2}|\overline{T0}\rangle$, and $U\overline{Y}_{1}\overline{Y}_{2}|\overline{T0}\rangle$. These brother states and $|S\rangle$ are mutually orthogonal, and their reduced density matrices are identical when restricted to systems of diameter $\leq C$. At most $3$ such brother states of $|S\rangle$ exist since these states are all invariant under the action of $n-2$ deformed generators $\{UA_{v}U^{\dagger}\}$, $\{UB_{p}U^{\dagger}\}$.
Now we show that the deformed code space $U\Pi U^{\dagger}$ is also a TSC, where $\Pi$ denotes the projector onto the original toric code. Note that the deformed generators $\{UA_{v}U^{\dagger}\}$, $\{UB_{p}U^{\dagger}\}$ may not be Pauli strings. We find the Pauli stabilizers of $U\Pi U^{\dagger}$ from the stabilizer group $G$ of $|S\rangle$. We claim that there must exist $n-2$ independent Pauli strings in $G$ with diameters $\leq C$, where “independent” means they cannot multiply to $\mathbb{I}_{2^{n}}$. Suppose for example that, there exists and exists at most $n-3$ such local Pauli strings, then we can construct $7$ brother states for $|S\rangle$ by finding $3$ Pauli strings $Q_{1},Q_{2},Q_{3}$ in $G$, satisfying that $Q_{1},Q_{2},Q_{3}$ together with the $n-3$ local Pauli strings generate $G$. The support of $Q_{1},Q_{2},Q_{3}$ must have diameter $>C$ by assumption. We can then examine the states stabilized by these $n-3$ local Pauli strings and ±$\pm Q_{1},\pm Q_{2},\pm Q_{3}$, get $7$ brother states for $|S\rangle$ and leading to a contradiction. Therefore, we proved the claim that there are $n-2$ independent $C$-local Pauli strings that stabilize $|S\rangle$. Since the deformed code states in $U\Pi U^{\dagger}$ are locally indistinguishable from $|S\rangle$, these $n-2$ local Pauli strings also fix other code states. Additionally, the constant-depth local circuit $U$ will degrade the distance of toric code, $L_{q}$, by a constant factor, so the deformed code family $U\Pi U^{\dagger}$ also has macroscopic distance.
Now we know that $U$ is a constant-depth local circuit connecting the toric code and another TSC such that the logical magic state $|\overline{T0}\rangle$ is mapped to a stabilizer state $|S\rangle$, thus is a logical stabilizer state in the deformed TSC. Now comes the key observation that this is impossible due to fundamental restriction on the constant-depth circuit mapping between $2$D TSCs, which is a consequence of the Bravyi–König theorem formulated in the context of fault-tolerant logical gates  Bravyi and König (2013). Here to prove “strong” LRM, we establish the following robust extension of the Bravyi–König theorem (detailed proof in Appendix B) which may be of independent interest in quantum error correction.
Theorem 2 (Robust Bravyi–König).
Suppose two $D$-dimensional ($D\geq 2$) TSCs are defined on the same lattice, both have $k$ logical qubits, with code projectors $\Pi_{1}=J_{1}J_{1}^{\dagger}$ and $\Pi_{2}=J_{2}J_{2}^{\dagger}$ respectively. If a constant-depth local unitary $U$ satisfies $\big{\|}U\Pi_{1}U^{\dagger}-\Pi_{2}\big{\|}=o(1)$, then the induced logical unitary satisfies
$\min_{Q\in\mathcal{C}_{k}^{(D)}}\big{\|}J_{2}^{\dagger}UJ_{1}-Q\big{\|}=o(1).$ (1)
Specifically, for $2$D TSCs, a constant-depth local circuit can only induce a logical Clifford between them, which makes it impossible to transform $|\overline{T0}\rangle$ into $|S\rangle$. Therefore, we conclude that $|\overline{T0}\rangle$ in toric code has LRM.
We comment that LRM can be generated by only one non-Clifford gate: $|\overline{T0}\rangle$ can be obtained by first preparing 2-qubit state $|T0\rangle$ with a non-Clifford gate $T=\operatorname{diag}(1,e^{i\pi/4})$, and then perform the Clifford encoding circuit for toric code. We also notice an intriguing property of topological nonstabilizerness of TCSs: tracing out qubits from the union of several contractable regions on the torus (see Fig. 1(a)) does not reduce the amount of nonstabilizerness in $|\overline{T0}\rangle$, because the erasure of such regions is correctable by Clifford recovery operations.
On the other hand, the hardness of erasing magic has useful implications for the important problem of transversal logical non-Clifford gates. We say $\{|\psi_{n}\rangle\}$ has $\mathrm{SRM}_{0}$, if their magic can be removed by tensor products of single-qubit gates. For example, product states, the logical $|T\rangle^{\otimes k}$ state in codes with transversal $T$ gates, and “no low-energy stabilizer states” constructed in Ref. Coble et al. (2023a) all have $\mathrm{SRM}_{0}$. For any $n$-qubit state $\rho$, unitary $U=U_{1}\otimes\cdots\otimes U_{n}$ and any subset $A\subset[n]$, we have $\sum_{P:\operatorname{supp}(P)=A}\operatorname{Tr}(P\rho)^{2}=\sum_{P:\operatorname{supp}(P)=A}\operatorname{Tr}(PU\rho U^{\dagger})^{2}$. Therefore, for $|\psi_{n}\rangle$ with $\mathrm{SRM}_{0}$, $\sum_{P:\operatorname{supp}(P)=A}\langle\psi_{n}|P|\psi_{n}\rangle^{2}$ are integers for all $A\subset[n]$. This provides an easily verified necessary condition for a state family to have $\mathrm{SRM}_{0}$. Take $|\psi_{n}\rangle$ be a logical nonstabilizer state in a stabilizer code, $\mathrm{SRM}_{0}$ will be a necessary condition for a stabilizer code to have the corresponding transversal logical gate. For example, consider logical state $|\overline{T0}\rangle$ in the toric code, and let $A$ be a handle which is the support of a logical $\overline{Z}_{1}$. Then by easy calculations, $\sum_{P:\operatorname{supp}(P)=A}\langle\overline{T0}|P|\overline{T0}\rangle^{2}=1/2$, which rules out $|\overline{T0}\rangle$ from having $\mathrm{SRM}_{0}$, thereby ruling out the probability of performing transversal logical $T$ on toric code.
IV Topological order and long-range magic phases
Fundamentally characterized by long-range entanglement, topological order represents an intrinsically quantum notion of phases of matter and has been of central interest in modern quantum many-body physics. Here we discuss how the strengthened concept of LRM yields novel perspectives on topological order. We frame our discussion in terms of qubit systems for clarity of presentation, though all results hold for qudits with arbitrary dimension. The toric code serves as a representative example of topological order, and all translation invariant 2D TSCs belong to the same phase as the decoupled copies of toric codes Bombin et al. (2012); Bombín (2014); Haah (2021). The LRM/SRM property of states in a topological order further reveals the many-body entanglement structure and can be instrumental for the classification of quantum phases.
The quantum double model Kitaev (2003b) and the string-net model Levin and Wen (2005) provide local Hamiltonian realizations of topological orders, describing topological orders by ground spaces of local Hamiltonians.
Definition 2.
We say a topological order is an SRM phase, if there exists a local Hamiltonian realization of it with a family of ground stabilizer states. We say a topological order is an (strong) LRM phase, if for any local Hamiltonian realization of it, all ground states exhibit (strong) LRM.
In Definition 2 we consider LRM/SRM defined with regard to geometrically local circuits to align with the physics literature, but it is worth stressing that all results also hold for all-to-all connectivity (by replacing TCS by constant weight stabilizer code).
Utilizing the TQO condition Bravyi et al. (2006, 2010) which asserts that the ground space describing a topological order forms a quantum error-correcting code with macroscopic distance, we can demonstrate that if a family of states in ground space exhibits SRM—that is, they are mapped to stabilizer states by constant-depth local circuits—then the entire ground space is mapped to a TSC by same circuits. This implies that the topological order can be described by this TSC. We arrive at the following theorem, with the detailed proof provided in Appendix C.
Theorem 3.
A topological order is a LRM phase iff it cannot be realized by any TSC, and is an SRM phase iff it can be realized by a TSC. If a topological order cannot be realized approximately up to $o(1)$ error in projector’s distance by any TSC, then it is a strong LRM phase.
For 2D topological orders, previous arguments relate TSC-realizability to whether the orders are Abelian or non-Abelian Potter and Vasseur (2016); Ellison et al. (2022). Combining these results with Theorem 3:
Remark.
Any 2D non-Abelian topological order is an LRM phase, and 2D Abelian topological order that admit gapped boundaries is an SRM phase.
Noting that the SRM/LRM phase is defined for arbitrary subspace families, it generalizes the Abelian/non-Abelian classification beyond 2D topological orders, encompassing higher-dimensional topological orders and local Hamiltonians with general connectivity.
Based on the difference in ground space degeneracy between a topological order and TSCs, we can further rigorously establish a stronger result:
Theorem 4.
Suppose the ground space dimension of a topological order is not a power of 2, then this topological order is a strong LRM phase.
The proof of this theorem is given in Appendix C.1. Intuitively, the “strong” property is ensured by the energy gap of conventional topological order. Specifically, if a family of ground states $\{|\phi_{n}\rangle\}$ of a topological order described by Hamiltonians $\{H^{(n)}\}$ satisfies $\big{\|}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=o(1/n)$ for constant-depth local circuits $\{U_{n}\}$ and stabilizer states $\{|S_{n}\rangle\}$, then $U_{n}$ maps the ground space of $H^{(n)}$ close to a TSC given by the code state $|S_{n}\rangle$. The $o(1/n)$ distance of $U_{n}|\phi_{n}\rangle$ and $|S_{n}\rangle$ implies all code states of the TSC have energy at most $o(1)$ above the ground energy of the deformed Hamiltonian $U_{n}H^{(n)}U_{n}^{\dagger}$. Since $\{H^{(n)}\}$ is gapped, the logical dimension of the TSC is at most the ground space degeneracy of $H^{(n)}$. The converse also holds similarly, implying that the dimension of the TSC equals the ground space degeneracy of $H^{(n)}$.
Example 2.
(Informal) For the entire phase of Fibonacci topological order on certain manifolds, all ground states exhibit strong LRM.
Consider the Fibonacci code defined on a hollow cylinder $\Sigma_{2}=S^{2}-(A^{1}\cup A^{2})$ with discs $A^{1}$ and $A^{2}$ both labeled by $1$ Koenig et al. (2010). The ribbon graphs on $\Sigma_{2}$ consistent with labeling $(1,1)$ span a subspace $\mathcal{H}_{\Sigma_{2}}^{(1,1)}$ with $\dim(\mathcal{H}_{\Sigma_{2}}^{(1,1)})=3$. The Fibonacci code realized by the ground space of the Levin–Wen Hamiltonian $H=-\sum_{\text{plaquettes }p}B_{p}-\sum_{\text{vertices }v}Q_{v}$ is isomorphic to $\mathcal{H}_{\Sigma_{2}}^{(1,1)}$ Koenig et al. (2010). Therefore, as the lattice becomes denser, all ground states of $H$ have strong LRM. Also, the degeneracy of Fibonacci code on closed higher-genus surfaces is not generally a power of $2$ Schotte et al. (2022), implying that their ground states also exhibit strong LRM.
V Long-range magic from correlations
Correlation functions play a crucial role in studying long-range entanglement  Zeng et al. (2015), serving to reveal and quantify the quantum correlations between distant points in a many-body system. Here, we establish a scheme for identifying LRM (including all-to-all cases) from two-point “magical” correlations. We begin by examining the constraints on correlations generated by SRM states and then utilize the violation of these constraints to demonstrate the LRM property in families such as ${|\mathrm{C}^{n-1}Z\rangle}$ and GHZ-type states.
Most qubit pairs in an SRM state are only correlated by a stabilizer state (see Fig. 1(b)), such correlations are rigid and can be prepared from EPR pairs. We have the following theorem, whose proof is given in Appendix D.1, and applies to constant-depth circuits with arbitrary connectivity.
Theorem 5.
(Informal) Suppose $\{|\psi_{n}\rangle\}$ is a family of SRM states, then any two-qubit reduced state of $|\psi_{n}\rangle$ can be obtained by local operations on $O(1)$ EPR pairs, if the backward light-cone of these two qubits do not intersect.
Using the violation of these constraints, an interesting example with LRM is given by the family of hypergraph states $|\mathrm{C}^{n-1}Z\rangle:=\mathrm{C}^{n-1}Z|+^{n}\rangle$ with $\mathrm{C}^{n-1}Z=\operatorname{diag}(1,\cdots,1,-1)$ being the $n$-qubit multi-controlled-$Z$ gate. They are also of strong experimental interest since multi-controlled-$Z$ gates are native gates on various important physical platforms for quantum computing Kim et al. (2022); Nguyen et al. (2024); Bluvstein et al. (2024); Wang et al. (2024). By measuring $X$ on arbitrary two qubits $\{k,l\}\subset[n]$, we get $\langle X_{k}X_{l}\rangle:=\langle\mathrm{C}^{n-1}Z|X_{k}X_{l}|\mathrm{C}^{n-1}Z\rangle=1-2^{2-n}$. On single qubits we also have $\langle X_{k}\rangle=\langle X_{l}\rangle=1-2^{2-n}$. Detailed derivations are left to Appendix D. We demonstrate in Proposition 18 in Appendix D that, for $P$ a single-qubit Pauli operator, $K\geq 1$ EPR pairs can generate a two-qubit state $\rho$ by local completely-positive trace preserving (CPTP) maps, with $\langle P\otimes P\rangle_{\rho}=b$ and $\langle P\otimes\mathbb{I}\rangle_{\rho}=\langle\mathbb{I}\otimes P\rangle_{\rho}=c$, if and only if $(b,c)$ lies within the feasible regions surrounded by the blue lines shown in Fig. 2. For $P=X$ and $\rho$ the two-point reduced density matrix of $|\mathrm{C}^{n-1}Z\rangle$, we have $b=c=1-2^{2-n}$. For arbitrary constant $K$, $(1-2^{2-n},1-2^{2-n})$ can always stay outside the feasible region when $n$ is sufficiently large.
Figure 2: The “feasible region” is the area bounded by the blue lines. If a bipartite two-qubit state $\rho$ has $(b,c)$ coordinate located outside the feasible region, then $\rho$ cannot be obtained by local CPTP maps on $K$ EPR pairs. The blue boundary lines consist of two linear functions and $2^{K}$ quadratic functions. We plot the cases for $K=2,3$ as examples, with the general cases being similar. The feasible region converges to a triangle when $K\rightarrow\infty$. The coordinates $(b,c)=(1-2^{2-n},1-2^{2-n})$ given by $|\mathrm{C}^{n-1}Z\rangle$ are plotted on the orange line $b=c$, which moves outside the feasible region for any fixed $K$ when $n$ is sufficiently large.
Corollary 6.
$\{|\mathrm{C}^{n-1}Z\rangle\}_{n=3}^{\infty}$ has LRM.
Suppose on the contrary that there exists $\{n_{q}\}_{q=1}^{\infty}\subset\mathbb{Z}_{>0}$ such that $\{|\mathrm{C}^{n_{q}-1}Z\rangle=U_{n_{q}}|S_{n_{q}}\rangle\}$ has SRM. Denote $M_{q}=\{j\in[n_{q}]:\mathcal{L}(j)\cap\mathcal{L}(1)\neq\emptyset\}$ be the set of qubits whose light-cone with regard to $U_{n_{q}}$ intersects with that of qubit $1$. We have $\left|M_{q}\right|=O(1)$. By Theorem 5, for all qubits $k\notin M_{q}$, the reduced state $\operatorname{Tr}_{\overline{\{1,k\}}}|\mathrm{C}^{n_{q}-1}Z\rangle\langle\mathrm{C}^{n_{q}-1}Z|$ can be generated by local CPTP maps on $O(1)$ EPR pairs. However, we know that when $n_{q}$ is sufficiently large, $\operatorname{Tr}_{\overline{\{1,k\}}}|\mathrm{C}^{n_{q}-1}Z\rangle\langle\mathrm{C}^{n_{q}-1}Z|$ cannot be generated by local CPTP maps on $O(1)$ EPR pairs, for all $k\neq 1$, leading to a contradiction. Therefore, $\{|\mathrm{C}^{n-1}Z\rangle\}$ has LRM.
Simpler examples include GHZ-type state families $\left\{\sqrt{1-\alpha}|0^{n}\rangle+\sqrt{\alpha}e^{i\theta}|1^{n}\rangle\right\}$, for which we have $\langle Z_{k}Z_{l}\rangle=1$ and $\langle Z_{k}\rangle=1-2\alpha$ for all positions $\{k,l\}\subset[n]$ and all values of $\theta$. When $1-2\alpha$ is not a finite-length binary decimal, the feasible region shown in Fig. 2 does not contain $(1,1-2\alpha)$ for arbitrary $K\geq 1$, implying that this family exhibits LRM.
VI Classicality of short-range magic and NLTM conjecture
A family of SRM states $\{|\psi_{n}\rangle=U_{n}|S_{n}\rangle\}$ admits an efficient classical description by storing the quantum circuits implementing $U_{n}$ and the stabilizer tableaux of $|S_{n}\rangle$. The expectation value of any $O(1)$-weight observable $O$ on $|\psi_{n}\rangle=U_{n}|S_{n}\rangle$ can be computed exactly on classical computers in time $\mathrm{poly}(n)$. This follows from the fact that $U_{n}^{\dagger}OU_{n}$ acts nontrivially on only $O(1)$ qubits, can be decomposed into the sum of $O(1)$ Pauli strings, and the expectation value of $|S_{n}\rangle$ on Pauli strings can be computed efficiently Aaronson and Gottesman (2004). Consequently, SRM states can serve as classical proofs for quantum PCP Hamiltonians Aharonov et al. . Based on this, we formulate the no low-energy trivial magic (NLTM) conjecture (also mentioned in a talk Nirkhe (2023)):
Conjecture 1 (NLTM).
There exists a constant $\epsilon>0$ and a family of $O(1)$-local commuting Hamiltonians $\{H^{(n)}=\sum_{i=1}^{m^{(n)}}H_{i}^{(n)}\}$ acting on $n$ qubits, with $m^{(n)}=\Theta(n)$ and $\|H_{i}^{(n)}\|_{\infty}\leq 1$, such that any family of states $\{|\psi_{n}\rangle\}$ satisfying $\langle\psi_{n}|H^{(n)}|\psi_{n}\rangle<\lambda_{\min}(H^{(n)})+\epsilon n$ has LRM.
As detailed in Appendix E, NLTM is a necessary consequence of the quantum PCP conjecture. A resolution of NLTM would provide further insights and evidence for quantum PCP. We extend the notion of SRM states to mixed SRM states, defined as the reduced density matrices of pure SRM states. Like pure SRM states, mixed SRM states can also serve as classical proofs. NLTM is strictly stronger than NLTS Freedman and Hastings ; Anshu et al. (2023), as SRM states, which serve as classical proofs, generalize short-range entangled states. In a sense, NLTM unifies different forms of efficient classical proofs enabled by trivial entanglement and classical computational complexity  Nirkhe (2022); Coble et al. (2023b), thereby providing a more refined condition for the quantum PCP conjecture. Given our findings of “no lowest-energy trivial magic” from LRM phases of matter, a potential path towards a full resolution of the NLTM conjecture is to extend LRM phases, e.g., Fibonacci-type topological orders beyond geometrical settings, in light of the use of good qLDPC codes in the NLTS proof Anshu et al. (2023).
SRM states can be prepared efficiently. If the graph representing the connectivity of qubits is connected, any $n$-qubit Clifford unitary can be synthesized using $O(n^{2})$ gates Aaronson and Gottesman (2004); Nash et al. (2020). Therefore, all $n$-qubit states with SRM can be prepared with $O(n^{2})$ gates, and states with gate complexity $\omega(n^{2})$ exhibit LRM. With the gate complexity upper bounded by $O(n^{2})$, it follows from Ref. Zhao et al. (2024) that all SRM states can be learned to $\epsilon$-precision in trace distance with high probability, using sample complexity $\widetilde{O}(n^{2}\epsilon^{-2})$. Note that even with all-to-all connectivity, there exist states that require $\Omega(2^{n})$ gates Plesch and Brukner (2011); Sun et al. (2023) to prepare. Haar random states exhibit LRM with probability approaching $1$ as $n\rightarrow\infty$. All-to-all random quantum circuits (Chen et al. (2024) Corollary 1.9) with $\omega(n^{6})$ gates can create a state with gate complexity $\omega(n^{2})$, and thus LRM, with probability approaching $1$.
VII Concluding remarks
Motivated by the intriguing interplay between entanglement and nonstabilizerness, we propose the concept of LRM and systematically explore its properties, manifestations, and significance, from various perspectives including quantum error-correcting codes, topological order, and correlation functions. Our study provides a systematic theory of LRM that builds on and generates valuable insights for wide-ranging areas in quantum computing, complexity, and many-body physics.
Various intriguing open questions remain. The relationship between LRM and nonlocal magic merits further study, and we provide a preliminary discussion in Appendix F. The study of local state transformations will further contribute to proving LRM from correlations, utilizing more information from the two-point reduced density matrix beyond merely the diagonal elements. From the perspective of LRM theory itself, valuable future tasks include e.g. developing quantitative characterizations for LRM (a natural possibility being magic variants of topological entanglement entropies), bounding the circuit depth required to erase LRM in various scenarios, and exploring LRM in mixed states. Also, the question of learning SRM states with polynomial computational complexity Huang et al. (2024); Landau and Liu (2024) warrants further investigation. Other directions with promising prospects include: relation with measurement-assisted complexity of many-body systems Tantivasadakarn et al. (2023); Gu et al. ; further applications for QEC and fault tolerance from code property and logical gate perspectives; a more fundamental understanding of the $1/n$ critical error scaling, potentially in relation to approximate QEC codes where similar phenomena emerge in fundamental ways Faist et al. (2020); Liu and Zhou (2023); Yi et al. (2024)). In addition, we anticipate the LRM theory to be a good starting point for exploring the profound interplay between many-body physics and computation. Further enriching the theory of LRM from these perspectives is expected to yield deeper insights into the diverse fields connected by LRM and the interactions between them.
Acknowledgements.
We thank Soonwon Choi, Tyler Ellison, Andi Gu, Yu-Jie Liu, Zimu Li, Zhengwei Liu, Zhenhuan Liu, Natalie Parham, Yuguo Shao, Jinmin Yi and Zishuo Zhao for valuable discussions and feedback. Z.-W.L. is supported in part by a startup funding from YMSC, Tsinghua University, and NSFC under Grant No. 12475023. While finalizing this paper, we noticed the release of an independent work Korbany et al. which considers similar notions of LRM but adopts different perspectives and focuses on different types of systems (1D matrix product states).
Appendix A Strong LRM from limitations of logical gates for topological stabilizer codes
We define two edges on a lattice as adjacent if they share a common vertex. We define the distance between two edges to be the length of the shortest path along the lattice connecting the two edges. For a set $E$ of edges on a lattice, the diameter $\operatorname{Diam}(E)$ is defined as the smallest integer $\delta$ such that every pair of edges in $E$ is within distance $\delta$ of each other. For an operator $O$ acting on qubits associated with the edges, we also denote $\operatorname{Diam}(O):=\operatorname{Diam}(\operatorname{supp}(O))$, where $\operatorname{supp}(O)$ denotes the support of $O$.
Lemma 7.
Suppose two $n$-qubit pure states $|\psi_{1}\rangle$ and $|\psi_{1}\rangle$ satisfies $\operatorname{Tr}_{\overline{A}}|\psi_{1}\rangle\langle\psi_{1}|=\operatorname{Tr}_{\overline{A}}|\psi_{2}\rangle\langle\psi_{2}|$ for some region $A\subset[n]$, and observable $O$ satisfies $\operatorname{supp}(O)\subset A$, then $\langle\psi_{1}|O|\psi_{1}\rangle=\langle\psi_{2}|O|\psi_{2}\rangle$.
Proof.
Since $\operatorname{Tr}_{\overline{\operatorname{supp}(O)}}|\psi_{1}\rangle\langle\psi_{1}|=\operatorname{Tr}_{\overline{\operatorname{supp}(O)}}|\psi_{2}\rangle\langle\psi_{2}|$, we know
$\langle\psi_{1}|O|\psi_{1}\rangle=$ $\operatorname{Tr}(|\psi_{1}\rangle\langle\psi_{1}|O)=\operatorname{Tr}\left(\operatorname{Tr}_{\overline{\operatorname{supp}(O)}}(|\psi_{1}\rangle\langle\psi_{1}|)O|_{\operatorname{supp}(O)}\right)$ (2)
$=$ $\operatorname{Tr}\left(\operatorname{Tr}_{\overline{\operatorname{supp}(O)}}(|\psi_{2}\rangle\langle\psi_{2}|)O|_{\operatorname{supp}(O)}\right)=\operatorname{Tr}(|\psi_{2}\rangle\langle\psi_{2}|O)=\langle\psi_{2}|O|\psi_{2}\rangle,$ (3)
where $O|_{\operatorname{supp}(O)}$ denotes the restriction of $O$ to its support. ∎
Lemma 8.
Suppose $H$ is a Hamiltonian acting on a finite-dimensional Hilbert space $\mathcal{H}$, with eigenvalues $E_{0}<E_{1}<E_{2}<\cdots$. Let $D_{0}$ be the degeneracy of the ground-state space with energy $E_{0}$. For an orthonormal state set $\{|\phi_{1}\rangle,\cdots,|\phi_{t}\rangle\}$, if $\langle\phi_{i}|H|\phi_{i}\rangle<E_{0}+\frac{E_{1}-E_{0}}{D_{0}+1}$ for $i=1,\cdots,t$, then $t\leq D_{0}$.
Proof.
Denote $\Pi_{0},\Pi_{1},\Pi_{2},\cdots$ be the projectors onto the eigenspaces of $E_{0},E_{1},E_{2},\cdots$. For $i=1,\cdots,t$ we have
$\langle\phi_{i}|H|\phi_{i}\rangle=$ $E_{0}\langle\phi_{i}|\Pi_{0}|\phi_{i}\rangle+\sum_{i\geq 1}E_{i}\langle\phi_{i}|\Pi_{i}|\phi_{i}\rangle$ (4)
$\geq$ $E_{0}\langle\phi_{i}|\Pi_{0}|\phi_{i}\rangle+E_{1}\sum_{i\geq 1}\langle\phi_{i}|\Pi_{i}|\phi_{i}\rangle$ (5)
$=$ $E_{0}\langle\phi_{i}|\Pi_{0}|\phi_{i}\rangle+E_{1}\langle\phi_{i}|(\mathbb{I}-\Pi_{0})|\phi_{i}\rangle$ (6)
$=$ $E_{1}-(E_{1}-E_{0})\langle\phi_{i}|\Pi_{0}|\phi_{i}\rangle.$ (7)
Combined with $\langle\phi_{i}|H|\phi_{i}\rangle<E_{0}+\frac{E_{1}-E_{0}}{D_{0}+1}$ we know $\langle\phi_{i}|\Pi_{0}|\phi_{i}\rangle>\frac{D_{0}}{D_{0}+1}$. Denote $\Pi^{\prime}=\sum_{i=1}^{t}|\phi_{i}\rangle\langle\phi_{i}|$, we have
$\operatorname{Tr}(\Pi^{\prime}\Pi_{0})=\sum_{i=1}^{t}\langle\phi_{i}|\Pi_{0}|\phi_{i}\rangle>\frac{tD_{0}}{D_{0}+1}.$ (8)
On the other hand, since
$\operatorname{Tr}(\Pi^{\prime}\Pi_{0})\leq\left\|\Pi^{\prime}\right\|_{\infty}\left\|\Pi_{0}\right\|_{1}=D_{0},$ (9)
we know $t<D_{0}+1$, that is, $t\leq D_{0}$. ∎
We prove the following theorem, which generalizes the toric code case discussed in the main text to general TSCs and its “strong” version.
Theorem 9.
For a $D$-dimensional ($D\geq 2$) TSC family and a $k$-qubit state $|\phi\rangle$, if $|\phi\rangle\notin\big{\{}V^{\dagger}|0^{k}\rangle\mid V\in\mathcal{C}_{k}^{(D)}\big{\}}$, where $k$ is the number of logical qubits, then the corresponding logical state $|\overline{\phi}\rangle$ has strong LRM.
Proof.
Suppose on the contrary that the logical state family $\{|\overline{\phi}_{n}\rangle\}$, where $n$ is the number of qubits, does not have strong LRM. Then, there exists a sequence $T\subset\mathbb{Z}_{>0}$, a family of constant-depth local circuits $\{U_{n}\}_{n\in T}$, a family of stabilizer states $\{|S_{n}\rangle\}_{n\in T}$, such that for $n\in T$ we have
$\big{\|}U_{n}|\overline{\phi}_{n}\rangle\langle\overline{\phi}_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=o(1/n).$ (10)
Denote projectors onto the code space of this family of TSC as $\{\Pi^{(n)}\}$. Suppose this family of TSC has local Pauli generators $\{Q_{1}^{(n)},\cdots,Q_{n-k}^{(n)}\}$. Denote $d_{n}$ be the code distance of $\{\Pi^{(n)}\}$, satisfying $d_{n}\rightarrow\infty$ as $n\rightarrow\infty$. Suppose the light-cone sizes of $\{U_{n}\}$ are upper bounded by a constant $\eta$.
Denote $C$ as a constant independent of $n$ that upper bounds the diameters of all deformed generators $U_{n}Q_{i}^{(n)}U_{n}^{\dagger}$. Define $\mathcal{B}_{n}$ as the set of states that are $C$-locally indistinguishable from $U_{n}|\overline{\phi}_{n}\rangle$,
$\mathcal{B}_{n}:=\Big{\{}|\psi\rangle\Big{|}\operatorname{Tr}_{\overline{J}}|\psi\rangle\langle\psi|=\operatorname{Tr}_{\overline{J}}(U_{n}|\overline{\phi}_{n}\rangle\langle\overline{\phi}_{n}|U_{n}^{\dagger}),\forall J\subset[n]\text{ with }\operatorname{Diam}(J)\leq C\Big{\}}.$ (11)
For an arbitrary $|\psi\rangle$ satisfying $\langle\psi|U_{n}\Pi^{(n)}U_{n}^{\dagger}|\psi\rangle=1$, $U_{n}^{\dagger}|\psi\rangle$ is a logical state of $\Pi_{n}$, and thus indistinguishable from $|\overline{\phi}_{n}\rangle$ when restricted to regions of size $\leq d_{n}-1$. Since the light-cone sizes of $U_{n}$ are upper bounded by $\eta$, we know $|\psi\rangle$ is indistinguishable from $U_{n}|\overline{\phi}_{n}\rangle$ when restricted to regions of size $\leq\lfloor(d_{n}-1)/\eta\rfloor$. $d_{n}\rightarrow\infty$ implies $|\psi\rangle\in\mathcal{B}_{n}$ when $n\in T$ is sufficiently large. On the other hand, since all $|\psi\rangle\in\mathcal{B}_{n}$ is $C$-locally indistinguishable from $U_{n}|\overline{\phi}_{n}\rangle$ and the diameters of $U_{n}Q_{i}^{(n)}U_{n}^{\dagger}$’s are upper bounded by $C$, by Lemma 7 we know
$\langle\psi|U_{n}Q_{i}^{(n)}U_{n}^{\dagger}|\psi\rangle=\langle\overline{\phi}_{n}|U_{n}^{\dagger}U_{n}Q_{i}^{(n)}U_{n}^{\dagger}U_{n}|\overline{\phi}_{n}\rangle=\langle\overline{\phi}_{n}|Q_{i}^{(n)}|\overline{\phi}_{n}\rangle=1$ (12)
for $i=1,\cdots,n-k$, implying $\langle\psi|U_{n}\Pi^{(n)}U_{n}^{\dagger}|\psi\rangle=1$. Therefore, $\mathcal{B}_{n}$ is exactly the set of pure states in the deformed code $U_{n}\Pi^{(n)}U_{n}^{\dagger}$:
$\mathcal{B}_{n}=\big{\{}|\psi\rangle\mid\langle\psi|U_{n}\Pi^{(n)}U_{n}^{\dagger}|\psi\rangle=1\big{\}}.$ (13)
Define $\mathcal{A}_{n}$ as the set of states that are $C$-locally indistinguishable from $|S_{n}\rangle$,
$\mathcal{A}_{n}:=\Big{\{}|\psi\rangle\Big{|}\operatorname{Tr}_{\overline{J}}|\psi\rangle\langle\psi|=\operatorname{Tr}_{\overline{J}}|S_{n}\rangle\langle S_{n}|,\forall J\subset[n]\text{ with }\operatorname{Diam}(J)\leq C\Big{\}}.$ (14)
Denote $G_{n}\subset\mathcal{P}_{n}$ be the stabilizer group of $|S_{n}\rangle$. Let $r_{n}$ be the maximum number of independent Pauli strings in $\{g\in G_{n}\mid\operatorname{Diam}(g)\leq C\}$. Pick such an independent set $\{P_{1}^{(n)},\cdots,P_{r_{n}}^{(n)}\}\subset\{g\in G_{n}\mid\operatorname{Diam}(g)\leq C\}$. We must have
$\{g\in G_{n}\mid\operatorname{Diam}(g)\leq C\}\subset\langle P_{1},\cdots,P_{r}\rangle,$ (15)
because otherwise we would be able to find $r+1$ independent Pauli strings in $\{g\in G_{n}\mid\operatorname{Diam}(g)\leq C\}$. Let
$\mathcal{A}_{n}^{\prime}:=\big{\{}|\psi\rangle\mid P_{i}^{(n)}|\psi\rangle=|\psi\rangle,i=1,\cdots,r_{n}\big{\}}=\big{\{}|\psi\rangle\mid\langle\psi|\widetilde{\Pi}^{(n)}|\psi\rangle=1\big{\}},$ (16)
where $\widetilde{\Pi}^{(n)}:=\prod_{i=1}^{r_{n}}\frac{\mathbb{I}+P_{i}^{(n)}}{2}$.
Now we show that $\mathcal{A}_{n}\subset\mathcal{A}_{n}^{\prime}$, and $\mathcal{A}_{n}$ contains a set of $2^{n-r_{n}}$ orthonormal pure states (actually $\mathcal{A}_{n}=\mathcal{A}_{n}^{\prime}$, as we will show later). For $|a\rangle\in\mathcal{A}_{n}$, since $\operatorname{Diam}(P_{i}^{(n)})\leq C$, by Lemma 7 we know $\langle a|P_{i}^{(n)}|a\rangle=\langle S_{n}|P_{i}^{(n)}|S_{n}\rangle=1$ for $i=1,\cdots,r_{n}$. This implies $|a\rangle\in\mathcal{A}_{n}^{\prime}$ and thus $\mathcal{A}_{n}\subset\mathcal{A}_{n}^{\prime}$.
We take $\{P_{r_{n}+1}^{(n)},\cdots,P_{n}^{(n)}\}\subset G_{n}$ such that $\{P_{1}^{(n)},\cdots,P_{n}^{(n)}\}$ are independent. We have $\langle P_{1}^{(n)},\cdots,P_{n}^{(n)}\rangle=G_{n}$. Now we omit the upper script $(n)$ for $P_{i}^{(n)}$ for brevity. For $j=r_{n}+1,\cdots,n$, since $P_{j}\notin\langle P_{1},\cdots,P_{r_{n}}\rangle$, by Eq. (15) we know $\operatorname{Diam}(P_{j})>C$. For $\mathbf{b}=(b_{r_{n}+1},\cdots,b_{n})\in\{0,1\}^{n-r_{n}}$, consider the states
$|\psi_{\mathbf{b}}^{(n)}\rangle\langle\psi_{\mathbf{b}}^{(n)}|$ $:=\prod_{i=1}^{r_{n}}\frac{\mathbb{I}+P_{i}}{2}\prod_{j=r_{n}+1}^{n}\frac{\mathbb{I}+(-1)^{b_{j}}P_{j}}{2}$ (17)
$=\frac{1}{2^{n}}\sum_{\mathbf{x}\in\{0,1\}^{n}}\prod_{i=1}^{r_{n}}P_{i}^{x_{i}}\prod_{j=r_{n}+1}^{n}((-1)^{b_{j}}P_{j})^{x_{j}},$ (18)
where $\mathbf{x}=(x_{1},\cdots,x_{n})$. These states are orthogonal to each other. In particular, we have $|\psi_{\mathbf{0}}^{(n)}\rangle=|S_{n}\rangle$. Consider an arbitrary subset $J\subset[n]$ with $\operatorname{Diam}(J)\leq C$. For $\mathbf{x}\in\{0,1\}^{n}$ such that $(x_{r_{n}+1},\cdots,x_{n})\neq 0^{n-r_{n}}$, we must have
$\operatorname{Diam}\left(\prod_{i=1}^{r_{n}}P_{i}^{x_{i}}\prod_{j=r_{n}+1}^{n}((-1)^{b_{j}}P_{j})^{x_{j}}\right)=\operatorname{Diam}\left(\prod_{i=1}^{r_{n}}P_{i}^{x_{i}}\prod_{j=r_{n}+1}^{n}P_{j}^{x_{j}}\right)>C,$ (19)
otherwise $\prod_{i=1}^{r_{n}}P_{i}^{x_{i}}\prod_{j=r_{n}+1}^{n}P_{j}^{x_{j}}\in\{g\in G_{n}\mid\operatorname{Diam}(g)\leq C\}\subset\langle P_{1},\cdots,P_{r_{n}}\rangle$, which would contradict the independence $\{P_{1},\cdots,P_{n}\}$. Therefore, for arbitrary $\mathbf{b}\in\{0,1\}^{n-r_{n}}$, when $(x_{r_{n}+1},\cdots,x_{n})\neq 0^{n-r_{n}}$ we have
$\operatorname{Tr}_{\overline{J}}\left(\prod_{i=1}^{r_{n}}P_{i}^{x_{i}}\prod_{j=r_{n}+1}^{n}((-1)^{b_{j}}P_{j})^{x_{j}}\right)=0.$ (20)
Denote $\mathbf{y}=(y_{1},\cdots,y_{r_{n}})$, for all $\mathbf{b}\in\{0,1\}^{n-r_{n}}$ we have
$\operatorname{Tr}_{\overline{J}}|\psi^{(n)}_{\mathbf{b}}\rangle\langle\psi^{(n)}_{\mathbf{b}}|=$ $\frac{1}{2^{n}}\sum_{\mathbf{x}\in\{0,1\}^{n}}\operatorname{Tr}_{\overline{J}}\left(\prod_{i=1}^{r_{n}}P_{i}^{x_{i}}\prod_{j=r_{n}+1}^{n}((-1)^{b_{j}}P_{j})^{x_{j}}\right)$ (21)
$=$ $\frac{1}{2^{n}}\sum_{\mathbf{y}\in\{0,1\}^{r_{n}}}\operatorname{Tr}_{\overline{J}}\left(\prod_{i=1}^{r_{n}}P_{i}^{y_{i}}\prod_{j=r_{n}+1}^{n}((-1)^{b_{j}}P_{j})^{0}\right)$ (22)
$=$ $\frac{1}{2^{n}}\sum_{\mathbf{y}\in\{0,1\}^{r_{n}}}\operatorname{Tr}_{\overline{J}}\left(\prod_{i=1}^{r_{n}}P_{i}^{y_{i}}\prod_{j=r_{n}+1}^{n}P_{j}^{0}\right)$ (23)
$=$ $\frac{1}{2^{n}}\sum_{\mathbf{x}\in\{0,1\}^{n}}\operatorname{Tr}_{\overline{J}}\left(\prod_{i=1}^{r_{n}}P_{i}^{x_{i}}\prod_{j=r_{n}+1}^{n}P_{j}^{x_{j}}\right)=\operatorname{Tr}_{\overline{J}}|S\rangle\langle S|,$ (24)
Therefore, we find a set of $2^{n-r_{n}}$ orthonormal pure states $\{|\psi^{(n)}_{\mathbf{b}}\rangle\mid\mathbf{b}\in\{0,1\}^{n-r_{n}}\}\subset\mathcal{A}_{n}$. By $\mathcal{A}_{n}\subset\mathcal{A}_{n}^{\prime}$ we know $\dim(\operatorname{span}(\mathcal{A}_{n}))=\dim(\operatorname{span}(\mathcal{A}_{n}^{\prime}))=2^{n-r_{n}}$.
Now we show that $r_{n}=n-k$ when $n\in T$ is sufficiently large. For the original TSC $\Pi^{(n)}$, define Hamiltonian $H^{(n)}=\sum_{i=1}^{n-k}\frac{\mathbb{I}-Q_{i}^{(n)}}{2}$, which has eigenvalues $\{0,1,\cdots,n-k\}$ and ground space projector $\Pi^{(n)}$. For the subspace $\widetilde{\Pi}^{(n)}$ corresponding to $\mathcal{A}_{n}^{\prime}$, we define $\widetilde{H}^{(n)}=\sum_{j=1}^{r_{n}}\frac{\mathbb{I}-P^{(n)}_{j}}{2}$, which has eigenvalues $\{0,1,\cdots,r_{n}\}$ and ground space projector $\widetilde{\Pi}^{(n)}$. Intuitively, $U_{n}|\overline{\phi}_{n}\rangle\approx|S_{n}\rangle$ implies $U_{n}\Pi^{(n)}U_{n}^{\dagger}\approx\widetilde{\Pi}^{(n)}$.
We first show that $r_{n}=n-k$ when $n\in T$ is sufficiently large. For all $\mathbf{b}\in\{0,1\}^{n-r_{n}}$, since $|\psi^{(n)}_{\mathbf{b}}\rangle\in\mathcal{A}_{n}$ and $\operatorname{Diam}(U_{n}Q_{i}^{(n)}U_{n}^{\dagger})\leq C$, we know
$\langle\psi^{(n)}_{\mathbf{b}}|U_{n}H^{(n)}U_{n}^{\dagger}|\psi^{(n)}_{\mathbf{b}}\rangle=\sum_{i=1}^{n-k}\frac{1-\langle\psi^{(n)}_{\mathbf{b}}|U_{n}Q_{i}^{(n)}U_{n}^{\dagger}|\psi^{(n)}_{\mathbf{b}}\rangle}{2}=\sum_{i=1}^{n-k}\frac{1-\langle S_{n}|U_{n}Q_{i}^{(n)}U_{n}^{\dagger}|S_{n}\rangle}{2}=\langle S_{n}|U_{n}H^{(n)}U_{n}^{\dagger}|S_{n}\rangle,$ (25)
therefore
$\left|\langle\psi_{\mathbf{b}}|U_{n}H^{(n)}U_{n}^{\dagger}|\psi_{\mathbf{b}}\rangle-\langle\overline{\phi}_{n}|H^{(n)}|\overline{\phi}_{n}\rangle\right|=$ $\left|\langle S_{n}|U_{n}H^{(n)}U_{n}^{\dagger}|S_{n}\rangle-\langle\overline{\phi}_{n}|H^{(n)}|\overline{\phi}_{n}\rangle\right|$ (26)
$=$ $\left|\operatorname{Tr}\big{[}H^{(n)}\big{(}U_{n}^{\dagger}|S_{n}\rangle\langle S_{n}|U_{n}-|\overline{\phi}_{n}\rangle\langle\overline{\phi}_{n}|\big{)}\big{]}\right|$ (27)
$\leq$ $\sum_{i=1}^{n-k}\left|\operatorname{Tr}\Big{[}\frac{\mathbb{I}-Q_{i}^{(n)}}{2}\big{(}U_{n}^{\dagger}|S_{n}\rangle\langle S_{n}|U_{n}-|\overline{\phi}_{n}\rangle\langle\overline{\phi}_{n}|\big{)}\Big{]}\right|$ (28)
$\leq$ $\sum_{i=1}^{n-k}\big{\|}\frac{\mathbb{I}-Q_{i}^{(n)}}{2}\big{\|}_{\infty}\big{\|}|S_{n}\rangle\langle S_{n}|-U_{n}|\overline{\phi}_{n}\rangle\langle\overline{\phi}_{n}|U_{n}^{\dagger}\big{\|}_{1}$ (29)
$\leq$ $(n-k)o(1/n)=o(1).$ (30)
Therefore, for arbitrary $\mathbf{b}\in\{0,1\}^{n-r_{n}}$ we have
$\langle\psi^{(n)}_{\mathbf{b}}|U_{n}H^{(n)}U_{n}^{\dagger}|\psi^{(n)}_{\mathbf{b}}\rangle\leq\langle\overline{\phi}_{n}|H^{(n)}|\overline{\phi}_{n}\rangle+o(1)=o(1).$ (31)
Since $k$ is a constant independent of $n$, by Lemma 8 we know $2^{n-r_{n}}\leq 2^{k}$ when $n\in T$ is sufficiently large. For arbitrary $|b\rangle\in\mathcal{B}_{n}$,
$\left|\langle b|\widetilde{H}^{(n)}|b\rangle-\langle S_{n}|\widetilde{H}^{(n)}|S_{n}\rangle\right|=$ $\left|\langle\overline{\phi}_{n}|U_{n}^{\dagger}\widetilde{H}^{(n)}U_{n}|\overline{\phi}_{n}\rangle-\langle S_{n}|\widetilde{H}^{(n)}|S_{n}\rangle\right|$ (32)
$=$ $\left|\operatorname{Tr}\big{[}\widetilde{H}^{(n)}\big{(}U_{n}|\overline{\phi}_{n}\rangle\langle\overline{\phi}_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{)}\big{]}\right|$ (33)
$\leq$ $\sum_{i=1}^{r_{n}}\left|\operatorname{Tr}\Big{[}\frac{\mathbb{I}-P_{i}^{(n)}}{2}\big{(}U_{n}|\overline{\phi}_{n}\rangle\langle\overline{\phi}_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{)}\Big{]}\right|$ (34)
$\leq$ $\sum_{i=1}^{r_{n}}\big{\|}U_{n}|\overline{\phi}_{n}\rangle\langle\overline{\phi}_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}$ (35)
$\leq$ $r_{n}o(1/n)=o(1).$ (36)
Therefore, for arbitrary $|b\rangle\in\mathcal{B}_{n}$ we have
$\langle b|\widetilde{H}^{(n)}|b\rangle\leq\langle S_{n}|\widetilde{H}^{(n)}|S_{n}\rangle+o(1)=o(1).$ (37)
We already know the ground space dimension $2^{n-r_{n}}$ of $\widetilde{H}^{(n)}$ is upper bounded by $2^{k}$. When $n\in T$ is sufficiently large, $\langle b|\widetilde{H}^{(n)}|b\rangle\leq\frac{1}{2^{k}+1}\leq 0+\frac{1-0}{2^{n-r_{n}}+1}$. By Lemma 8 we know $2^{k}\leq 2^{n-r_{n}}$. We conclude that $r_{n}=n-k$ when $n\in T$ is sufficiently large.
Now we show $U_{n}\Pi^{(n)}U_{n}^{\dagger}\approx\widetilde{\Pi}^{(n)}$. Since $\widetilde{H}^{(n)}\geq\mathbb{I}-\widetilde{\Pi}^{(n)}$, for arbitrary $|b\rangle\in\mathcal{B}_{n}$,
$\langle b|\widetilde{H}^{(n)}|b\rangle\geq\langle b|(\mathbb{I}-\widetilde{\Pi}^{(n)})|b\rangle=1-\langle b|\widetilde{\Pi}^{(n)}|b\rangle.$ (38)
By Eq. (37) we know
$\langle b|\widetilde{\Pi}^{(n)}|b\rangle\geq 1-\langle b|\widetilde{H}^{(n)}|b\rangle=1-o(1).$ (39)
We take $\{|b_{1}\rangle,\cdots,|b_{2^{k}}\rangle\}$ be a set of orthogonal basis of $\operatorname{span}(\mathcal{B}_{n})$, satisfying $\sum_{i=1}^{2^{k}}|b_{i}\rangle\langle b_{i}|=U_{n}\Pi^{(n)}U_{n}^{\dagger}$, then
$\operatorname{Tr}\big{(}U_{n}\Pi^{(n)}U_{n}^{\dagger}\widetilde{\Pi}^{(n)}\big{)}=\sum_{i=1}^{2^{k}}\langle b_{i}|\widetilde{\Pi}^{(n)}|b_{i}\rangle\geq 2^{k}(1-o(1)).$ (40)
Therefore we have
$\big{\|}U_{n}\Pi^{(n)}U_{n}^{\dagger}-\widetilde{\Pi}^{(n)}\big{\|}_{2}^{2}=$ $\operatorname{Tr}\left(\big{(}U_{n}\Pi^{(n)}U_{n}^{\dagger}-\widetilde{\Pi}^{(n)}\big{)}^{2}\right)$ (41)
$=$ $\operatorname{Tr}\big{(}U_{n}\Pi^{(n)}U_{n}^{\dagger}\big{)}+\operatorname{Tr}\big{(}\widetilde{\Pi}^{(n)}\big{)}-2\operatorname{Tr}\big{(}U_{n}\Pi^{(n)}U_{n}^{\dagger}\widetilde{\Pi}^{(n)}\big{)}$ (42)
$\leq$ $2^{k}+2^{k}-2\cdot 2^{k}(1-o(1))=2^{k+1}o(1)=o(1),$ (43)
implying $\big{\|}U_{n}\Pi^{(n)}U_{n}^{\dagger}-\widetilde{\Pi}^{(n)}\big{\|}_{2}=o(1)$.
Denote $\widetilde{d}_{n}$ be the distance of the stabilizer code $\widetilde{\Pi}^{(n)}$. Recall that $d_{n}$ is the distance of $\Pi_{n}$, and $\eta$ upper bounds the light-cone sizes of $\{U_{n}\}$. Therefore, the deformed code $U_{n}\Pi^{(n)}U_{n}^{\dagger}$ has distance $\geq\lfloor\frac{d_{n}-1}{\eta}\rfloor+1$. Now we show that when $n\in T$ is sufficiently large, $\widetilde{\Pi}^{(n)}$ is a code with distance
$\widetilde{d}_{n}\geq\lfloor\frac{d_{n}-1}{\eta}\rfloor+1,$ (44)
which tends to infinity as $n\rightarrow\infty$. We remark that this implies all states in $\mathcal{A}_{n}^{\prime}$ are $C$-locally indistinguishable, thus we have $\mathcal{A}_{n}=\mathcal{A}_{n}^{\prime}$.
Suppose on the contrary that there exists a subsequence $T^{\prime}\subset T$, such that $\widetilde{d}_{n}\leq\lfloor\frac{d_{n}-1}{\eta}\rfloor$ for all $n\in T^{\prime}$. Denote $K_{n}$ be the stabilizer group corresponding to $\widetilde{\Pi}^{(n)}$. Then there exists a family of Pauli strings $\{E_{n}\}_{n\in T^{\prime}}$, where $E_{n}\in\mathcal{P}_{n}$ has phase $+1$ and satisfies $\mathrm{wt}(E_{n})=\widetilde{d}_{n}\leq\lfloor\frac{d_{n}-1}{\eta}\rfloor$, such that $E_{n}$ is a logical error for $\widetilde{\Pi}^{(n)}$. This implies that $\mathbb{I}_{2}^{\otimes n},i\mathbb{I}_{2}^{\otimes n},-\mathbb{I}_{2}^{\otimes n},-i\mathbb{I}_{2}^{\otimes n}\notin\{E_{n}h\mid h\in K_{n}\}$. We have
$\widetilde{\Pi}^{(n)}E_{n}\widetilde{\Pi}^{(n)}=E_{n}\widetilde{\Pi}^{(n)}\text{, and }\operatorname{Tr}(E_{n}\widetilde{\Pi}^{(n)})=\frac{1}{2^{n-k}}\sum_{h\in K_{n}}\operatorname{Tr}(E_{n}h)=0.$ (45)
Since $U_{n}\Pi^{(n)}U_{n}^{\dagger}$ has distance $\geq\lfloor\frac{d_{n}-1}{\eta}\rfloor+1$, we know there exist a constant $c_{n}\in\mathbb{R}$ with $\left|c_{n}\right|\leq 1$ such that
$(U_{n}\Pi^{(n)}U_{n}^{\dagger})E_{n}(U_{n}\Pi^{(n)}U_{n}^{\dagger})=c_{n}U_{n}\Pi^{(n)}U_{n}^{\dagger}.$ (46)
Therefore, we have
$\big{\|}E_{n}\widetilde{\Pi}^{(n)}-c_{n}U_{n}\Pi^{(n)}U_{n}^{\dagger}\big{\|}_{2}$ (47)
$=$ $\big{\|}\widetilde{\Pi}^{(n)}E_{n}\widetilde{\Pi}^{(n)}-(U_{n}\Pi^{(n)}U_{n}^{\dagger})E_{n}(U_{n}\Pi^{(n)}U_{n}^{\dagger})\big{\|}_{2}$ (48)
$\leq$ $\big{\|}\widetilde{\Pi}^{(n)}E_{n}\widetilde{\Pi}^{(n)}-\widetilde{\Pi}^{(n)}E_{n}(U_{n}\Pi^{(n)}U_{n}^{\dagger})\big{\|}_{2}+\big{\|}\widetilde{\Pi}^{(n)}E_{n}(U_{n}\Pi^{(n)}U_{n}^{\dagger})-(U_{n}\Pi^{(n)}U_{n}^{\dagger})E_{n}(U_{n}\Pi^{(n)}U_{n}^{\dagger})\big{\|}_{2}$ (49)
$\leq$ $\big{\|}\widetilde{\Pi}^{(n)}\big{\|}_{\infty}\big{\|}E_{n}\big{\|}_{\infty}\big{\|}\widetilde{\Pi}^{(n)}-U_{n}\Pi^{(n)}U_{n}^{\dagger}\big{\|}_{2}+\big{\|}\widetilde{\Pi}^{(n)}-U_{n}\Pi^{(n)}U_{n}^{\dagger}\big{\|}_{2}\big{\|}E_{n}\big{\|}_{\infty}\big{\|}U_{n}\Pi^{(n)}U_{n}^{\dagger}\big{\|}_{\infty}$ (50)
$=$ $2\big{\|}\widetilde{\Pi}^{(n)}-U_{n}\Pi^{(n)}U_{n}^{\dagger}\big{\|}_{2}=o(1).$ (51)
This implies
$\big{\|}c_{n}E_{n}\widetilde{\Pi}^{(n)}-\widetilde{\Pi}^{(n)}\big{\|}_{2}=\big{\|}c_{n}\widetilde{\Pi}^{(n)}-E_{n}\widetilde{\Pi}^{(n)}\big{\|}_{2}$ (52)
$\leq$ $\big{\|}c_{n}\widetilde{\Pi}^{(n)}-c_{n}U_{n}\Pi^{(n)}U_{n}^{\dagger}\big{\|}_{2}+\big{\|}c_{n}U_{n}\Pi^{(n)}U_{n}^{\dagger}-E_{n}\widetilde{\Pi}_{n}\big{\|}_{2}=o(1).$ (53)
However, this is impossible since $\operatorname{Tr}(cE_{n}\widetilde{\Pi}^{(n)})=0$ and $\operatorname{Tr}(\widetilde{\Pi}^{(n)})=2^{k}$. Therefore, we conclude that $\widetilde{d}_{n}\geq\lfloor\frac{d_{n}-1}{\eta}\rfloor+1$ when $n\in T$ is sufficiently large.
When $n\in T$ is sufficiently large, the stabilizer code family $\widetilde{\Pi}^{(n)}$ has properties:
  1. 1.
    It is defined on the same manifold as the original TSC $\Pi^{(n)}$.
  2. 2.
    It has local Pauli generators $P^{(n)}_{1},\cdots,P^{(n)}_{n-k}$, with diameters $\leq C$.
  3. 3.
    Its distance can be arbitrarily large when $n\rightarrow\infty$.
Therefore, $\{\widetilde{\Pi}^{(n)}\}$ is a TSC family.
Denote $\Pi^{(n)}=J_{1}^{(n)}J_{1}^{(n)\dagger}$ and $\widetilde{\Pi}^{(n)}=J_{2}^{(n)}J_{2}^{(n)\dagger}$, where $J_{1}^{(n)}$ and $J_{2}^{(n)}$ are Clifford encoding isometries. We have $|\overline{\phi}_{n}\rangle=J_{1}^{(n)}|\phi\rangle$. The constant depth local circuit $U_{n}$ maps the original TCS $\Pi^{(n)}$ to $U_{n}\Pi^{(n)}U_{n}^{\dagger}$, which is near the newly defined TSC $\widetilde{\Pi}^{(n)}$: $\big{\|}U_{n}\Pi^{(n)}U_{n}^{\dagger}-\widetilde{\Pi}^{(n)}\big{\|}_{2}=o(1)$. By Theorem 13, the robust form of Bravyi–König theorem, there exists a sequence of $k$-qubit unitaries $\{V^{(n)}\}\subset\mathcal{C}_{k}^{(D)}$ such that the induced linear transform $J_{2}^{(n)\dagger}U_{n}J_{1}^{(n)}$ (not a unitary in general) satisfies
$\big{\|}J_{2}^{(n)\dagger}U_{n}J_{1}^{(n)}-V^{(n)}\big{\|}_{2}=o(1).$ (54)
Therefore, we have
$\big{\|}J_{2}^{(n)\dagger}U_{n}J_{1}^{(n)}|\phi\rangle\langle\phi|\big{(}J_{2}^{(n)\dagger}U_{n}J_{1}^{(n)}\big{)}^{\dagger}-V^{(n)}|\phi\rangle\langle\phi|V^{(n)\dagger}\big{\|}_{1}$ (55)
$\leq$ $2\big{\|}J_{2}^{(n)\dagger}U_{n}J_{1}^{(n)}-V^{(n)}\big{\|}_{1}\leq 2\cdot 2^{k/2}\big{\|}J_{2}^{(n)\dagger}U_{n}J_{1}^{(n)}-V^{(n)}\big{\|}_{2}=o(1).$ (56)
Also,
$\big{\|}J_{2}^{(n)\dagger}U_{n}J_{1}^{(n)}|\phi\rangle\langle\phi|\big{(}J_{2}^{(n)\dagger}U_{n}J_{1}^{(n)}\big{)}^{\dagger}-J_{2}^{(n)\dagger}|S_{n}\rangle\langle S_{n}|J_{2}^{(n)}\big{\|}_{1}=\big{\|}U_{n}|\overline{\phi}_{n}\rangle\langle\overline{\phi}_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=o(1/n).$ (57)
Therefore,
$\big{\|}J_{2}^{(n)\dagger}|S_{n}\rangle\langle S_{n}|J_{2}^{(n)}-V^{(n)}|\phi\rangle\langle\phi|V^{(n)\dagger}\big{\|}_{1}\leq o(1)+o(1/n)=o(1).$ (58)
$J_{2}^{(n)\dagger}|S_{n}\rangle$ is a $k$-qubit stabilizer state. Note that $|\phi\rangle\notin\mathcal{C}_{k}^{(D)\dagger}|0^{k}\rangle:=\big{\{}W^{\dagger}|0^{k}\rangle\mid W\in\mathcal{C}_{k}^{(D)}\big{\}}$. From Ref. Zeng et al. (2008), we know $\mathcal{C}_{k}^{(D)\dagger}\mathcal{C}^{(2)}_{k}=\mathcal{C}_{k}^{(D)\dagger}$, thus $|\phi\rangle\notin\mathcal{C}_{k}^{(D)\dagger}|0^{k}\rangle=\mathcal{C}_{k}^{(D)\dagger}\mathcal{C}^{(2)}_{k}|0^{k}\rangle$, implying $V|\phi\rangle\notin\mathcal{C}^{(2)}_{k}|0^{k}\rangle$. Therefore, $V|\phi\rangle$ is a $k$-qubit nonstabilizer state, implying that Eq. (58) is impossible.
Finally, we can conclude that $\{|\overline{\phi}_{n}\rangle\}$ has strong LRM. ∎
Appendix B A robust form of the Bravyi–König theorem
In this section, we denote $\left\|A\right\|:=\left\|A\right\|_{2}=\sqrt{\operatorname{Tr}(A^{\dagger}A)}$ for a matrix $A$. For convenience, we define $\mathcal{C}_{k}^{(1)}:=\{e^{i\theta}P\mid\theta\in\mathbb{R},P\in\mathcal{P}_{k}\}$ in this section, this doesn’t change the definition of $\mathcal{C}_{k}^{(j)}$ for $j\geq 2$.
Lemma 10.
For square matrices $A,B_{1},\cdots,B_{m}$ of the same size, if for $j=1,\cdots,m$ we have $\big{\|}[A,B_{j}]\big{\|}\leq\epsilon_{j}$ and $\left\|B_{j}\right\|_{\infty}\leq 1$, then
$\big{\|}[A,B_{1}\cdots B_{m}]\big{\|}\leq\epsilon_{1}+\cdots+\epsilon_{m}.$ (59)
Proof.
We have
$\big{\|}[A,B_{1}\cdots B_{m}]\big{\|}=$ $\big{\|}AB_{1}\cdots B_{m}-B_{1}\cdots B_{m}A\big{\|}$ (60)
$\leq$ $\sum_{i=1}^{m}\big{\|}B_{1}\cdots B_{i-1}(AB_{i})B_{i+1}\cdots B_{m}-B_{1}\cdots B_{i-1}(B_{i}A)B_{i+1}\cdots B_{m}\big{\|}$ (61)
$=$ $\sum_{i=1}^{m}\big{\|}B_{1}\cdots B_{i-1}(AB_{i}-B_{i}A)B_{i+1}\cdots B_{m}\big{\|}$ (62)
$\leq$ $\sum_{i=1}^{m}\left\|B_{1}\right\|_{\infty}\cdots\left\|B_{i-1}\right\|_{\infty}\left\|AB_{i}-B_{i}A\right\|\left\|B_{i+1}\right\|_{\infty}\cdots\left\|B_{m}\right\|_{\infty}$ (63)
$\leq$ $\epsilon_{1}+\cdots+\epsilon_{m}.$ (64)
Lemma 11.
Suppose a unitary $M$ acts on $k$ qubits. If for some $\epsilon\geq 0$ sufficiently small,
$\max_{P\in\mathcal{G}_{k}}~{}\min\left\{\big{\|}[M,P]\big{\|}^{2},~{}\big{\|}\{M,P\}\big{\|}^{2}\right\}\leq\epsilon,$ (65)
where $\mathcal{G}_{k}:=\{X_{1},\cdots,X_{k},Z_{1},\cdots,Z_{k}\}$, then there exists a $Q\in\mathcal{C}_{k}^{(1)}$ such that
$\left\|M-Q\right\|^{2}\leq 2^{-1}k\epsilon+2^{-k-2}k^{2}\epsilon^{2}=O(\epsilon).$ (66)
Proof.
Decompose $M$ as $M=\sum_{R\in\mathcal{P}_{k}^{+}}c_{R}R$, where $\mathcal{P}_{k}^{+}$ denotes the set of $4^{k}$ $k$-qubit Pauli strings with phase $+1$, and $c_{R}=\frac{1}{2^{k}}\operatorname{Tr}(RM)\in\mathbb{C}$. Since $M$ is unitary, we have $\operatorname{Tr}(M^{\dagger}M)=\operatorname{Tr}(\mathbb{I}_{2^{k}})=2^{k}$, implying $\sum_{R\in\mathcal{P}_{k}^{+}}\left|c_{R}\right|^{2}=1$. For arbitrary $P\in\mathcal{P}_{k}$,
$[M,P]$ $=\sum_{R\in\mathcal{P}_{k}^{+}}c_{R}[R,P]=\sum_{R\in\mathcal{P}_{k}^{+},RP=-PR}2c_{R}RP,$ (67)
$\{M,P\}$ $=\sum_{R\in\mathcal{P}_{k}^{+}}c_{R}\{R,P\}=\sum_{R\in\mathcal{P}_{k}^{+},RP=PR}2c_{R}RP.$ (68)
Therefore we have
$\big{\|}[M,P]\big{\|}^{2}$ $=\operatorname{Tr}([M,P]^{\dagger}[M,P])=2^{k}\sum_{R\in\mathcal{P}_{k}^{+},RP=-PR}4\left|c_{R}\right|^{2},$ (69)
$\big{\|}\{M,P\}\big{\|}^{2}$ $=\operatorname{Tr}(\{M,P\}^{\dagger}\{M,P\})=2^{k}\sum_{R\in\mathcal{P}_{k}^{+},RP=PR}4\left|c_{R}\right|^{2}.$ (70)
Consider $Q^{\prime}\in\mathcal{P}_{k}^{+}$ defined as following: For $i=1,\cdots,k$,
  1. 1.
    If $\big{\|}[M,X_{i}]\big{\|}^{2}\leq\epsilon$ and $\big{\|}[M,Z_{i}]\big{\|}^{2}\leq\epsilon$, let $Q^{\prime}|_{i}=\mathbb{I}$,
  2. 2.
    If $\big{\|}[M,X_{i}]\big{\|}^{2}\leq\epsilon$ and $\big{\|}\{M,Z_{i}\}\big{\|}^{2}\leq\epsilon$, let $Q^{\prime}|_{i}=X$,
  3. 3.
    If $\big{\|}\{M,X_{i}\}\big{\|}^{2}\leq\epsilon$ and $\big{\|}[M,Z_{i}]\big{\|}^{2}\leq\epsilon$, let $Q^{\prime}|_{i}=Z$,
  4. 4.
    If $\big{\|}\{M,X_{i}\}\big{\|}^{2}\leq\epsilon$ and $\big{\|}\{M,Z_{i}\}\big{\|}^{2}\leq\epsilon$, let $Q^{\prime}|_{i}=Y$,
where $X_{i}$ is the $k$-qubit Pauli string which is $X$ on the $i$-th qubit and $\mathbb{I}$ elsewhere, and $Q^{\prime}|_{i}$ is the $i$-th Pauli operator of $Q^{\prime}$. For the case 2, $\big{\|}[M,X_{i}]\big{\|}^{2}\leq\epsilon$ and $\big{\|}\{M,Z_{i}\}\big{\|}^{2}\leq\epsilon$, we have
$\big{\|}[M,X_{i}]\big{\|}^{2}$ $=2^{k}\sum_{R\in\mathcal{P}_{k}^{+},RX_{i}=-X_{i}R}4\left|c_{R}\right|^{2}=2^{k+2}\sum_{R\in\mathcal{P}_{k}^{+},R|_{i}=Y,Z}\left|c_{R}\right|^{2}\leq\epsilon,$ (71)
$\big{\|}\{M,Z_{i}\}\big{\|}^{2}$ $=2^{k}\sum_{R\in\mathcal{P}_{k}^{+},RZ_{i}=Z_{i}R}4\left|c_{R}\right|^{2}=2^{k+2}\sum_{R\in\mathcal{P}_{k}^{+},R|_{i}=I,Z}\left|c_{R}\right|^{2}\leq\epsilon,$ (72)
thus $\sum_{R\in\mathcal{P}_{k}^{+},R|_{i}\neq X}\left|c_{R}\right|^{2}=\sum_{R\in\mathcal{P}_{k}^{+},R|_{i}=I,Y,Z}\left|c_{R}\right|^{2}\leq 2\cdot 2^{-k-2}\epsilon=2^{-k-1}\epsilon$. Other cases are similar, and we obtain
$\sum_{R\in\mathcal{P}_{k}^{+},R|_{i}\neq Q^{\prime}|_{i}}\left|c_{R}\right|^{2}\leq 2^{-k-1}\epsilon.$ (73)
Notice that $\{R\in\mathcal{P}_{k}^{+}\mid R\neq Q^{\prime}\}=\bigcup_{i=1}^{k}\big{\{}R\in\mathcal{P}_{k}^{+}\big{|}~{}R|_{i}\neq Q^{\prime}|_{i}\big{\}}$, thus
$\sum_{R\in\mathcal{P}_{k}^{+},R\neq Q^{\prime}}\left|c_{R}\right|^{2}\leq\sum_{i=1}^{k}\sum_{R\in\mathcal{P}_{k}^{+},R|_{i}\neq Q^{\prime}|_{i}}\left|c_{R}\right|^{2}\leq 2^{-k-1}k\epsilon.$ (74)
Therefore, $\left|c_{Q^{\prime}}\right|^{2}\geq 1-2^{-k-1}k\epsilon$. We write $c_{Q^{\prime}}=e^{i\theta}\left|c_{Q^{\prime}}\right|$, and take $Q=e^{i\theta}Q^{\prime}\in\mathcal{C}_{k}^{(1)}$. For $\epsilon\geq 0$ sufficiently small we $0\leq 1-2^{-k-1}k\epsilon\leq 1$, so $\left|c_{Q^{\prime}}\right|\geq\sqrt{1-2^{-k-1}k\epsilon}\geq 1-2^{-k-1}k\epsilon$, implying $\left|c_{Q^{\prime}}-e^{i\theta}\right|^{2}=(1-\left|c_{Q^{\prime}}\right|)^{2}\leq(2^{-k-1}k\epsilon)^{2}$. Since $M-Q=\sum_{R\in\mathcal{P}_{k}^{+}}(c_{R}-e^{i\theta}\delta_{R=Q^{\prime}})R$, we conclude that
$\left\|M-Q\right\|^{2}=$ $2^{k}\sum_{R\in\mathcal{P}_{k}^{+}}\left|c_{R}-e^{i\theta}\delta_{R=Q^{\prime}}\right|^{2}$ (75)
$=$ $2^{k}\sum_{R\in\mathcal{P}_{k}^{+},R\neq Q^{\prime}}\left|c_{R}\right|^{2}+2^{k}\left|c_{Q^{\prime}}-e^{i\theta}\right|^{2}$ (76)
$\leq$ $2^{-1}k\epsilon+2^{-k-2}k^{2}\epsilon^{2}.$ (77)
Lemma 12.
Suppose $M$ is a unitary matrix acting on $k$ qubits, and $j\geq 1$. If for some $\delta\geq 0$ sufficiently small, and for every $P\in\mathcal{P}_{k}$, there exists $A_{P}\in\mathcal{C}^{(j)}_{k}$ such that $\left\|MPM^{\dagger}-A_{P}\right\|\leq\delta$, then there exists $B\in\mathcal{C}^{(j+1)}_{k}$, such that
$\left\|M-B\right\|=O(\delta).$ (78)
Proof.
For arbitrary $P,Q\in\mathcal{P}_{k}$ we have
$\left\|A_{P}A_{Q}-MPQM^{\dagger}\right\|$ $\leq\left\|A_{P}A_{Q}-A_{P}MQM^{\dagger}\right\|+\left\|A_{P}MQM^{\dagger}-MPM^{\dagger}MQM^{\dagger}\right\|$ (79)
$\leq\big{\|}A_{P}\big{\|}_{\infty}\big{\|}A_{Q}-MQM^{\dagger}\big{\|}+\big{\|}A_{P}-MPM^{\dagger}\big{\|}\big{\|}MQM^{\dagger}\big{\|}_{\infty}\leq 2\delta.$ (80)
If $[P,Q]=0$, then
$\left\|A_{P}A_{Q}-A_{Q}A_{P}\right\|\leq\left\|A_{P}A_{Q}-MPQM^{\dagger}\right\|+\left\|MPQM^{\dagger}-MQPM^{\dagger}\right\|+\left\|MQPM^{\dagger}-A_{Q}A_{P}\right\|\leq 4\delta.$ (81)
If $\{P,Q\}=0$, then
$\big{\|}A_{P}A_{Q}+A_{Q}A_{P}\big{\|}\leq\left\|A_{P}A_{Q}-MPQM^{\dagger}\right\|+\left\|MPQM^{\dagger}+MQPM^{\dagger}\right\|+\left\|A_{Q}A_{P}-MQPM^{\dagger}\right\|\leq 4\delta.$ (82)
For all $j\geq 1$, $\mathcal{C}^{(j)}_{k}$ is finite up to global phases. Therefore, the sets
$\big{\{}\left\|C_{1}C_{2}-C_{2}C_{1}\right\|\mid C_{1},C_{2}\in\mathcal{C}^{(j)}_{k}\big{\}}\text{ and }\big{\{}\left\|C_{1}C_{2}+C_{2}C_{1}\right\|\mid C_{1},C_{2}\in\mathcal{C}^{(j)}_{k}\big{\}}$ (83)
are finite subsets of $\mathbb{R}$. Therefore, when $\delta$ is sufficiently small,
$[P,Q]$ $=0\Rightarrow\left\|A_{P}A_{Q}-A_{Q}A_{P}\right\|=0\iff[A_{P},A_{Q}]=0,$ (84)
$\{P,Q\}$ $=0\Rightarrow\left\|A_{P}A_{Q}+A_{Q}A_{P}\right\|=0\iff\{A_{P},A_{Q}\}=0.$ (85)
Denote $\mathcal{G}_{k}:=\{X_{1},\cdots,X_{k},Z_{1},\cdots,Z_{k}\}$, which generates $\mathcal{P}_{k}$ up to phases. For all $P\in\mathcal{G}_{k}$, $MPM^{\dagger}$ has eignevalues half $+1$ half $-1$. Since $\mathcal{C}^{(j)}_{k}$ is a set of unitaries which is finite up to global phases, when $\delta$ is sufficiently small, $\left\|MPM^{\dagger}-A_{P}\right\|\leq\delta$ ensures that $A_{P}\in\mathcal{C}^{(j)}_{k}$ must equal to some matrix $A_{P}^{\prime}\in\mathcal{C}^{(j)}_{k}$ which has eigenvalues half $+1$ half $-1$, up to a small global phase, otherwise $\left\|MPM^{\dagger}-A_{P}\right\|$ cannot be arbitrarily small. We write $A_{P}^{\prime}=e^{-i\theta_{P}}A_{P}$, with small $\left|\theta_{P}\right|$. $A_{P}^{\prime}$ is unitary and has real eigenvalues, thus it is Hermitian.
Since the eigenvalues of $A_{P}$ are half $e^{i\theta_{P}}$ half $-e^{i\theta_{P}}$, by Hoffman-Wielandt inequality Hoffman and Wielandt (2003) we know $\left\|MPM^{\dagger}-A_{P}\right\|^{2}\geq 2^{k}\left|1-e^{i\theta_{P}}\right|^{2}$, thus we have $\left|1-e^{i\theta_{P}}\right|\leq 2^{-k/2}\delta$.
Now we obtain $\{A_{P}^{\prime}\mid P\in\mathcal{G}_{k}\}\subset\mathcal{C}^{(j)}_{k}$, a set of “Pauli like” matrices, satisfying
  1. 1.
    For all $P,Q\in\mathcal{G}_{k}$, $[P,Q]=0\iff[A_{P}^{\prime},A_{Q}^{\prime}]=0$ , and $\{P,Q\}=0\iff\{A_{P}^{\prime},A_{Q}^{\prime}\}=0$;
  2. 2.
    For all $P\in\mathcal{G}_{k}$, $A_{P}^{\prime}$ is a Hermitian matrix satisfying $(A_{P}^{\prime})^{2}=\mathbb{I}$.
By these relations, we now show that there exists a unitary $B$ such that $BPB^{\dagger}=A_{P}^{\prime}$, for all $P\in\mathcal{G}_{k}$.
$\mathcal{G}_{k}$ is a minimal generating set for the matrix algebra $M_{n}(\mathbb{C})$ by multiplication and linear combination, with relations: I. $P^{2}=\mathbb{I},\forall P\in\mathcal{G}_{k}$; II. $[Z_{i},Z_{j}]=0,[X_{i},X_{j}]=0,\forall i,j\in[k]$; III. $[Z_{i},X_{j}]=0,\forall i\neq j$; IV. $\{Z_{i},X_{i}\}=0,\forall i\in[k]$. Since $\{A_{P}^{\prime}\mid P\in\mathcal{G}_{k}\}\subset\mathcal{C}^{(j)}_{k}\subset M_{n}(\mathbb{C})$ is another set of generators with the same relations, we know it also generates $M_{n}(\mathbb{C})$, and the map $\varphi:\mathcal{G}_{k}\rightarrow\{A_{P}^{\prime}\mid P\in\mathcal{G}_{k}\}$ defined by $\varphi(P)=A_{P}^{\prime}$ induces an automorphism $\varphi:M_{n}(\mathbb{C})\rightarrow M_{n}(\mathbb{C})$ of $M_{2^{k}}(\mathbb{C})$.
By the Skolem–Noether theorem (see e.g. Ref. Lorenz (2007)), we know every automorphism of the matrix algebra $M_{n}(\mathbb{C})$ is inner (i.e., has the form $\varphi(x)=u^{-1}xu$). Therefore there exists an invertible matrix $\widetilde{B}\in M_{n}(\mathbb{C})$, such that $\varphi(\cdot)=\widetilde{B}\cdot\widetilde{B}^{-1}$.
For all $P\in\mathcal{G}_{k}$, since $(A_{P}^{\prime})^{\dagger}=A_{P}^{\prime}$, we have
$(\widetilde{B}^{\dagger}\widetilde{B})P(\widetilde{B}^{\dagger}\widetilde{B})^{-1}=\widetilde{B}^{\dagger}(\widetilde{B}P\widetilde{B}^{-1})(\widetilde{B}^{\dagger})^{-1}=\widetilde{B}^{\dagger}A_{P}^{\prime}(\widetilde{B}^{\dagger})^{-1}=(\widetilde{B}^{-1}A_{P}^{\prime}\widetilde{B})^{\dagger}=P^{\dagger}=P.$ (86)
Therefore for all $P\in\mathcal{G}_{k}$ we have $(\widetilde{B}^{\dagger}\widetilde{B})P=P(\widetilde{B}^{\dagger}\widetilde{B})$, implying $\widetilde{B}^{\dagger}\widetilde{B}=b\mathbb{I}$ for some $b>0$. Define $B=b^{-1/2}\widetilde{B}$, we have $B^{\dagger}B=\mathbb{I}$, and $\varphi(\cdot)=\widetilde{B}\cdot\widetilde{B}^{-1}=B\cdot B^{-1}=B\cdot B^{\dagger}$. Therefore $BPB^{\dagger}=A_{P}^{\prime}$ for all $P\in\mathcal{G}_{k}$.
Now we show that $B\in\mathcal{C}^{(j+1)}_{k}$. For arbitrary $R\in\mathcal{P}_{k}$, write $R=e^{i\varphi}P_{1}\cdots P_{s}$ where $P_{1},\cdots,P_{s}\in\mathcal{G}_{k}$ and $\theta\in\mathbb{R}$. We have
$BRB^{\dagger}=e^{i\varphi}(BP_{1}B^{\dagger})\cdots(BP_{s}B^{\dagger})=e^{i\varphi}A_{P_{1}}^{\prime}\cdots A_{P_{s}}^{\prime}.$ (87)
There exists $A_{R}\in\mathcal{C}^{(j)}_{k}$ close to $MRM^{\dagger}$, thus
$\left\|BRB^{\dagger}-A_{R}\right\|\leq$ $\left\|BRB^{\dagger}-MRM^{\dagger}\right\|+\left\|MRM^{\dagger}-A_{R}\right\|$ (88)
$=$ $\left\|A_{P_{1}}^{\prime}\cdots A_{P_{s}}^{\prime}-A_{P_{1}}\cdots A_{P_{s}}\right\|+\left\|MRM^{\dagger}-A_{R}\right\|$ (89)
$\leq$ $\sum_{i=1}^{s}\left|1-e^{i\theta_{P_{i}}}\right|\left\|\mathbb{I}\right\|+\delta\leq k2^{-k/2}\delta 2^{k/2}+\delta=(k+1)\delta.$ (90)
Since $BRB^{\dagger}=e^{i\varphi}A_{P_{1}}^{\prime}\cdots A_{P_{s}}^{\prime}\in\big{(}\mathcal{C}^{(j)}_{k}\big{)}^{s}$ and $A_{R}\in\mathcal{C}^{(j)}_{k}$, where $\big{(}\mathcal{C}^{(j)}_{k}\big{)}^{s}=\mathcal{C}^{(j)}_{k}\cdots\mathcal{C}^{(j)}_{k}$ contains $\mathcal{C}^{(j)}_{k}$ and is a finite set up to global phases, for sufficiently small $\delta$, $BRB^{\dagger}$ must equal to $A_{R}$ up to global phases. By $A_{R}\in\mathcal{C}^{(j)}_{k}$ we know $BRB^{\dagger}\in\mathcal{C}^{(j)}_{k}$. Therefore we have $B\mathcal{P}_{k}B^{\dagger}\subset\mathcal{C}^{(j)}_{k}$, that is, $B\in\mathcal{C}^{(j+1)}_{k}$.
Now we upper bound the distance between $M$ and $B$. For any $P\in\mathcal{G}_{k}$,
$\big{\|}(M^{\dagger}B)P-P(M^{\dagger}B)\big{\|}=\big{\|}BPB^{\dagger}-MPM^{\dagger}\big{\|}=\big{\|}A_{P}^{\prime}-MPM^{\dagger}\big{\|}$ (91)
$\leq\big{\|}A_{P}^{\prime}-A_{P}\big{\|}+\big{\|}A_{P}-MPM^{\dagger}\big{\|}\leq\delta+\delta=2\delta.$ (92)
By Lemma 11, we conclude that
$\big{\|}B-M\big{\|}=\big{\|}M^{\dagger}B-\mathbb{I}\big{\|}=O(\delta).$ (93)
Theorem 13 (Robust Bravyi–König).
Suppose two families of $D$-dimensional ($D\geq 2$) TSCs are defined on the same lattice $\Lambda^{(n)}$, both have $k$ logical qubits, with code projectors $\Pi_{1}^{(n)}=J_{1}^{(n)}J_{1}^{(n)\dagger}$ and $\Pi_{2}^{(n)}=J_{2}^{(n)}J_{2}^{(n)\dagger}$ respectively, where $n$ is the number of physical qubits and $J_{1}^{(n)},J_{2}^{(n)}$ are Clifford encoding isometries. If a family of constant-depth local unitaries $\{U^{(n)}\}$ satisfies $\big{\|}U^{(n)}\Pi_{1}^{(n)}U^{(n)\dagger}-\Pi_{2}^{(n)}\big{\|}_{2}=o(1)$, then there exist $\{Q^{(n)}\}\subset\mathcal{C}_{k}^{(D)}$ such that the induced linear transform $J_{2}^{(n)\dagger}U^{(n)}J_{1}^{(n)}$ satisfies $\big{\|}J_{2}^{(n)\dagger}U^{(n)}J_{1}^{(n)}-Q^{(n)}\big{\|}_{2}=o(1)$.
Proof.
We omit superscripts $(n)$ for brevity and implicitly consider these families in the following proof.
As in Ref. Bravyi and König (2013) (arXiv version, page 5), we partition the lattice $\Lambda$ into $D+1$ regions $\Lambda=\bigcup_{j=1}^{D+1}\Lambda_{j}$. $\Lambda_{j}$ corresponds to $(j-1)$-dimensonal simplices in a triangulation of $\Lambda$. Each region $\Lambda_{j}$ is a disjoint union of chunks of size $O(1)$ separated by distance $\Omega(1)$. Each region $\Lambda_{j}$ is correctable.
As in Ref. Bravyi and König (2013), take $n$-qubit Pauli strings $P_{1},P_{2},\cdots,P_{D}$, such that $[P_{1},\Pi_{1}]=0$ and $[P_{i},\Pi_{2}]=0$ for $i=2,\cdots,D$. We can ensure that $\operatorname{supp}(P_{j})$ does not overlap with $\Lambda_{j}$ together with its $\rho$-neighborhood for some constant $\rho$. We can take $P_{1},P_{2},\cdots,P_{D}$ such that $J_{1}^{\dagger}P_{1}J_{1}$ and $J_{2}^{\dagger}P_{2}J_{2},\cdots,J_{2}^{\dagger}P_{D}J_{2}$ be arbitrary elements in $\mathcal{P}_{k}$.
Define $K_{1}=UP_{1}U^{\dagger}$, and $K_{j}=P_{j}^{\dagger}K_{j-1}P_{j}K_{j-1}^{\dagger}$ for $j=2,\cdots,D$. Same as the proof in Ref. Bravyi and König (2013), we know $K_{j}$ acts trivially on $\bigcup_{i=1}^{j}\Lambda_{i}$. Therefore, $K_{D}$ is supported on the correctable region $\Lambda_{D+1}$.
Since $\big{\|}\Pi_{1}U^{\dagger}-U^{\dagger}\Pi_{2}\big{\|}=\big{\|}U\Pi_{1}-\Pi_{2}U\big{\|}=\big{\|}U\Pi_{1}U^{\dagger}-\Pi_{2}\big{\|}=o(1)$, we know that
$\big{\|}[\Pi_{2},K_{1}]\big{\|}=$ $\left\|\Pi_{2}UP_{1}U^{\dagger}-UP_{1}U^{\dagger}\Pi_{2}\right\|$ (94)
$\leq$ $\left\|\Pi_{2}UP_{1}U^{\dagger}-U\Pi_{1}P_{1}U^{\dagger}\right\|+\left\|UP_{1}\Pi_{1}U^{\dagger}-UP_{1}U^{\dagger}\Pi_{2}\right\|$ (95)
$\leq$ $\big{\|}\Pi_{2}U-U\Pi_{1}\big{\|}\big{\|}P_{1}U^{\dagger}\big{\|}_{\infty}+\big{\|}UP_{1}\big{\|}_{\infty}\big{\|}\Pi_{1}U^{\dagger}-U^{\dagger}\Pi_{2}\big{\|}=o(1).$ (96)
For $j=2,\cdots,D$, there are $2^{j-1}$ $K_{1}$’s in $K_{j}$, by Lemma 10,
$\big{\|}[K_{j},\Pi_{2}]\big{\|}\leq 2^{j-1}\big{\|}[K_{1},\Pi_{2}]\big{\|}=o(1).$ (97)
Since $K_{D}$ is a correctable error, there exists $c\in\mathbb{C}$ with $\left|c\right|\leq 1$ such that $\Pi_{2}K_{D}\Pi_{2}=c\Pi_{2}$. Now we show that $c\approx\pm 1$ and $K_{D}\Pi_{2}\approx\pm\Pi_{2}$. We have
$\left\|K_{D}\Pi_{2}-c\Pi_{2}\right\|=$ $\left\|K_{D}\Pi_{2}\Pi_{2}-\Pi_{2}K_{D}\Pi_{2}\right\|\leq\left\|K_{D}\Pi_{2}-\Pi_{2}K_{D}\right\|\left\|\Pi_{2}\right\|_{\infty}=o(1),$ (98)
$\left\|\Pi_{2}K_{D}-c\Pi_{2}\right\|=$ $\left\|\Pi_{2}\Pi_{2}K_{D}-\Pi_{2}K_{D}\Pi_{2}\right\|\leq\left\|\Pi_{2}\right\|_{\infty}\left\|\Pi_{2}K_{D}-K_{D}\Pi_{2}\right\|=o(1).$ (99)
By $P_{D}^{2}=e^{i\theta}\mathbb{I}$, we have
$\left\|\Pi_{2}-c^{2}\Pi_{2}\right\|=\left\|(\Pi_{2}K_{D-1}P_{D}K_{D-1}^{\dagger})(K_{D-1}P_{D}K_{D-1}^{\dagger}\Pi_{2})-(cP_{D}\Pi_{2})(cP_{D}\Pi_{2})\right\|$ (100)
$\leq$ $\big{\|}\Pi_{2}K_{D-1}P_{D}K_{D-1}^{\dagger}-cP_{D}\Pi_{2}\big{\|}\big{\|}K_{D-1}P_{D}K_{D-1}^{\dagger}\Pi_{2}\big{\|}_{\infty}+\big{\|}cP_{D}\Pi_{2}\big{\|}_{\infty}\big{\|}K_{D-1}P_{D}K_{D-1}^{\dagger}\Pi_{2}-cP_{D}\Pi_{2}\big{\|}$ (101)
$\leq$ $\big{\|}\Pi_{2}K_{D-1}P_{D}K_{D-1}^{\dagger}-cP_{D}\Pi_{2}\big{\|}+\left|c\right|\big{\|}K_{D-1}P_{D}K_{D-1}^{\dagger}\Pi_{2}-cP_{D}\Pi_{2}\big{\|}$ (102)
$=$ $\left\|\Pi_{2}K_{D}-c\Pi_{2}\right\|+\left|c\right|\left\|K_{D}\Pi_{2}-c\Pi_{2}\right\|=o(1).$ (103)
Therefore $\left|1-c^{2}\right|\leq\left\|\Pi_{2}\right\|^{-1}o(1)=o(1)$, implying either $\left|c-1\right|=o(1)$ or $\left|c+1\right|=o(1)$, and
either $\text{either }\left\|K_{D}\Pi_{2}-\Pi_{2}\right\|=o(1)\text{ or }\left\|K_{D}\Pi_{2}+\Pi_{2}\right\|=o(1).$ (104)
We write this as $\left\|K_{D}\Pi_{2}\pm\Pi_{2}\right\|=o(1)$. By $\Pi_{2}=J_{2}J_{2}^{\dagger}$ we know this is equivalent to
$\big{\|}J_{2}^{\dagger}K_{D}J_{2}\pm\mathbb{I}\big{\|}=o(1).$ (105)
Denote $\hat{U}=J_{2}^{\dagger}UJ_{1}$, $\hat{P}_{1}=J_{1}^{\dagger}P_{1}J_{1}$, and $\hat{P}_{i}=J_{2}^{\dagger}P_{i}J_{2}$ for $i=2,\cdots,D$. Note that $\hat{P}_{1},\cdots,\hat{P}_{D}\in\mathcal{P}_{k}$, while $\hat{U}$ is a linear transform on $k$ qubits that is not a unitary in general. Also define $k$-qubit linear transformations $\hat{K}_{1}=\hat{U}\hat{P}_{1}\hat{U}^{\dagger}$, and $\hat{K}_{j}=\hat{P}_{j}^{\dagger}\hat{K}_{j-1}\hat{P}_{j}\hat{K}_{j-1}^{\dagger}$ for $j=2,\cdots,D$. They are not unitaries in general.
Now we show that $\big{\|}\hat{K}_{D}-J_{2}^{\dagger}K_{D}J_{2}\big{\|}=o(1)$. We have
$\big{\|}\hat{K}_{1}-J_{2}^{\dagger}K_{1}J_{2}\big{\|}=$ $\big{\|}J_{2}^{\dagger}UJ_{1}J_{1}^{\dagger}P_{1}J_{1}J_{1}^{\dagger}U^{\dagger}J_{2}-J_{2}^{\dagger}UP_{1}U^{\dagger}J_{2}\big{\|}$ (106)
$=$ $\big{\|}J_{2}^{\dagger}(U\Pi_{1})P_{1}(\Pi_{1}U^{\dagger})J_{2}-J_{2}^{\dagger}(\Pi_{2}U)P_{1}(U^{\dagger}\Pi_{2})J_{2}\big{\|}$ (107)
$\leq$ $2\big{\|}U\Pi_{1}-\Pi_{2}U\big{\|}=o(1)$ (108)
Suppose for $i\geq 1$ we have $\big{\|}\hat{K}_{i}-J_{2}^{\dagger}K_{i}J_{2}\big{\|}=o(1)$, then
$\big{\|}\hat{K}_{i+1}-J_{2}^{\dagger}K_{i+1}J_{2}\big{\|}=$ $\big{\|}J_{2}^{\dagger}P_{i+1}^{\dagger}J_{2}\hat{K}_{i}J_{2}^{\dagger}P_{i+1}J_{2}\hat{K}_{i}^{\dagger}-J_{2}^{\dagger}P_{i+1}^{\dagger}K_{i}P_{i+1}K_{i}^{\dagger}J_{2}\big{\|}$ (109)
$\leq$ $\big{\|}J_{2}^{\dagger}P_{i+1}^{\dagger}J_{2}\hat{K}_{i}J_{2}^{\dagger}P_{i+1}J_{2}\hat{K}_{i}^{\dagger}-J_{2}^{\dagger}P_{i+1}^{\dagger}J_{2}(J_{2}^{\dagger}K_{i}J_{2})J_{2}^{\dagger}P_{i+1}J_{2}(J_{2}^{\dagger}K_{i}^{\dagger}J_{2})\big{\|}$ (110)
$+\big{\|}J_{2}^{\dagger}P_{i+1}^{\dagger}J_{2}(J_{2}^{\dagger}K_{i}J_{2})J_{2}^{\dagger}P_{i+1}J_{2}(J_{2}^{\dagger}K_{i}^{\dagger}J_{2})-J_{2}^{\dagger}P_{i+1}^{\dagger}K_{i}P_{i+1}K_{i}^{\dagger}J_{2}\big{\|}$ (111)
$\leq$ $2\big{\|}\hat{K}_{i}-J_{2}^{\dagger}K_{i}J_{2}\big{\|}+\big{\|}J_{2}^{\dagger}P_{i+1}^{\dagger}K_{i}P_{i+1}(\Pi_{2}K_{i}^{\dagger})J_{2}-J_{2}^{\dagger}P_{i+1}^{\dagger}K_{i}P_{i+1}(K_{i}^{\dagger}\Pi_{2})J_{2}\big{\|}$ (112)
$\leq$ $2\big{\|}\hat{K}_{i}-J_{2}^{\dagger}K_{i}J_{2}\big{\|}+\big{\|}\Pi_{2}K_{i}^{\dagger}-K_{i}^{\dagger}\Pi_{2}\big{\|}=o(1).$ (113)
Therefore, we get $\big{\|}\hat{K}_{D}-J_{2}^{\dagger}K_{D}J_{2}\big{\|}=o(1)$ by induction, implying
$\big{\|}\hat{K}_{D}\pm\mathbb{I}\big{\|}=o(1).$ (114)
We have
$\big{\|}\hat{U}\hat{U}^{\dagger}-\mathbb{I}\big{\|}=\big{\|}J_{2}^{\dagger}(U\Pi_{1})U^{\dagger}J_{2}-J_{2}^{\dagger}(\Pi_{2}U)U^{\dagger}J_{2}\big{\|}\leq\big{\|}U\Pi_{1}-\Pi_{2}U\big{\|}=o(1),$ (115)
thus there exists a unitary matrix $\hat{U}^{\prime}$ such that $\|\hat{U}-\hat{U}^{\prime}\|=o(1)$. For $i=1,\cdots,D$, we define $\hat{K}_{i}^{\prime}$ by replacing $\hat{U}$ (or $\hat{U}^{\dagger}$) in $\hat{K}_{i}$ by $\hat{U}^{\prime}$ (or $\hat{U}^{\prime\dagger}$), thus $\hat{K}_{i}^{\prime},i=1,\cdots,D$ are unitaries. We have $\|\hat{K}_{i}-\hat{K}_{i}^{\prime}\|=o(1)$ for all $i=1,\cdots,D$.
By $\big{\|}\hat{K}_{D}\pm\mathbb{I}\big{\|}=o(1)$ implies $\big{\|}\hat{K}_{D}^{\prime}\pm\mathbb{I}\big{\|}=o(1)$, thus
$\big{\|}\hat{P}_{D}^{\dagger}\hat{K}_{D-1}^{\prime}\pm\hat{K}_{D-1}^{\prime}\hat{P}_{D}^{\dagger}\big{\|}=\big{\|}\hat{K}_{D}^{\prime}\pm\mathbb{I}\big{\|}=o(1).$ (116)
Since $\hat{P}_{D}$ can be an arbitrary element in $\mathcal{P}_{k}$, by Lemma 11 we know there exists $Q_{1}\in\mathcal{C}_{k}^{(1)}$ such that
$\big{\|}\hat{K}_{D-1}^{\prime}-Q_{1}\big{\|}=o(1).$ (117)
Now for $21\leq j\leq D-$ 2, suppose we have shown that there exists $Q_{j}\in\mathcal{C}_{k}^{(j)}$ such that $\big{\|}\hat{K}_{D-j}^{\prime}-Q_{j}\big{\|}=o(1)$, then
$\big{\|}\hat{K}_{D-j-1}^{\prime}\hat{P}_{D-j}\hat{K}_{D-j-1}^{\prime\dagger}-\hat{P}_{D-j}Q_{j}\big{\|}=\big{\|}\hat{K}_{D-j}^{\prime}-Q_{j}\big{\|}=o(1).$ (118)
Since $\hat{P}_{D-j}$ can be an arbitrary element in $\mathcal{P}_{k}$, and $\hat{P}_{D-j}Q_{j}\in\mathcal{C}_{k}^{(1)}\mathcal{C}_{k}^{(j)}=\mathcal{C}_{k}^{(j)}$, by Lemma 12 we know there exists $Q_{j+1}\in\mathcal{C}_{k}^{(j+1)}$ such that
$\big{\|}\hat{K}_{D-j-1}^{\prime}-Q_{j+1}\big{\|}=o(1).$ (119)
In particular, there exists $Q_{D-1}\in\mathcal{C}_{k}^{(D-1)}$ such that
$\big{\|}\hat{U}^{\prime}\hat{P}_{1}\hat{U}^{\prime\dagger}-Q_{D-1}\big{\|}=\big{\|}\hat{K}_{1}^{\prime}-Q_{D-1}\big{\|}=o(1).$ (120)
Since $\hat{P}_{1}$ can be arbitrary element in $\mathcal{P}_{k}$, by Lemma 12 we know that there exists $Q_{D}\in\mathcal{C}_{k}^{(D)}$ such that $\big{\|}\hat{U}^{\prime}-Q_{D}\big{\|}=o(1)$. Together with $\|\hat{U}-\hat{U}^{\prime}\|=o(1)$ we know
$\big{\|}J_{2}^{\dagger}UJ_{1}-Q_{D}\big{\|}=\big{\|}\hat{U}-Q_{D}\big{\|}=o(1).$ (121)
Appendix C Topological order and LRM phases
Theorem 14.
The following statements hold:
  1. 1.
    A topological order is an SRM phase iff it can be realized by a TSC.
  2. 2.
    A topological order is an LRM phase iff it cannot be realized by any TSC.
  3. 3.
    If a topological order cannot be realized approximately by any TSC up to $o(1)$ error in projector’s distance, then it is a strong LRM phase.
Proof.
Statement 1.
($\Leftarrow$) Since TCS has a family of ground stabilizer states, we know this topological order is an SRM phase.
($\Rightarrow$) Suppose this topological order is defined on a $D$-dimensional manifold $(D\geq 2)$. Since this topological order is an SRM phase, we know it can be realized by a family of local Hamiltonians $\{H^{(n)}=\sum_{i\in I^{(n)}}H_{i}^{(n)}\}$ acting on $n$ qubits located on this manifold, with a family of stabilizer ground states $\{|S^{(n)}\rangle\}$.
Denote $C$ as a constant independent of $n$ that upper bounds the diameters of local terms $\{H_{i}^{(n)}\}$. Let $\mathcal{A}_{n}$ be the set of states that are $C$-locally indistinguishable from $|S^{(n)}\rangle$, defined as
$\mathcal{A}_{n}:=\Big{\{}|\psi\rangle\Big{|}\operatorname{Tr}_{\overline{J}}|\psi\rangle\langle\psi|=\operatorname{Tr}_{\overline{J}}|S^{(n)}\rangle\langle S^{(n)}|,\forall J\subset[n]\text{ with }\operatorname{Diam}(J)\leq C\Big{\}}.$ (122)
We show that $\mathcal{A}_{n}$ is exactly the set of pure ground states of $H^{(n)}$. On one hand, for an arbitrary ground state $|\phi\rangle$ of $H^{(n)}$, by the TQO condition we know the reduced density matrices are the same locally, implying $|\phi\rangle\in\mathcal{A}_{n}$. On the other hand, for any $|a\rangle\in\mathcal{A}_{n}$, by Lemma 7 we know $\langle a|H_{i}^{(n)}|a\rangle=\langle S^{(n)}|H_{i}^{(n)}|S^{(n)}\rangle$. Therefore, $\langle a|H^{(n)}|a\rangle=\langle S^{(n)}|H^{(n)}|S^{(n)}\rangle$, and we have
$\mathcal{A}_{n}=\{\text{pure ground states of }H^{(n)}\}.$ (123)
Let $r_{n}$ be the maximum number of independent Pauli strings in $\{h\in G_{n}\mid\operatorname{Diam}(h)\leq C\}$, where $G_{n}$ is the stabilizer group of $|S^{(n)}\rangle$. Pick such an independent set $\{P_{1}^{(n)},\cdots,P_{r_{n}}^{(n)}\}\subset\{h\in G_{n}\mid\operatorname{Diam}(h)\leq C\}$, we must have
$\{h\in G_{n}\mid\operatorname{Diam}(h)\leq C\}\subset\langle P_{1}^{(n)},\cdots,P_{r_{n}}^{(n)}\rangle,$ (124)
since otherwise we can find $r_{n}+1$ independent Pauli strings in $\{h\in G_{n}\mid\operatorname{Diam}(h)\leq C\}$. Let
$\mathcal{A}^{\prime}_{n}:=\{|\psi\rangle\mid P_{i}^{(n)}|\psi\rangle=|\psi\rangle,i=1,\cdots,r_{n}\}.$ (125)
Now we show that $\mathcal{A}_{n}^{\prime}=\mathcal{A}_{n}$. For $|a\rangle\in\mathcal{A}_{n}$, we have $\langle a|P_{i}^{(n)}|a\rangle=\langle S^{(n)}|P_{i}^{(n)}|S^{(n)}\rangle=1$, that is, $P_{i}^{(n)}|a\rangle=|a\rangle$, thus $\mathcal{A}_{n}\subset\mathcal{A}^{\prime}_{n}$. We take $\{P_{r_{n}+1}^{(n)},\cdots,P_{n}^{(n)}\}\subset G$ such that $\{P_{1}^{(n)},\cdots,P_{n}^{(n)}\}$ are independent, so $\langle P_{1}^{(n)},\cdots,P_{n}^{(n)}\rangle=G_{n}$. For $j=r+1,\cdots,n$, since $P_{j}^{(n)}\notin\langle P_{1}^{(n)},\cdots,P_{r_{n}}^{(n)}\rangle$, we have $\operatorname{Diam}(P_{j}^{(n)})>C$. For $\mathbf{b}=(b_{r_{n}+1},\cdots,b_{n})\in\{0,1\}^{n-r_{n}}$, consider the states
$|\psi_{\mathbf{b}}^{(n)}\rangle\langle\psi_{\mathbf{b}}^{(n)}|$ $:=\prod_{i=1}^{r_{n}}\frac{\mathbb{I}+P_{i}^{(n)}}{2}\prod_{j=r_{n}+1}^{n}\frac{\mathbb{I}+(-1)^{b_{j}}P_{j}^{(n)}}{2}.$ (126)
They are orthogonal to each other. In particular, we have $|S^{(n)}\rangle=|\psi_{\mathbf{0}}^{(n)}\rangle$. Following the same procedure as in the proof of Theorem 9, we obtain $\{|\psi_{\mathbf{b}}^{(n)}\rangle\mid\mathbf{b}\in\{0,1\}^{n-r_{n}}\}\subset\mathcal{A}_{n}$, which means $\dim(\operatorname{span}(\mathcal{A}_{n}))\geq 2^{n-r_{n}}$. Note that $\dim(\operatorname{span}(\mathcal{A}^{\prime}_{n}))=2^{n-r_{n}}$, and $\operatorname{span}(\mathcal{A}_{n})\subset\operatorname{span}(\mathcal{A}^{\prime}_{n})$. We deduce that $\operatorname{span}(\mathcal{A})=\operatorname{span}(\mathcal{A}^{\prime}_{n})$. By Eq. (123) and Eq. (125) we know $\mathcal{A}_{n}$ and $\mathcal{A}^{\prime}_{n}$ are the sets of unit vectors in $\operatorname{span}(\mathcal{A}_{n})$ and $\operatorname{span}(\mathcal{A}^{\prime}_{n})$ respectively, which implies $\mathcal{A}_{n}=\mathcal{A}^{\prime}_{n}$.
Therefore, we know that the ground space of $H^{(n)}$ must have degeneracy $2^{k}$ for some integer $k\geq 1$, and $r_{n}=n-k$ is a constant independent of $n$.
To conclude, $\{P_{1}^{(n)},\cdots,P_{n-k}^{(n)}\}$ gives a TSC, since
  1. 1.
    It is defined on the same manifold as the Hamiltonian $H^{(n)}$.
  2. 2.
    Pauli generators $P_{1}^{(n)},\cdots,P_{n-k}^{(n)}$ have diameters $\leq C$.
  3. 3.
    It has macroscopic distance.
Therefore, this topological order can be realized by a TSC.
Statement 2.
Suppose a topological order is not an LRM phase, then there exists a family of local Hamiltonians $\{H^{(n)}=\sum_{i\in I^{(n)}}H_{i}^{(n)}\}$ acting on $n$ qubits that realizes of this topological order, and a family of ground states $\{|\phi^{(n)}\rangle=U^{(n)}|S^{(n)}\rangle\}_{n\in T}$, $T\subset\mathbb{Z}_{>0}$ exhibiting SRM. Since a constant-depth local circuit does not change phase, we know local Hamiltonians $\{U^{(n)}H^{(n)}U^{(n)\dagger}\}_{n\in T}$ also realize this topological order. Since $\{U^{(n)}H^{(n)}U^{(n)\dagger}\}_{n\in T}$ has a family of ground stabilizer state $\{|S^{(n)}\rangle\}_{n\in T}$, we know this topological order is an SRM phase.
Therefore, we get: A topological order is not an LRM phase $\iff$ it is an SRM phase.
Taking “not” on two sides of Statement 1 gets Statement 2.
Statement 3.
Implicit in the proof of Theorem 9.
C.1 Strong LRM from dimension mismatch
Theorem 15.
Suppose a topological order is realized by a family of local Hamiltonians $\{H^{(n)}=\sum_{i\in I^{(n)}}H_{i}^{(n)}\}$ acting on $n$ qubits located on a $D$-dimensional manifold $(D\geq 2)$. Suppose $|I^{(n)}|=\Theta(n)$, $\|H_{i}^{(n)}\|_{\infty}\leq 1$. If the dimension of the ground space of $\{H^{(n)}\}$ is not a power of 2, then any family of ground states $\big{\{}|\phi_{n}\rangle\big{\}}$ of $\{H^{(n)}\}$ has strong LRM.
Proof.
Assume for contradiction that there exists a subsequence $T\subset\mathbb{Z}_{>0}$, a family of ground states $\{|\phi_{n}\rangle\}_{n\in T}$, a family of constant-depth local circuits $\{U_{n}\}_{n\in T}$, and a family of stabilizer states $\{|S_{n}\rangle\}_{n\in T}$, such that for $n\in T$,
$\big{\|}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=o(1/n).$ (127)
Denote $C$ as a constant independent of $n$ that upper bounds the diameters of all deformed local terms $UH_{i}^{(n)}U^{\dagger}$. Define $\mathcal{B}_{n}$ as the set of states that are $C$-locally indistinguishable from $U_{n}|\phi_{n}\rangle$,
$\mathcal{B}_{n}:=$$\Big{\{}|\psi\rangle\Big{|}\operatorname{Tr}_{\overline{J}}|\psi\rangle\langle\psi|=\operatorname{Tr}_{\overline{J}}(U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}),\forall J\subset[n]\text{ with }\operatorname{Diam}(J)\leq C\Big{\}}$ (128)
[…]
where the second equality follows from the same arguments as the proof of Proposition LABEL:prop:TopologicalOrder_SRM_then_TSC.
Let $r_{n}$ be the maximum number of independent Pauli strings in $\{h\in G_{n}\mid\operatorname{Diam}(h)\leq C\}$, where $G_{n}$ is the stabilizer group of $|S_{n}\rangle$. Pick such an independent set $\{P_{1}^{(n)},\cdots,P_{r_{n}}^{(n)}\}\subset\{h\in G_{n}\mid\operatorname{Diam}(h)\leq C\}$. Let
$\mathcal{A}_{n}:=\Big{\{}|\psi\rangle\Big{|}\operatorname{Tr}_{\overline{J}}|\psi\rangle\langle\psi|=\operatorname{Tr}_{\overline{J}}|S_{n}\rangle\langle S_{n}|,\forall J\subset[n]\text{ with }\operatorname{Diam}(J)\leq C\Big{\}},$ (130)
and
$\mathcal{A}_{n}^{\prime}:=$$\big{\{}|\psi\rangle\mid P_{j}^{(n)}|\psi\rangle=|\psi\rangle,j=1,\cdots,r_{n}\big{\}}$ (131)
$=$ $\Big{\{}\text{pure ground states of }\widetilde{H}^{(n)}:=-\sum_{j=1}^{r_{n}}P_{j}^{(n)}\Big{\}}.$ (132)
By the proof of Theorem 9, we know $\mathcal{A}_{n}\subset\mathcal{A}_{n}^{\prime}$, and $\mathcal{A}_{n}$ contains a set of $2^{n-r_{n}}$ orthonormal pure states $\{|\psi^{(n)}_{\mathbf{b}}\rangle\mid\mathbf{b}\in\{0,1\}^{n-r_{n}}\}$, implying $\dim(\operatorname{span}(\mathcal{A}_{n}))=\dim(\operatorname{span}(\mathcal{A}_{n}^{\prime}))=2^{n-r_{n}}$.
We claim that $\dim(\operatorname{span}(\mathcal{B}_{n}))=2^{n-r_{n}}$ when $n\in T$ is sufficiently large. For all $\mathbf{b}\in\{0,1\}^{n-r_{n}}$, since $|\psi^{(n)}_{\mathbf{b}}\rangle\in\mathcal{A}_{n}$
$\left|\langle\psi^{(n)}_{\mathbf{b}}|U_{n}H^{(n)}U_{n}^{\dagger}|\psi^{(n)}_{\mathbf{b}}\rangle-\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle\right|=$ $\left|\langle S_{n}|U_{n}H^{(n)}U_{n}^{\dagger}|S_{n}\rangle-\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle\right|$ (133)
$=$ $\left|\operatorname{Tr}\big{[}H^{(n)}\big{(}U_{n}^{\dagger}|S_{n}\rangle\langle S_{n}|U_{n}-|\phi_{n}\rangle\langle\phi_{n}|\big{)}\big{]}\right|$ (134)
$\leq$ $\sum_{i\in I^{(n)}}\left|\operatorname{Tr}\big{[}H_{i}^{(n)}\big{(}U_{n}^{\dagger}|S_{n}\rangle\langle S_{n}|U_{n}-|\phi_{n}\rangle\langle\phi_{n}|\big{)}\big{]}\right|$ (135)
$\leq$ $\sum_{i\in I^{(n)}}\big{\|}H_{i}^{(n)}\big{\|}_{\infty}\big{\|}|S_{n}\rangle\langle S_{n}|-U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}\big{\|}_{1}$ (136)
$\leq$ $|I^{(n)}|o(1/n)=o(1).$ (137)
Therefore, for arbitrary $|\psi^{(n)}_{\mathbf{b}}\rangle$ we have
$\langle a|U_{n}H^{(n)}U_{n}^{\dagger}|a\rangle\leq\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle+o(1)=\lambda_{\min}(U_{n}H^{(n)}U_{n}^{\dagger})+o(1).$ (138)
Since $\{U_{n}H^{(n)}U_{n}^{\dagger}\}$ is topological ordered and thus gapped, by Lemma 8 we know $2^{n-r_{n}}\leq\dim(\operatorname{span}(\mathcal{B}_{n}))$ when $n\in T$ is sufficiently large. For arbitrary $|b\rangle\in\mathcal{B}_{n}$,
$\left|\langle b|\widetilde{H}^{(n)}|b\rangle-\langle S_{n}|\widetilde{H}^{(n)}|S_{n}\rangle\right|=$ $\left|\langle\phi_{n}|U_{n}^{\dagger}\widetilde{H}^{(n)}U_{n}|\phi_{n}\rangle-\langle S_{n}|\widetilde{H}^{(n)}|S_{n}\rangle\right|$ (139)
$=$ $\left|\operatorname{Tr}\big{[}\widetilde{H}^{(n)}\big{(}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{)}\big{]}\right|$ (140)
$\leq$ $\sum_{j=1}^{r_{n}}\left|\operatorname{Tr}\big{[}P_{j}^{(n)}\big{(}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{)}\big{]}\right|$ (141)
$\leq$ $\sum_{j=1}^{r_{n}}\big{\|}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}$ (142)
$\leq$ $r_{n}o(1/n)=o(1).$ (143)
Therefore, for arbitrary $|b\rangle\in\mathcal{B}_{n}$ we have
$\langle b|\widetilde{H}^{(n)}|b\rangle\leq\langle S_{n}|\widetilde{H}^{(n)}|S_{n}\rangle+o(1)=\lambda_{\min}(\widetilde{H}^{(n)})+o(1).$ (144)
Since $\{\widetilde{H}^{(n)}\}$ is gapped, and we already know its ground space dimension $2^{n-r_{n}}\leq\operatorname{dim}(\operatorname{span}(\mathcal{B}_{n}))$, by Lemma 8 we know $\dim(\operatorname{span}(\mathcal{B}_{n}))\leq 2^{n-r_{n}}$ when $n\in T$ is sufficiently large.
Therefore, when $n\in T$ is sufficiently large, we have
$2^{n-r_{n}}=\dim(\operatorname{span}(\mathcal{B}_{n}))=\text{ground space dimension of $\{H^{(n)}\}$}.$ (145)
Since we assume the ground space dimension of $\{H^{(n)}\}$ is not a power of 2, we get a contradiction. This allows us to conclude that for arbitrary ground state family $\big{\{}|\phi_{n}\rangle\big{\}}$ of $\{H^{(n)}\}$, any family of constant-depth local circuits $\{U_{n}\}$ and any family of stabilizer states $\{|S_{n}\rangle\}$, we have
$\big{\|}U_{n}|\phi_{n}\rangle\langle\phi_{n}|U_{n}^{\dagger}-|S_{n}\rangle\langle S_{n}|\big{\|}_{1}=\Omega(1/n),$ (146)
that is, $\big{\{}|\phi_{n}\rangle\big{\}}$ has strong LRM. ∎
Appendix D LRM from correlation
D.1 Two-qubit reduced density matrices of SRM states
For a quantum circuit $U$ acting on $n$ qubits and a subset $A\subset[n]:=\{1,2,\cdots,n\}$, we define the backward light-cone $\mathcal{L}(A)\subset[n]$ as the set of qubits that can influence the output of $U$ at $A$.
Theorem 16.
For a family of SRM states $\{|\psi_{n}\rangle=U_{n}|S_{n}\rangle\}$, where $\{U_{n}\}$ are constant-depth local circuits and $\{|S_{n}\rangle\}$ are stabilizer states, the following holds: For all subsystems $A,B\subset[n]$ of size $O(1)$ satisfying $\mathcal{L}(A)\cap\mathcal{L}(B)=\emptyset$, the bipartite reduced state $\operatorname{Tr}_{\overline{A\cup B}}|\psi_{n}\rangle\langle\psi_{n}|$ can be obtained by local operations on $O(1)$ EPR pairs $|\Phi^{+}\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$ shared by the two parties, where $\overline{A\cup B}:=[n]-(A\cup B)$.
To prove this theorem, note that (see Fig. 1(b))
$\operatorname{Tr}_{\overline{A\cup B}}|\psi_{n}\rangle\langle\psi_{n}|=\mathcal{E}_{A}\otimes\mathcal{E}_{B}\big{(}\operatorname{Tr}_{\overline{\mathcal{L}(A)\cup\mathcal{L}(B)}}|S_{n}\rangle\langle S_{n}|\big{)},$ (147)
where $\mathcal{E}_{A}(\sigma):=\operatorname{Tr}_{\mathcal{L}(A)-A}(U_{A}\sigma U_{A}^{\dagger})$ is a completely-positive trace preserving (CPTP) map form $\mathcal{L}(A)$ to $A$, with $U_{A}$ being the circuit within the light-cone of $A$ (colored orange in Fig. 1(b)), and $\mathcal{E}_{B}$ is defined similarly. $\operatorname{Tr}_{\overline{\mathcal{L}(A)\cup\mathcal{L}(B)}}|S_{n}\rangle\langle S_{n}|$ is a reduced state of a stabilizer state, and thus is a bipartite stabilizer mixed state Aaronson and Gottesman (2004) on $M=\left|\mathcal{L}(A)\right|+\left|\mathcal{L}(B)\right|=O(1)$ qubits. It can only create rigid correlations between $\mathcal{L}(A)$ and $\mathcal{L}(B)$, similar to those of EPR pairs, thereby restricting the possible states that can be created by performing local operations on it. By Lemma 17 in Appendix D.2, $\operatorname{Tr}_{\overline{\mathcal{L}(A)\cup\mathcal{L}(B)}}|S_{n}\rangle\langle S_{n}|$ can be created by local operations on $O(1)$ EPR pairs, and thus so can $\operatorname{Tr}_{\overline{A\cup B}}|\psi_{n}\rangle\langle\psi_{n}|$.
D.2 Bipartite stabilizer mixed states
Given the generating set $\{g_{1},\cdots,g_{r}\}$ for an $n$-qubit stabilizer code, where $r\leq n$ and $g_{1},\cdots,g_{r}$ are independent, mutually commuting Pauli strings such that $\langle g_{1},\cdots,g_{r}\rangle$ does not contain $-\mathbb{I}$, the corresponding stabilizer mixed state Aaronson and Gottesman (2004) is defined as $\frac{1}{2^{r}}\prod_{i=1}^{r}(\mathbb{I}+g_{i})$, which is the uniform distribution over all states within the code space. Stabilizer mixed states are in one-to-one correspondence with stabilizer codes.
Lemma 17.
A bipartite stabilizer mixed state $\rho$ on $N$ qubits can be obtained by local operations on at most $N$ EPR pairs.
Proof.
By Theorem 1 in Audenaert and Plenio (2005), every bipartite stabilizer mixed state $\rho_{AB}$ on $N=N_{A}+N_{B}$ qubits with $r$ independent generators is locally unitary equivalent to $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes p}\otimes\rho^{\prime}$, where $|\Phi^{+}\rangle=\frac{1}{\sqrt{2}}(|00\rangle+|11\rangle)$ is EPR pair shared by $A$ and $B$, and
$\rho^{\prime}=\frac{1}{2^{N-2p}}\prod_{l=2p+1}^{r}(\mathbb{I}+h_{l})$ (148)
is a bipartite separable stabilizer mixed state acting on $(N_{A}-p)+(N_{B}-p)$ qubits, with $h_{l}$ a tensor product of single-qubit $\mathbb{I}$ and $X$, acting on $N-2p$ qubits. The number of EPR pairs $p$ is upper bounded by $p\leq\min(\lfloor r/2\rfloor,N_{A},N_{B})$.
[…]
D.3 Generate correlations by EPR pairs
We take $P=Z$ without loss of generality.
Proposition 18.
For integer $K\geq 1$, $K$ EPR pairs can generate a two-qubit state $\rho$ by local CPTP maps, with $\langle Z\otimes Z\rangle_{\rho}=b$ and $\langle Z\otimes\mathbb{I}\rangle_{\rho}=\langle\mathbb{I}\otimes Z\rangle_{\rho}=c$, if and only if $(b,c)$ lies in the region given by
$1+2c+b\geq 0,$ (149)
$1-2c+b\geq 0,$ (150)
$\frac{1+2c+b}{4}2^{K}\leq l+(\frac{1+c}{2}2^{K}-l)^{2},\text{ for}l=0,1,\cdots,2^{K}-1.$ (151)
See Fig. 2 for the region given by constraints Eqs. (149)–(151).
Proof.
($\Rightarrow$) Suppose $\rho=\mathcal{E}_{A}\otimes\mathcal{E}_{B}\big{(}|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\big{)}$ for some local CPTP maps $\mathcal{E}_{A}$ and $\mathcal{E}_{B}$. Measure $\rho$ in the computational basis, we get $|00\rangle$ with probability $(1+2c+b)/4$, get $|11\rangle$ with probability $(1-2c+b)/4$, and get $|01\rangle$ or $|10\rangle$ with probability $(1-b)/4$. Therefore, we get a classical correlation Jain et al. (2013)
$\frac{1}{4}\begin{bmatrix}1+2c+b&1-b\\ 1-b&1-2c+b\\ \end{bmatrix}.$ (152)
To ensure (152) a valid correlation, each element should be non-negative. By measuring local POVMs $\{\mathcal{E}_{A}^{\dagger}(|0\rangle\langle 0|),\mathcal{E}_{A}^{\dagger}(|1\rangle\langle 1|)\}$ and $\{\mathcal{E}_{B}^{\dagger}(|0\rangle\langle 0|),\mathcal{E}_{B}^{\dagger}(|1\rangle\langle 1|)\}$ on $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}$, we can get the correlation (152). All $2\times 2$ classical correlations $M$ that can be obtained by $K$ EPR pairs, can be obtained by measuring local POVMs $\{E,\mathbb{I}_{2^{K}}-E\}$ and $\{F,\mathbb{I}_{2^{K}}-F\}$ acting on $K$ qubits on $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}$, and can be written as
$M=$ $\begin{bmatrix}\operatorname{Tr}\left(E\otimes F|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\right)&\operatorname{Tr}\left(E\otimes(\mathbb{I}_{2^{K}}-F)|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\right)\\ \operatorname{Tr}\left((\mathbb{I}_{2^{K}}-E)\otimes F|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\right)&\operatorname{Tr}\left((\mathbb{I}_{2^{K}}-E)\otimes(\mathbb{I}_{2^{K}}-F)|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}\right)\end{bmatrix}$ (153)
$=$ $2^{-K}\begin{bmatrix}\operatorname{Tr}\left(E^{\mathrm{T}}F\right)&\operatorname{Tr}\left(E\right)-\operatorname{Tr}\left(E^{\mathrm{T}}F\right)\\ \operatorname{Tr}\left(F\right)-\operatorname{Tr}\left(E^{\mathrm{T}}F\right)&2^{K}-\operatorname{Tr}\left(E\right)-\operatorname{Tr}\left(F\right)+\operatorname{Tr}\left(E^{\mathrm{T}}F\right)\end{bmatrix}.$ (154)
Letting (154) $=$ (152), we get
$\operatorname{Tr}(E)=\frac{1+c}{2}2^{K},$ (155)
$\operatorname{Tr}(F)=\frac{1+c}{2}2^{K},$ (156)
$\operatorname{Tr}(E^{\mathrm{T}}F)=\frac{1+2c+b}{4}2^{K},$ (157)
$0\leq E,F\leq\mathbb{I}_{2^{K}}.$ (158)
Note that $\operatorname{Tr}(E^{\mathrm{T}}F)=\operatorname{Tr}((E^{*})^{\dagger}F)\leq\left\|E^{*}\right\|_{2}\left\|F\right\|_{2}=\left\|E\right\|_{2}\left\|F\right\|_{2}$, where $E^{*}$ is the complex conjugate of $E$. For $l=0,1,\cdots,2^{K}-1$, when $l\leq\frac{1+c}{2}2^{K}\leq l+1$, we have $\left\|E\right\|_{2},\left\|F\right\|_{2}\leq\sqrt{l+(\frac{1+c}{2}2^{K}-l)^{2}}$. Therefore, we obtain
$\operatorname{Tr}(E^{\mathrm{T}}F)\leq l+(\frac{1+c}{2}2^{K}-l)^{2},$ (159)
giving the constraint $\frac{1+2c+b}{4}2^{K}\leq l+(\frac{1+c}{2}2^{K}-l)^{2}$. These $2^{K}$ constraints also imply $1-b\geq 0$.
($\Leftarrow$) Suppose $(b,c)$ lies in the region given by Eqs. (149)–(151). We claim that we can generate the correlation (152), by measuring some POVMs $\{E,\mathbb{I}_{2^{K}}-E\}$ and $\{F,\mathbb{I}_{2^{K}}-F\}$ on $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}$, that is, we can find $E$ and $F$ satisfying the constraints in Eqs. (155)–(158).
For $l=0,1,\cdots,2^{K-1}-1$, when $c$ satisfies $l\leq\frac{1+c}{2}2^{K}\leq l+1$, we have
$0\leq\frac{1+2c+b}{4}2^{K}\leq l+(\frac{1+c}{2}2^{K}-l)^{2}.$ (160)
Since $\operatorname{Tr}(E)=\operatorname{Tr}(F)=\frac{1+c}{2}2^{K}\leq 2^{K-1}$, we can take $E^{*}$ and $F$ such their support are disjoint and get $\operatorname{Tr}(E^{\mathrm{T}}F)=0$. By taking $E=F=\operatorname{diag}(1,\cdots,1,\frac{1+c}{2}2^{K}-l,0,\cdots,0)$ with $l$ 1’s, we get $\operatorname{Tr}(E^{\mathrm{T}}F)=l+(\frac{1+c}{2}2^{K}-l)^{2}$. Moreover, $\operatorname{Tr}(E^{\mathrm{T}}F)$ takes arbitrary values in the region $[0,l+(\frac{1+c}{2}2^{K}-l)^{2}]$ by appropriately choosing $0\leq E,F\leq\mathbb{I}_{2^{K}}$. Therefore, we get $E$ and $F$ satisfying the constraints in Eqs. (155)–(158).
For $l=2^{K-1},1,\cdots,2^{K}-1$, when $c$ satisfies $l\leq\frac{1+c}{2}2^{K}\leq l+1$, we have
$c2^{K}\leq\frac{1+2c+b}{4}2^{K}\leq l+(\frac{1+c}{2}2^{K}-l)^{2}.$ (161)
By taking $E=\operatorname{diag}(1,\cdots,1,c,\cdots,c)$ and $F=\operatorname{diag}(c,\cdots,c,1,\cdots,1)$ with $2^{K-1}$ 1’s and $2^{K-1}$ $c$’s, we get $\operatorname{Tr}(E)=\operatorname{Tr}(F)=\frac{1+c}{2}2^{K}$ and $\operatorname{Tr}(E^{\mathrm{T}}F)=c2^{K}$. By taking $E=F=\operatorname{diag}(1,\cdots,1,\frac{1+c}{2}2^{K}-l,0,\cdots,0)$ with $l$ 1’s, we get $\operatorname{Tr}(E^{\mathrm{T}}F)=l+(\frac{1+c}{2}2^{K}-l)^{2}$. Moreover, $\operatorname{Tr}(E^{\mathrm{T}}F)$ takes arbitrary values in the region $[c2^{K},l+(\frac{1+c}{2}2^{K}-l)^{2}]$ by appropriately choosing $0\leq E,F\leq\mathbb{I}_{2^{K}}$. Therefore, we get $E$ and $F$ satisfying the constraints in Eqs. (155)–(158).
Given $(b,c)$, measuring the corresponding POVMs $\{E,\mathbb{I}_{2^{K}}-E\}$ and $\{F,\mathbb{I}_{2^{K}}-F\}$ on $|\Phi^{+}\rangle\langle\Phi^{+}|^{\otimes K}$, get the correlation (152), and then prepare the diagonal state
$\rho:=\frac{1+2c+b}{4}|00\rangle\langle 00|+\frac{1-b}{4}|01\rangle\langle 01|+\frac{1-b}{4}|10\rangle\langle 10|+\frac{1-2c+b}{4}|00\rangle\langle 00|$ (162)
conditioned on the measurement results. We have $\langle Z\otimes Z\rangle_{\rho}=b$ and $\langle Z\otimes\mathbb{I}\rangle_{\rho}=\langle\mathbb{I}\otimes Z\rangle_{\rho}=c$, and $\rho$ is prepared by local CPTP maps on $K$ EPR pairs. ∎
D.4 $|\mathrm{C}^{n-1}Z\rangle$
The reduced density matrix of $|\mathrm{C}^{n-1}Z\rangle$ Wei and Liu (2024) on two qubits $\{k,l\}\subset[n]$ is
$\operatorname{Tr}_{\overline{\{k,l\}}}|\mathrm{C}^{n-1}Z\rangle\langle\mathrm{C}^{n-1}Z|=\big{(}1-\frac{1}{2^{n-2}}\big{)}|++\rangle\langle++|+\frac{1}{2^{n-2}}|\mathrm{C}Z\rangle\langle\mathrm{C}Z|.$ (163)
We have $\langle\mathrm{C}Z|X\otimes X|\mathrm{C}Z\rangle=0$, thus
$\langle X_{k}X_{l}\rangle=\operatorname{Tr}\left(X\otimes X\operatorname{Tr}_{\overline{\{k,l\}}}|\mathrm{C}^{n-1}Z\rangle\langle\mathrm{C}^{n-1}Z|\right)=1-\frac{1}{2^{n-2}}.$ (164)
On one qubit we have
$\operatorname{Tr}_{\overline{\{k\}}}|\mathrm{C}^{n-1}Z\rangle\langle\mathrm{C}^{n-1}Z|=\big{(}1-\frac{1}{2^{n-1}}\big{)}|+\rangle\langle+|+\frac{1}{2^{n-1}}|-\rangle\langle-|,$ (165)
thus
$\langle X_{k}\rangle=1-\frac{1}{2^{n-1}}-\frac{1}{2^{n-1}}=1-\frac{1}{2^{n-2}}.$ (166)
Appendix E NLTM as a strengthened necessary condition for quantum PCP
Here we present a straightforward argument that the validity of NLTM conjecture is necessary for quantum PCP to hold. That is, as remarked in the main text, NLTM is a stronger version of NLTS that gets closer to quantum PCP since it takes into account more general classical witnesses. This statement was also mentioned in a talk Nirkhe (2023).
Proposition 19.
Suppose $\mathsf{NP}\neq\mathsf{QMA}$ and the quantum PCP conjecture holds, then the NLTM conjecture holds.
The proof is adapted from Ref. Nirkhe (2022). Various necessary preliminary definitions can be found therein.
Proof.
Assume the contrary, namely, assume that NLTM is false while $\mathsf{NP}\neq\mathsf{QMA}$ and the quantum PCP conjecture holds.
Take an arbitrary QMA decision problem $L$. If the quantum PCP conjecture is true, there is a polynomial-time quantum reduction that maps an input instance $x$ of $L$ to a local Hamiltonian $H^{(n)}=\sum_{i=1}^{m^{(n)}}H_{i}^{(n)}$ acting on $n$ qubits ($H^{(n)}$ depends on $x$, $n$ relies on the size of $x$), with $m^{(n)}=\Theta(n)$ and $\|H_{i}^{(n)}\|_{\infty}\leq 1$. There are two energy thresholds $a^{(n)}$ and $b^{(n)}$ with $b^{(n)}-a^{(n)}\geq\gamma n$ for some constant $\gamma>0$. The promise for $H^{(n)}$ is:
  1. 1.
    YES instance: If $x\in L$, then $\lambda_{\min}(H^{(n)})\leq a^{(n)}$;
  2. 2.
    NO instance: If $x\notin L$, then $\lambda_{\min}(H^{(n)})\geq b^{(n)}$.
By false-NLTM assumption, we now show that $L\in\mathsf{NP}$.
Consider a family of YES instance $\{x\}$ and the corresponding $\{H^{(n)}\}$. For constant $\gamma>0$, by false-NLTM, for this family of Hamiltonian $\{H^{(n)}\}$, there exists a family of states $\{|\phi_{n}\rangle\}$ with
$\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle<\lambda_{\min}(H^{(n)})+\gamma n\leq a^{(n)}+\gamma n=b^{(n)}$ (167)
that has SRM. Suppose they can be written as $|\phi_{n}\rangle=U_{n}|S_{n}\rangle$ for families of constant-depth circuits $\{U_{n}\}$ and stabilizer states $\{|S_{n}\rangle\}$. The classical description of $U_{n}$ and $|S_{n}\rangle$ can serve as the NP witness for YES instance $H^{(n)}$. To see this, first notice that $\langle\phi_{n}|H^{(n)}|\phi^{(n)}\rangle<b^{(n)}$ is a convincing proof that $\lambda_{\min}(H^{(n)})\leq a^{(n)}$, due to the promise. Also, $\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle$ can be computed efficiently classically by the classical description of $U_{n}$ and $|S_{n}\rangle$, because
$\langle\phi_{n}|H^{(n)}|\phi_{n}\rangle=\sum_{i=1}^{m^{(n)}}\langle S_{n}|U_{n}^{\dagger}H_{i}^{(n)}U_{n}|S_{n}\rangle,$ (168)
where $\big{|}\operatorname{supp}\big{(}U_{n}^{\dagger}H_{i}^{(n)}U_{n}\big{)}\big{|}=O(1)$ implies $\langle S_{n}|U_{n}^{\dagger}H_{i}^{(n)}U_{n}|S_{n}\rangle$ can be computed efficiently by e.g. decompose $U_{n}^{\dagger}H_{i}^{(n)}U_{n}$ into the sum of $O(1)$ Pauli strings Aaronson and Gottesman (2004).
For a family of NO instance $\{x\}$ and the corresponding $\{H^{(n)}\}$, since $\lambda_{\min}(H^{(n)})\geq b^{(n)}$, arbitrary NP witnesses will not be accepted.
Therefore we conclude that $L\in\mathsf{NP}$. Since $L\in\mathsf{QMA}$ is taken arbitrarily, we have $\mathsf{QMA}\subset\mathsf{NP}$, contrary to the assumption that $\mathsf{NP}\neq\mathsf{QMA}$. ∎
Appendix F LRM and nonlocal magic
Given a partition of the whole system into several subsystems, we say a nonstabilizer state has completely nonlocal magic if its local reduced density matrices on subsystems have no magic. Here we point out that LRM is a stronger notion compared to completely nonlocal magic.
Suppose $\{|\psi_{n}\rangle\}$ has LRM, then for an arbitrary disjoint local partition $[n]=R_{1}\cup\cdots\cup R_{t_{n}}$, with $\left|R_{i}\right|=O(1)$, we can use a constant-depth local circuit $U^{(n)}=U_{1}\otimes\cdots\otimes U_{t_{n}}$ to rotate the reduced density matrix on each $R_{i}$ into a matrix that is diagonal in the computational basis, thereby exhibiting no magic on each $R_{i}$. Meanwhile, $U^{(n)}|\psi_{n}\rangle$ remains magical, because $U^{(n)}$ cannot remove all the magic of $|\psi_{n}\rangle$ when $n$ is sufficiently large.
On the other hand, some families of states with SRM can have completely nonlocal magic. For example, the logical $|T\rangle^{\otimes k}$ state in 3D color codes can be obtained by performing a single layer of non-Clifford gates on logical $|+\rangle^{\otimes k}$, a stabilizer state, thus having SRM. Note that its magic is completely nonlocal because it exhibits no magic when restricted to regions smaller than the code distance.